### v1

# Catalog of outbursts of neutron star LMXBs

Abstract
Many X-ray binaries are transiently accreting. Having statistics on their recurrence times is helpful to address questions related to binary evolution and populations, as well as the physics of binary systems. We compile a catalog of known outbursts of neutron star (identified through bursts or pulsations) low-mass X-ray binaries, until late 2023. Most outbursts are taken from the literature, but we also identify some outbursts from public X-ray monitoring lightcurves. We find 109 outbursts not previously identified in the literature; most are from the frequent transients GRS 1747-312 and the Rapid Burster MXB 1730-335, though we suspect that two outbursts from Liller 1 may be from another transient, besides the Rapid Burster. We also find new outbursts for 10 other systems, and verify substantial quiescent intervals for XMM J174457-2850.3, XMMU J174716.1-281048, and AX J1754.2-2754. Outburst detection has been relatively efficient since 1996 for outbursts above $F_{X}$(2-10)$=3\times 10^{-10}$ ergs/s/cm2. While several systems have many known outbursts, 40 of the 85 systems we track have zero or one recorded outburst between 1996 and 2023. This suggests that faint Galactic Center X-ray binaries may be neutron star X-ray binaries, though we cannot completely rule out the proposition that most neutron star X-ray binaries undergo frequent outbursts below all-sky monitor detection limits.
Low-mass X-ray binary stars – neutron stars –catalogs – accretion
facilities: RXTE(ASM and PCA), Swift(XRT), MAXI(GSC), Integral(JEM-X), CXO, XMM-Newton
1 Introduction
The total number of mass-transferring neutron star low-mass X-ray binaries in the Galaxy is thought to be of the order of $10^{4}$ (Portegies Zwart et al., 1997; Kiel & Hurley, 2006; Jonker et al., 2011; van Haaften et al., 2015). However, only a small fraction of these systems have been detected, typically during outbursts in which their X-ray luminosity increases by a factor of $10^{3}$$10^{7}$ compared to quiescence. Understanding the full population of these systems, and their binary evolution, is helpful for us to understand the physics of binary systems (e.g. Podsiadlowski et al., 2002; Postnov & Yungelson, 2014; Van et al., 2019), and also to understand the products of their evolution, such as millisecond pulsars (Alpar & Shaham, 1985; Tauris & Savonije, 1999; Tauris et al., 2012), which may produce the excess gamma-rays seen from the Galactic bulge (Wang et al., 2005; Gonthier et al., 2018). One question of current interest is the nature of X-ray sources of intermediate luminosity ($L_{X}$(2-10 keV)$\sim 10^{31}-10^{33}$ erg/s) detected near the Galactic Centre (Muno et al., 2005; Hailey et al., 2018; Mori et al., 2021; Maccarone et al., 2022). The objects in this region that show no outbursts have been argued to be mostly black hole LMXBs, based on an argument that neutron star LMXBs do not have long enough recurrence times. Having detailed statistics on X-ray outbursts from neutron stars can be helpful in beginning to answer questions such as these.
Large numbers of transient X-ray binaries have been detected by a wide range of X-ray surveys. For many of these X-ray binaries, the nature of the accreting object is not confidently known. However, many transient X-ray binaries have secure identifications of the accretor as neutron stars, typically through the identification of thermonuclear X-ray bursts from the neutron star surface (Lewin et al., 1993), or through detection of pulsation from the neutron star spin (Wijnands & van der Klis, 1998). Verification of the opposite case, that the accretor is a black hole, is harder, requiring optical or infrared spectroscopy to measure, through radial velocities, the mass of the unseen companion, and show it is above the generally accepted maximum neutron star mass of 3 M (Casares & Jonker, 2014).
A number of catalogs of X-ray binaries have been produced over the years. The catalogs of Liu et al. (2001, 2007) have been particularly well-cited, but are now 17 years old. Ritter & Kolb (2003) have an extensive and updated catalog of X-ray binaries and cataclysmic variables with known periods, but do not include objects without known (or suggested) orbital periods. More recent catalogs have been published of black hole X-ray binaries, one focused on compiling detailed lightcurves (Tetarenko et al., 2016b), and one focusing on multiwavelength studies (Corral-Santana et al., 2016). We are interested in the particular problem of understanding the frequency of X-ray outbursts from neutron star LMXBs, but found that we lacked a catalog to answer this question. The catalog of Avakyan et al. 2023 was released when this work was largely complete, and while it is very useful, it also does not list outbursts. Yan & Yu (2015) and Lin et al. (2019) construct tables of outburst frequency for a number of X-ray binaries with known orbital periods, but these works only include 14 and 19 NS LMXBs respectively. Yan & Yu (2015) do include a detailed listing of the outbursts they include, which is helpful. We identify several more outbursts from 4U 1608-52, XTE J1709-267, MXB 1730-33, XTE J1739-285, 1A 1744–361, SAX J1748.9-2021, and Aquila X-1 in the time range these authors cover, typically fainter outbursts.
§ 2 gives details of the construction of our catalog. In § 3, we make a first use of this catalog to address the question of the nature of intermediate luminosity X-ray sources near the Galactic Centre.
2 Catalog
To construct such a catalog, we used two clear methods to verify the presence of a NS in an LMXB; presence of thermonuclear bursts, and/or presence of coherent millisecond pulsations. We took lists of confirmed neutron star LMXBs from the MINBAR X-ray burst catalog (Galloway et al., 2020), which identified 115 known X-ray bursters using data from BeppoSAX, RXTE, and INTEGRAL (1996-2012), plus the literature.
We omit 42 NS LMXBs in the MINBAR catalog that appear to be persistent. We list these persistent, verified bursting NS XRBs in Table D, in Appendix D.
We omit a few transient bursters; the position of XB 1940-04 (Murakami et al., 1983) is very poorly known. The eccentric X-ray binary Cir X-1 (or 3A 1516-569) has outbursts that reflect its orbital period, not disk outbursts (Murdin et al., 1980). We exclude 1A 1742-289 (or A1742-289), as it is not clear that the burst activity seen from MXB 1743-29 in the Galactic Center in 1976 (Lewin et al., 1976a) can be attributed to this radio-bright (Davies et al., 1976) source (we note that this is definitely not the same object as AX J1745.6-2901; Kennea & Skinner 1996). This leaves 70 sources from the MINBAR catalog.
We add to the MINBAR list the known recent bursters 4U 1730-22, MAXI J1816-195, MAXI J0556-332, MAXI J0911-655, Terzan 6 X-2 (see Appendix § B), and SRGA J144459.2-604207 111as updated by Jean in ’t Zand; https://personal.sron.nl/~jeanz/bursterlist.html.  [^2] We include SRGA J144459.2-604207 in our list for completeness, but otherwise our summary of outbursts stops in 2023. We also add known transient accreting millisecond pulsars that are not known bursters, from the Patruno & Watts (2021) catalog and more recent literature, XTE J0929-314, NGC 6440 X-2, XTE J1751-305, Swift J1756.9-2508, XTE J1807-294, IGR J17494-3030, IGR J16597-3704, and MAXI J1957+032. Finally, we include the transient slow (2.14 Hz) accreting pulsar GRO J1744-28. (We also list the few persistent LMXBs that are pulsars in Table D.) This gives us a total of 85 transient neutron star low-mass X-ray binaries.
We performed literature searches for this list of transients, and identify the known X-ray outbursts from each system. For each outburst, we attempt to determine the peak observed 2-10 keV X-ray luminosity. This will have some significant uncertainty, since not every outburst is caught at its peak, some outbursts are observed only in other energy ranges, and many distances are not well-known. We attempt to use the best-fit spectrum to convert measurements to 2-10 keV fluxes, but we only attempt this with simple power-law spectra (we do not try to convert complex Comptonized models). Where information is lacking, we use the best literature $N_{H}$ estimate and an assumed photon index of 2.1 to make conversions, or use literature estimates of the 2-10 keV flux in mCrab (assuming 1 Crab = $2\times 10^{-8}$ ergs cm-2 s-1). A few bursters have recorded bursts without any recorded outbursts (Gotthelf & Kulkarni, 1997; Cornelisse et al., 2002b); we list these in Table C in Appendix C.
We present the catalog of known outbursts of neutron star LMXBs in Table 1, for which we provide a sample here. The full table is in the electronic version of the journal (or Appendix A for the arXiv version). We include a line for each recorded outburst, with a literature reference for each outburst if available. Several objects with repeated bright outbursts can be easily spotted in the public lightcurves of the RXTE PCA Bulge Scan (Swank & Markwardt, 2001) [^3] https://asd.gsfc.nasa.gov/Craig.Markwardt//galscan/main.html, the RXTE All-Sky Monitor (Levine et al., 1996) [^4] http://xte.mit.edu; comments at http://xte.mit.edu/XTE/xte_anno.html and/or in MAXI GSC observations (Matsuoka et al., 2009; Suguzaki et al., 2011). We have added some obvious outbursts visible in these surveys to the list (e.g. outbursts from XTE J1709-267 in 2005, 2009, 2018/19, and 2022). The line for defining an ‘outburst’ can be unclear; for simplicity we require an increase of a factor of 10 from the quiescent flux (if known), and $L_{X}>10^{34}$ erg/s.
We separate the objects with many outbursts into separate tables, placed in Appendix B. For the most frequently repeating objects with bright outbursts (of order once per year or more), 4U 1608-52, MXB 1730-33 (the Rapid Burster in Liller 1), Aquila X-1, and GRS 1747-312 (the eclipser in Terzan 6), we use the three public lightcurves above, as well as Swift/XRT public lightcurves (Evans et al., 2007) [^5] https://www.swift.ac.uk/user_objects/, and the INTEGRAL JEM-X Galactic Bulge Monitoring lightcurves (Kuulkers et al., 2007) [^6] http://integral.esac.esa.int/BULGE/ to develop the best possible list of outbursts from 1996 to 2023, along with literature identification of outbursts before 1996. For the objects with the most frequent outbursts, we also record the MJD date of the outburst peak (as best we can tell). We place the outbursts of NGC 6440 X-2 in a separate table in Appendix B as well, as their high frequency requires MJD dates to differentiate them.
When identifying additional outbursts, we generally search for multiple points that are above 3 sigma. The MAXI baseline may often vary substantially, so we additionally impose that the potential outburst must rise significantly above the quiescent baseline.
We find additional outbursts, not mentioned in the literature, for 12 systems; XTE J1709-267 (4), XTE J1701-407 (1), XTE J1739-285 (2), 1A 1744-361 (2), XMMU J174716.1-281048 (1), AX J1754.2-2754 (3), SAX J1810.8-2609 (1), 4U 1608-52 (3), MXB 1730-335 (aka the Rapid Burster, 52), GRS 1747-312 (37), NGC 6440 X-2 (2), and Aquila X-1 (1), for a total of 109 newly identified outbursts.
We give comments on each X-ray binary’s outburst history in Appendices A and B, including the reference we use for the distance, and any discrepancies we identify in the literature regarding their outburst histories. Here we highlight a few key points;
$\bullet$ XMM J174457-2850.3, which showed frequent low-level outbursting behavior between 2001 and 2012, has seen its activity subside, with only one short outburst since 2012.
$\bullet$ We use Chandra, XMM-Newton, and Swift/XRT observations to verify that XMMU J174716.1-281048 declined from a $\sim$10-year-long outburst in 2012, reached full quiescence by 2018, showed another (short, $<$3 months) outburst in 2021, and then returned again to quiescence.
$\bullet$ We use Swift/XRT and Chandra observations to show that AX J1754.2-2754 has had $L_{X}$(2-10 keV)$<2\times 10^{33}$ erg/s for much of the past eight years, in contrast with its quasi-persistent behavior in 1999-2017 (Bassa et al., 2008).
$\bullet$ SAX J1806.5-2215’s long outburst, which started in Feb. 2011, ended sometime between July 2017 and January 2018.
$\bullet$ We catalog 52 outbursts from MXB 1730-335 (aka the Rapid Burster) that have not previously been identified in the literature. Some of these may not be produced by the Rapid Burster, but by other LMXBs in the host cluster Liller 1; another such outbursting transient has been verified (Homan et al., 2018), and we do identify two unusually short intervals (28 and 37 days) between outbursts that are good candidates for other transients. However, the relatively regular outbursts, and detection of Type II bursts in most cases where instruments are sensitive to them, indicate most outbursts are due to the Rapid Burster. Masetti (2002) noted that the average outburst interval decreased from 200 days to 100 days around 2000, with the peak flux also dropping by a factor of 2. Since 2000, the average outburst interval has varied between 100 and 170 days, and the peak $L_{X}$ has remained around $2\times 10^{37}$ erg/s.
$\bullet$ We catalog 37 previously unidentified outbursts from GRS 1747-312 (in the globular cluster Terzan 6). The outburst interval averages 138 days, with significant scatter, but regular enough to indicate that most outbursts are indeed due to a single source. A second transient bursting source, Terzan 6 X-2, has recently been identified in Terzan 6 Saji et al. (2016); van den Berg et al. (2024), but appears not to reach such high fluxes as GRS 1747-312.
In §3 we use the information in these tables to draw some general conclusions about our sensitivity to X-ray outbursts, and apply this information to consideration of interesting X-ray sources near the Galactic Center.
NS XRB Begin End Peak $F_{X}$ $N_{H}$ Index $L_{X}$ d References
2-10 keV cm-2 2-10 keV kpc
IGR J00291+5934 1998/11 1998/11 4.5e-10 - - 9.6e35 4.2 Remillard (2004)
IGR J00291+5934 2001/9 2001/9 3.9e-10 - - 8.2e35 4.2 Remillard (2004)
IGR J00291+5934 2004/12 2004/12 6.1e-10 7e21 1.7 1.3e36 4.2 Eckert et al. (2004); Galloway et al. (2005)
IGR J00291+5934 2008/8 2008/10 6.1e-10 - 1.7 1.3e36 4.2 Lewis et al. (2010)
IGR J00291+5934 2015/8 2015/8 5.6e-10 1e22 1.7 1.2e36 4.2 De Falco et al. (2017)
MAXI J0556-332 2011/1 2012/5 1.9e-09 - - 4.4e38 43.6 Matsumura et al. (2011); Sugizaki et al. (2013)
MAXI J0556-332 2012/10 2012/12 4e-10 3.2e20 - 9e37 43.6 Sugizaki et al. (2012); Parikh et al. (2017a)
MAXI J0556-332 2016/1 2016/3 7.7e-10 - - 1.8e38 43.6 Negoro et al. (2016a); Russell et al. (2016)
MAXI J0556-332 2020/04 2020/11 8.6e-10 - - 2.0e38 43.6 Page et al. (2022)
EXO 0748-676 1984/7 2008/8 1.1e-9 1.2e21 1.9 7e36 7.4 Parmar et al. (1986); Degenaar et al. (2009)
GS 0836-429 1970/12 1972/1 1.1e-9 $<$3.5e22 1.1e37 $<$9.2 Markert et al. (1977)
GS 0836-429 1990/10 1991/5 3.8e-10 2.2e22 1.5 3.9e36 $<$9.2 Aoki et al. (1992); Belloni et al. (1993)
GS 0836-429 2003/1 2003/5 1.0e-9 2.7e22 1.5 1.0e37 $<$9.2 Aranzana et al. (2016)
GS 0836-429 2003/9 2004/6 1.3e-9 2.7e22 1.5 1.3e37 $<$9.2 Aranzana et al. (2016)
MAXI J0911-655 2016/2 2023/10 1.7e-10 2.7e21 1.27 1.9e36 9.5 Sanna et al. (2017b); Ng et al. (2021a),MAXI
XTE J0929-314 2002/4 2002/4 6.7e-10 7.6e20 1.7-2 $>$2.9e36 $>$6 Galloway et al. (2002); Juett et al. (2003)
SAX J1324.5-6313 2015/8 2015/8 1.3e-10 1.5e22 - 1.0e36 $<$6.2 Negoro et al. (2015a); Cornelisse et al. (2002a)
MAXI J1421-613 2014/01 2014/1 1.9e-9 4.8e22 2.1 2e36 3 Serino et al. (2015); Nobukawa et al. (2020)
SRGA J144459.2-604207 2022/1 2022/1 4e-10 - - 3.4e36 8.5 Negoro et al. (2024)
SRGA J144459.2-604207 2023/12 2023/12 6e-10 - - 5.2e36 8.5 Negoro et al. (2024); Sguera & Sidoli (2024)
SRGA J144459.2-604207 2024/2 2024/3 2.7e-9 2.9e22 1.8 2.3e37 8.5 Molkov et al. (2024); Ng et al. (2024)
Cen X-4 1969/07 1969/7 6e-7 - - 1.2e38 1.3 Evans et al. (1970); Kuulkers et al. (2009)
Cen X-4 1979/05 1979/5 9.6e-8 9e20 - 1.9e37 1.3 Kaluzienski et al. (1980); Kuulkers et al. (2009)
Table 1: Published outbursts of verified neutron star LMXBs. Max $F_{X}$ is highest reported flux in each outburst, transformed to 2-10 keV, generally using the power-law spectral index and $N_{H}$ noted. We assume 1 Crab is $F_{X}$(2-10)=$2\times 10^{-8}$ erg s-1 cm-2. Peak observed intrinsic $L_{X}$ in 2-10 keV, for the assumed distance (listed). The full Table is given in the appendix (for arXiv), or online (for journal); a portion is reproduced here.
3 Discussion
3.1 General trends
In Fig. 1 we show the peak fluxes from all transient neutron star LMXBs (a few lie outside our bounds), with the most frequent outbursters indicated with different colors. In Fig. 2, we show the number of outbursts per year. We first review key information about the relevant monitoring instruments.
RXTE/ASM monitoring produced a dramatic increase in detected outbursts, from January 1996 until 2009-2011  [^7] The RXTE/ASM sensitivity decayed from 2009 until RXTE was turned off in 2011.. RXTE/ASM achieved roughly 15 mCrab ($F_{X}=3\times 10^{-10}$ ergs/cm2/s) 3-sigma sensitivity per day, in 2-10 keV (Levine et al., 1996). Indeed, detections from 1996 on now did reach $3\times 10^{-10}$, though completeness may have been a little higher, perhaps $6\times 10^{-10}$ ergs/cm2/s. The RXTE/ASM covered most sources daily, except for about 30 days each year when the Sun was nearest for sources in the ecliptic.
From 1996 to 2002, in each spring and fall, BeppoSAX conducted regular monitoring of the Galactic Bulge with its Wide-Field Camera (WFC), with a 40o by 40o field of view, 5’ angular resolution, and daily sensitivity of about 15 mCrab in 2-28 keV (Heise et al., 1999; in’t Zand et al., 2004). BeppoSAX’s WFC discovered a large number of faint ($L_{X}<10^{37}$ erg/s) X-ray transients in the Galactic Bulge, and also several X-ray bursts from transients that were too faint for the WFC to detect individually (Cornelisse et al., 2002b).
From 1999 to 2011, RXTE also conducted monitoring of the Galactic Bulge using slews of the PCA instrument (Swank & Markwardt, 2001). In practice, these typically reached a sensitivity of 5 mCrab, $10^{-10}$ ergs/cm2/s, 2.5-10 keV. Notably, the RXTE/PCA was able to detect many faint outbursts of NGC 6440 X-2 in 2009-2011.
The INTEGRAL Galactic Bulge hard X-ray monitoring program was particularly productive, with 12’ angular resolution and 5-15 mCrab daily sensitivity in 20-60 keV (Kuulkers et al., 2007), though it could observe the Bulge only 40% of the year, in the spring and fall. Twelve of these transients have principal names starting with IGR (only XTE has more, with 13), suggesting INTEGRAL’s impact. There appears (Fig. 2) to be a small decline in detections in the past few years, which may be due to declining sensitivity of INTEGRAL’s IBIS/ISGRI instrument. [^8] The Integral Galactic Bulge Monitoring Program’s website, http://integral.esac.esa.int/BULGE/, presents 18-40 keV IBIS/ISGRI lightcurves for sources, which show increasing noise from 2016 and stop in 2018.
Swift/BAT’s hard X-ray (15-50 keV) monitoring program has operated since 2006, with a daily 16 mCrab detection limit (Krimm et al., 2013). MAXI/GSC’s 2-20 keV monitor, with a 9 mCrab daily 3-sigma sensitivity (Sugizaki et al., 2013), has been operating since 2009.
From Fig. 2, we see that 1996 marks a dramatic change in detected outburst frequency. A significant peak appears around 2009-2011 in all outbursts. Fig. 2 of Tetarenko et al. (2016b), transient outbursts of BH LMXBs, reveals a similar peak in 2007-2012. However, this peak is not significant in brighter NS LMXB outbursts with $F_{X}>3\times 10^{-10}$ ergs/s/cm2, suggesting that fainter outbursts were more detectable in that time frame (e.g. by the RXTE PCA bulge monitoring program; which detected, e.g., the NGC 6440 X-2 outbursts at $1-2\times 10^{-10}$ ergs/s/cm2; Heinke et al. 2010).
We then construct a histogram of fluxes from all outbursts since 1996 (Fig. 3, blue triangles). This is sharply peaked at $F_{X}=10^{-9}$ ergs/s/cm2, which might suggest that our data becomes highly incomplete below $F_{X}=10^{-9}$ ergs/s/cm2. However, if we exclude the four most frequent outbursters (Aquila X-1, 4U 1608-52, MXB 1730-335, and GRS 1747-312), all of which have relatively long orbital periods, high mass transfer rates, and high outburst fluxes, we see a rather different histogram (Fig. 3, red squares), which does not peak until the bin $F_{X}=2-4\times 10^{-10}$ ergs/s/cm2. This suggests that significant incompleteness is happening in the next lower bin, covering $F_{X}=1-2\times 10^{-10}$ ergs/s/cm2.
We can then look at the distribution of $L_{X}$ in detected outbursts (Fig. 4). (We must keep in mind that the majority of these $L_{X}$ values are likely overestimates, due to the assumed distances being often upper limits, if the observed bursts were Eddington-limited.) The $L_{X}$ distribution looks generally similar to the distribution of fluxes, due to the concentration of transient LMXBs in the Galactic Bulge, but due to the variation in distance, the peak is less sharp. An equivalent plot for BH LMXB systems (Fig. 15 in Tetarenko et al. 2016b [^9] Including a shift down by roughly a factor 3 to move from bolometric to 2-10 keV luminosity.) peaks close to $L_{X}\sim 10^{39}$ erg/s, with a tail extending down to $10^{35}$ erg/s. The difference between BHs and NSs may be attributed to the bias towards longer orbital periods among BHs, which are correlated with higher peak $L_{X}$ (Wu et al., 2010). Black hole X-ray binaries may have longer orbital periods with main sequence donor stars, because unstable mass transfer sets in for mass ratios greater than about 0.8-1.0 for typical stellar structures on the lower main sequence; and at very short periods, even the expected peak luminosities leave black holes in radiatively inefficient states, leading to a strong selection effect against these objects (e.g. Arur & Maccarone 2018). Another selection effect is that verifying a black hole in an LMXB typically requires phase-resolved spectroscopy, which is more difficult in a short-period, thus fainter, system; while verifying a neutron star uses X-ray bursts or pulsations, which aren’t tightly correlated with peak $L_{X}$.
Intriguingly, there is a sharper peak in the $1-2\times 10^{36}$ erg/s bin than appears in the flux histogram. If this peak were due entirely to limited sensitivity below it, we would expect this peak to be less sharp in the $L_{X}$ histogram than the $F_{X}$ histogram. It is possible (but not certain) that the number of outbursts falls off below $10^{36}$ erg/s (”very faint” X-ray binary outbursts). A gap in the (persistent) X-ray luminosity function between $10^{34}$ and $10^{36}$ erg/s has previously been identified in globular clusters (Hertz & Grindlay, 1983; Verbunt & Lewin, 2006; Bahramian et al., 2020b), where the sources above, and just below, this gap consist predominantly of neutron star LMXBs. A dropoff below $L_{X}\sim 10^{36}$ erg/s has also been suggested in the bulge of M31 (Voss & Gilfanov, 2007), in globular clusters in nearby galaxies (Zhang et al., 2011), and possibly (but with inadequate data) for the Milky Way (Sazonov et al., 2006; Bahramian et al., 2021). On the other hand, extensive Swift, XMM-Newton and Chandra monitoring of the Galactic Center identified 28 outbursts from 13 LMXBs, between 2000 and 2014, including 11 between $L_{X}=10^{34}$ erg/s and $10^{35}$ erg/s, which argues strongly against any cutoff (Wijnands et al., 2006; Degenaar et al., 2015). This apparent discrepancy may be reconciled if most persistently-bright systems have $L_{X}>10^{36}$ erg/s, but transient outbursts are often fainter, so that a given snapshot would find a gap, but a list of peak outburst luminosities would not. Following the standard disk-instability model, the only LMXBs that could remain persistently below $10^{36}$ erg/s will be a small range of periods for ultracompact systems in’t Zand et al. (e.g. 2007); van Haaften et al. (e.g. 2013). If this is indeed the case, then more sensitive X-ray monitoring programs (e.g. proposed for STROBE-X, Ray et al. 2019) would find substantially more X-ray outbursts.
Figure 1: Peak fluxes by year from transient neutron star LMXBs. Green, cyan, blue, magenta, and black triangles indicate outbursts by 4U 1608-52, MXB 1730-335, GRS 1747-312, NGC 6440 X-2, and Aquila X-1. Red squares indicate all other transient LMXB outbursts.
Figure 2: Number of total outbursts per year (blue triangles) and outbursts with $F_{X}>3\times 10^{-10}$ ergs/cm2/s per year (red dots), from transient neutron star LMXBs. The effective periods in which the RXTE/ASM, RXTE/PCA bulge scans, INTEGRAL galactic bulge monitoring, Swift/BAT monitor, and MAXI/GSC monitoring were effective are indicated. Since 1996, detections of bright outbursts have been roughly constant in time.
Figure 3: Number of outbursts by peak flux since 1996, for all transient neutron star LMXBs (blue triangles), and for all except the four bright regular outbursting systems: Aquila X-1, 4U 1608-52, MXB 1730-335, and GRS 1747-312 (red squares).
Figure 4: Number of outbursts by peak $L_{X}$ since 1996, for all transient neutron star LMXBs (blue triangles), and for all except the four bright regular outbursting systems: Aquila X-1, 4U 1608-52, MXB 1730-335, and GRS 1747-312 (red squares).
Figure 5: Number of neutron star LMXBs with a given number of total outbursts since 1996. Off the scale to the right are 4U 1608-52 (26), Aquila X-1 (35), GRS 1747-312 (69), and MXB 1730-335 (74).
3.2 Application to the Galactic Center
Hailey et al. (2018) identified a dozen Chandra X-ray sources within a few parsecs of the Galactic Center that have non-thermal X-ray spectra (best fit with a power law with photon index between 1.5 and 2.9), which along with their $L_{X}$ values indicates these are quiescent LMXBs. Hailey et al. (2018) and Mori et al. (2021) suggest that these objects must be black hole LMXBs, arguing that all neutron star LMXBs have short recurrence times, generally less than 5 years, while black hole LMXBs can have longer recurrence times. Maccarone et al. (2022) argued that a number of known neutron star LMXB systems do have longer recurrence times, and that the non-thermal quiescent Galactic Center X-ray sources may be dominated by neutron star LMXBs. Mori et al. (2021) and Mori et al. (2022) argued that most neutron star LMXBs undergo frequent outbursts, based on the history of the known NS transients studied in the Galactic Center. This argument assumes that the confirmed neutron star transients are representative of the underlying population, and not subject to biases; for instance, that the more frequent outbursters are easier to confirm (e.g. they produce more X-ray bursts).
Clearer conclusions on the recurrence times of neutron star LMXBs require a thorough compilation of known outbursts, which we performed in this paper, allowing us to comment on the fractions of (observed) neutron star LMXBs with different recurrence times. These conclusions are, however, subject to various systematic uncertainties, due principally to the limited sensitivity and angular resolution of X-ray all-sky monitors.
A key question we can ask is what fraction of transient neutron star LMXBs have a recurrence time greater than 10 years, from the data available. We can answer this by checking for gaps of at least 10 years between 1996 and 2022 (the time frame with the best sensitivity). Among the 85 systems we can study, only 24 do not show such a gap, suggesting that 72% could hide for 10 years or longer. Naive direct application of this to the Galactic Centre suggests that for every NS LMXB with a shorter recurrence time (e.g. the 6 known in the Galactic Centre), there are at least three times more systems with longer recurrence periods. This is a lower limit, which does not account for systems that have been quiescent over the past 30+ years, such as the quiescent LMXBs observed in many globular clusters (e.g. Heinke et al., 2003b; Guillot et al., 2011; Bahramian et al., 2015).
This is a very rough estimate, subject to numerous caveats. Most importantly, all-sky monitors have limited sensitivity. We infer that we are missing most outbursts below $L_{X}\sim 2.3\times 10^{36}*(d/(8{\rm kpc}))^{2}$ erg/s; even brighter outbursts for $\sim$10of the year when the Sun is nearby; and some (not easily quantifiable) due to crowding with bright nearby sources. The Swift campaigns of the Galactic Center show that the three verified NS LMXBs there show fainter outbursts as well as bright outbursts; Degenaar et al. (2015) list 4 outbursts with peak $L_{X}>2.3\times 10^{36}$ erg/s, and 11 outbursts with peaks below this. Since 1996, our larger databases give 8 and 22, respectively. There are also nine other transients, showing very-faint X-ray outbursts (peak $L_{X}<10^{36}$ erg/s), that have been detected in the Galactic Centre area since 2000, which have not been confirmed as neutron stars through X-ray bursts, nor show evidence for being black holes (Degenaar et al., 2015). It is generally difficult to attribute X-ray bursts observed near the Galactic Center with RXTE (Galloway et al., 2008), and very faint X-ray binaries may be expected to burst only rarely (Peng et al., 2007; Lin et al., 2019). It seems likely that some (or most) of these very faint X-ray binaries are neutron stars, given that known black hole outbursts tend to be brighter than known neutron star outbursts (Yan & Yu, 2015). However, we have observed few short-period black hole systems, and it is possible that these are relatively faint; cf. Knevitt et al. 2014. These 9 very-faint transients have each had only one or two outbursts within the past 23 years, suggesting relatively long recurrence times ($>$20 years), and that a larger population of such transients may exist there.
For most neutron star transients elsewhere, we cannot constrain their history of fainter outbursts peaking at $L_{X}<2\times 10^{36}$ erg/s. Only for three nearby neutron stars accreting from low-mass companions are we able to rule out outbursts since 1996 above $L_{X}>7\times 10^{34}$ ergs/s; the LMXB Cen X-4 (at 1.3 kpc), and the two transitional millisecond pulsars PSR J1023+0038 (Archibald et al., 2009) and PSR J12270-4859 (Roy et al., 2015), both at 1.4 kpc. It is unclear if these two transitional millisecond pulsars undergo regular, or even any, major accretion outbursts, as IGR 18245-2452 did. These three objects suggest that most neutron stars accreting from low-mass companions remain in a quiescent state for long periods.
Cen X-4’s X-ray properties are fully consistent with the nonthermal Galactic Center sources described by Mori et al. (2021); Cen X-4’s thermal emission would be completely absorbed by the $N_{H}>10^{23}$ at the Galactic Center, while its harder, variable power-law tail with photon index 1-1.7 and 2-10 keV $L_{X}=5-50\times 10^{31}$ (Cackett et al., 2010; Chakrabarty et al., 2014) resembles the properties of the Galactic Center sources. Incidentally, the same is true for many other qLMXBs, including some detected in globular clusters (e.g. EXO 1745-248 and Swift J174805.3-244637 in Terzan 5; Wijnands et al. 2005a; Bahramian et al. 2014). While it is true that most identified globular cluster qLMXBs that have not experienced outbursts lack significant hard spectral tails (Mori et al., 2022), this is likely a selection effect, as globular cluster qLMXBs without detected outbursts are generally selected by their extremely soft spectra (e.g. Heinke et al., 2003b; Guillot et al., 2011). The deepest cluster X-ray observations, of 47 Tuc, indeed identified two strong candidate qLMXBs with significant power-law tails (index $\sim$2, $L_{X}=2-3\times 10^{31}$ erg/s), and no recorded outbursts (Heinke et al., 2005). 47 Tuc is far enough from the ecliptic that it is never blocked from view for all-sky X-ray monitors, and at 4.5 kpc, 20 mCrab comes to $10^{36}$ erg/s, such that even relatively faint outbursts would have been observed.
Among the other transient neutron star LMXBs with long ($>$10-year) periods between known outbursts, we can identify several more with substantial power-law components in their spectra; XTE J0929-314 (Wijnands et al., 2005b), SAX J1324.5-6313 (Cornelisse et al., 2002a; Bahramian & Heinke, 2015), XB 1733-30 (CX2 in Terzan 1; Cackett et al. 2006; Bozzo et al. 2020, GRO J1744-28 (Doroshenko et al., 2020), and 1RXS J180408.9-342058 (Parikh et al., 2017b).
We note that the XB-NEWS group has conducted frequent optical monitoring of Cen X-4 since 2008 without a proper outburst, although they did observe a slow optical brightening culminating with a failed outburst in late 2020, peaking at $L_{X}\sim 3\times 10^{33}$ erg/s (Baglio et al., 2022); this serves to illustrate that the XB-NEWS optical monitoring (Lewis et al., 2008; Russell et al., 2019) is effective at identifying outbursts. XB-NEWS monitors 50 systems weekly, including a number of other NS LMXB transients detectable in the optical (such as IGR J00291+59, XTE J2123-058, XTE J0929-314, XTE J1814-338, and HETE J1900.1-2455; Lewis et al. 2008). Careful analysis of the XB-NEWS archive may enable determination of how many fainter outbursts we are missing.
4 Conclusions
We have compiled a list of known neutron star transient LMXBs, from the lists of known bursters and accreting pulsars. We have searched the literature to produce a list of known outbursts from these systems, for each of which we record the date, approximate 2-10 keV peak flux and (inferred, given the best-estimate distance) luminosity, and one (or a few) relevant references. For a few objects, we have identified outbursts not identified in the literature, using public (generally all-sky) lightcurves. The purpose of this list of outbursts is to facilitate more in-depth analyses of the outburst records. Such studies could be done using only the information from these tables, in which case it is important that such investigators bear in mind the non-uniformity and limited sensitivity of these lists, or could use the information and references here as a jumping-off point for more detailed study of one or more objects.
We use the information assembled to make a few new conclusions. For instance, we consider arguments about the distance to 4U 1608-52, and identify periods of prolonged quiescence in AX J1754.2-2754, XMM J174457-2850.3, and XMMU J174716.1-281048 (aka IGR J17464-2811). We then look at the distribution of outbursts with time, and flux, to identify a rough empirical flux completeness limit for X-ray binary outbursts since 1996 of $F_{X}=3\times 10^{-10}$ ergs/s/cm2 (though a number of outbursts have been identified below this limit, and some above have been missed).
Finally, we look at the distribution of the frequency of recorded outbursts from these transients. While four transients average roughly one to three outbursts per year, nearly half (40 of 85 systems) have zero or one recorded outburst between 1996 and 2023. It is certainly possible that many of these have shown fainter outbursts below our detection level during this period. However, relatively stringent limits on outbursts from the nearby transient LMXB Cen X-4 for several decades suggest that there are transient neutron star LMXBs with properties consistent with the intriguing Galactic Center X-ray binaries identified by Hailey et al. (2018).
Swift/XRT lightcurves supplied by the UK Swift Science Centre at the University of Leicester (Evans et al., 2007). RXTE/ASM results provided by the ASM/RXTE teams at MIT and at the RXTE SOF and GOF at NASA’s GSFC. RXTE/PCA results from NASA Goddard Space Flight Center, https://asd.gsfc.nasa.gov/Craig.Markwardt//galscan/main.html. INTEGRAL/JEM-X lightcurves from Kuulkers et al. (2007) using the OSA10.1 software, distributed by the ISDC (Courvoisier et al., 2003). MAXI results provided by RIKEN, JAXA and the MAXI team (Matsuoka et al., 2009). This paper made very extensive use of the Astrophysical Data Service (https://ui.adsabs.harvard.edu), ArXiv (http://arxiv.org), and the Astronomer’s Telegram (http://www.astronomerstelegram.org). CH is supported by NSERC Discovery Grant RGPIN-2023-04264. JZ has been supported by a Mitacs Globalink Research Internship. In Appendix A we give the full table (Table A) of data on neutron star LMXB outbursts, except for the frequent outbursters, which are given in Appendix B. Appendix C tabulates information on bursters without recorded outbursts, and Appendix D tabulates the neutron star X-ray binaries that we consider to be persistent systems (not further discussed in this paper).
Appendix A Outburst details
[…]
NS XRB Begin End Peak $F_{X}$ $N_{H}$ Index $L_{X}$ d References
2-10 keV cm-2 2-10 keV kpc
IGR J00291+5934 1998/11 1998/11 4.5e-10 - - 9.6e35 4.2 Remillard (2004)
IGR J00291+5934 2001/9 2001/9 3.9e-10 - - 8.2e35 4.2 Remillard (2004)
IGR J00291+5934 2004/12 2004/12 6.1e-10 7e21 1.7 1.3e36 4.2 Eckert et al. (2004); Galloway et al. (2005)
IGR J00291+5934 2008/8 2008/10 6.1e-10 - 1.7 1.3e36 4.2 Lewis et al. (2010)
IGR J00291+5934 2015/8 2015/8 5.6e-10 1e22 1.7 1.2e36 4.2 De Falco et al. (2017)
MAXI J0556-332 2011/1 2012/5 1.9e-09 - - 4.4e38 43.6 Matsumura et al. (2011); Sugizaki et al. (2013)
MAXI J0556-332 2012/10 2012/12 4e-10 3.2e20 - 9e37 43.6 Sugizaki et al. (2012); Parikh et al. (2017a)
MAXI J0556-332 2016/1 2016/3 7.7e-10 - - 1.8e38 43.6 Negoro et al. (2016a); Russell et al. (2016)
MAXI J0556-332 2020/04 2020/11 8.6e-10 - - 2.0e38 43.6 Page et al. (2022)
EXO 0748-676 1984/7 2008/8 1.1e-9 1.2e21 1.9 7e36 7.4 Parmar et al. (1986); Degenaar et al. (2009)
GS 0836-429 1970/12 1972/1 1.1e-9 $<$3.5e22 1.1e37 $<$9.2 Markert et al. (1977)
GS 0836-429 1990/10 1991/5 3.8e-10 2.2e22 1.5 3.9e36 $<$9.2 Aoki et al. (1992); Belloni et al. (1993)
GS 0836-429 2003/1 2003/5 1.0e-9 2.7e22 1.5 1.0e37 $<$9.2 Aranzana et al. (2016)
GS 0836-429 2003/9 2004/6 1.3e-9 2.7e22 1.5 1.3e37 $<$9.2 Aranzana et al. (2016)
MAXI J0911-655 2016/2 2023/10 1.7e-10 2.7e21 1.27 1.9e36 9.5 Sanna et al. (2017b); Ng et al. (2021a),MAXI
XTE J0929-314 2002/4 2002/4 6.7e-10 7.6e20 1.7-2 $>$2.9e36 $>$6 Galloway et al. (2002); Juett et al. (2003)
SAX J1324.5-6313 2015/8 2015/8 1.3e-10 1.5e22 - 1.0e36 $<$6.2 Negoro et al. (2015a); Cornelisse et al. (2002a)
MAXI J1421-613 2014/01 2014/1 1.9e-9 4.8e22 2.1 2e36 3 Serino et al. (2015); Nobukawa et al. (2020)
SRGA J144459.2-604207 2022/1 2022/1 4e-10 - - 3.4e36 8.5 Negoro et al. (2024)
SRGA J144459.2-604207 2023/12 2023/12 6e-10 - - 5.2e36 8.5 Negoro et al. (2024); Sguera & Sidoli (2024)
SRGA J144459.2-604207 2024/2 2024/3 2.7e-9 2.9e22 1.8 2.3e37 8.5 Molkov et al. (2024); Ng et al. (2024)
Cen X-4 1969/07 1969/7 6e-7 - - 1.2e38 1.3 Evans et al. (1970); Kuulkers et al. (2009)
Cen X-4 1979/05 1979/5 9.6e-8 9e20 - 1.9e37 1.3 Kaluzienski et al. (1980); Kuulkers et al. (2009)
MAXI J1621-501 2017/09 2019/01 3.1e-9 6.1e22 2.7 2.6e37 8.4 Gorgone et al. (2019); Chenevez et al. (2018)
MAXI J1647-227 2012/06 2012/06 9.2e-10 4.3e21 1.5 1.1e37 10 Onodera et al. (2014); Kennea et al. (2012)
MXB 1658-298 1978/3 1979/7 1.9e-9 - - 2.3e37 10 Cominsky et al. (1983)
MXB 1658-298 1999/09 1999/09 6.5e-10 - - 7.8e36 10 Heise et al. (1999); Oosterbroek et al. (2001)
MXB 1658-298 2015/09 2015/09 5.5e-10 2.7e21 - 6.6e36 10 Sharma et al. (2018)
IGR J16597-3704 2017/10 2017/10 2.2e-10 1.5e22 1.46 2.2e36 9.1 Sanna et al. (2018)
XTE J1701-407 2008/6 2011/7 7.3e-10 3.4e22 2.0 3.3e36 6.2 Falanga et al. (2009); Linares et al. (2009)
XTE J1701-407 2011/09 2011/09 3.8e-10 3e22 1.4 1.8e36 6.2 Degenaar et al. (2011a)
XTE J1701-407 2016/08 2017/4 3.5e-10 - - 1.6e36 6.2 Swift/XRT
XTE J1701-407 2018/03 2018/03 3.5e-10 - - 1.6e36 6.2 Barthelmy et al. (2018),Swift/XRT
XTE J1701-407 2019/04 2019/04 4e-10 3.2e22 1.9 1.8e36 6.2 Iwakiri et al. (2019)
XTE J1701-407 2020/01 2020/03 1.8e-10 - - 8.2e35 6.2 Onori et al. (2020); Palmer et al. (2020)
XTE J1701-462 2005/12 2007/8 4.3e-8 2.7e22 1.9 4e38 8.8 Lin et al. (2009b); Kennea et al. (2006)
XTE J1701-462 2022/9 2023/3 1.4e-8 3e22 - 1.3e38 8.8 Cocchi et al. (2023),Swift/XRT
IGR J17062-6143 2005/12 2022/9* 1.4e-10 2.3e21 - 8.7e35 7.3 Remillard & Levine (2008b); Bult et al. (2021a)
XTE J1709-267 1990/3 1990/8 4.3e-9 - - 3.7e37 8.5 Yamauchi (2010)
XTE J1709-267 1997/1 1997/4 6.7e-9 - - 5.8e37 8.5 Cocchi et al. (1998)
XTE J1709-267 2002/3 2002/4 7.3e-9 4.4e21 - 6.3e37 8.5 Jonker et al. (2003)
XTE J1709-267 2004/5 2004/5 7.7e-10 - - 6.6e36 8.5 Markwardt et al. (2004)
XTE J1709-267 2005/5 2005/5 2e-10 - - 1.7e36 8.5 RXTE PCA Scan
XTE J1709-267 2007/11 2008/9 1.9e-9 - - 1.6e37 8.5 Remillard (2007), RXTE PCA Scan
XTE J1709-267 2009/3 2009/5 2.5e-10 - - 2.2e36 8.5 RXTE PCA Scan
XTE J1709-267 2010/7 2010/7 1.9e-9 - - 1.6e37 8.5 Negoro et al. (2010a)
XTE J1709-267 2012/8 2012/8 3.7e-9 - - 3.2e37 8.5 Sanchez-Fernandez et al. (2012)
XTE J1709-267 2013/8 2013/8 1e-9 - - 8.9e36 8.5 Negoro et al. (2013a)
XTE J1709-267 2016/6 2016/6 4.5e-9 - - 3.9e37 8.5 Nakahira et al. (2016)
XTE J1709-267 2017/8 2017/8 3.9e-10 - - 3.4e36 8.5 Negoro et al. (2017)
XTE J1709-267 2018/12 2019/1 2.2e-9 - - 1.9e37 8.5 MAXI
XTE J1709-267 2020/9 2020/12 2.2e-9 - - 1.9e37 8.5 Adachi et al. (2020)
XTE J1709-267 2022/6 2022/6 1.3e-9 - - 1.1e37 8.5 MAXI
XTE J1709-267 2023/10 2023/11 1.4e-9 - - 1.2e37 8.5 Negoro et al. (2023),MAXI
2S 1711-339 1975/2 1976/9 3.2e-9 1.5e22 2.2 2.2e37 $<$7.5 Cornelisse et al. (2002b)
2S 1711-339 1998/7 1999/5 1.1e-9 - - 7.2e36 $<$7.5 Remillard (1998)
2S 1711-339 2004/3 2004/3 4.5e-10 - - 3.0e36 $<$7.5 Markwardt & Swank (2004a)
2S 1711-339 2010/2 2010/2 4.3e-10 - - 2.9e36 $<$7.5 Negoro et al. (2010b)
2S 1711-339 2016/11 2016/11 1.1e-9 - - 7.7e36 $<$7.5 Negoro et al. (2016b)
1M 1716-315 1971/9 1984/6 1.1e-9 2.1e21 2 6.2e36 6.9 Jonker et al. (2007); Makishima et al. (1981)
IGR J17191-2821 2007/3 2007/5 1.8e-9 3.4e21 1.75 2.7e37 $<$11 Altamirano et al. (2010c); Markwardt et al. (2007)
IGR J17191-2821 2012/6 2012/6 9.3e-10 - - 1.3e37 $<$11 Sugimoto et al. (2012a)
XTE J1723-376 1999/2 1999/3 1.9e-9 4e22 - 1.4e37 8 Marshall & Markwardt (1999); Marshall et al. (1999)
4U 1730-22 1972/8 1972/8 2.9e-9 - - 1.7e37 6.9 Chen et al. (1997)
4U 1730-22 2021/6 2023/8* 2.4e-9 3.9e21 - 1.4e37 6.9 Chen et al. (2023); Bult et al. (2022a)
KS 1731-260 1988/10 2001/2 5.2e-9 1.1e22 - 3.0e37 7 Wijnands et al. (2001b)
XB 1733-30 1980/7 1999/4 3.3e-9 1.8e22 2.1 1.3e37 5.7 Guainazzi et al. (1999); Baumgardt & Vasiliev (2021)
XB 1733-30 2021/4 2023/7* 2e-9 2.2e22 2 7.7e36 5.7 Homan et al. (2023)
Swift J1734.5-3027 2013/9 2013/9 3.6e-10 9.4e21 2.36 2.2e36 7.2 Bozzo et al. (2015)
IGR J17379-3747 2004/2 2004/2 1.3e-10 - - 1.1e36 8.5 Markwardt et al. (2008); Chelovekov & Grebenev (2010)
IGR J17379-3747 2005/3 2005/3 2.2e-10 - - 1.9e36 8.5 Bult et al. (2019a)
IGR J17379-3747 2008/9 2008/9 2.7e-10 - - 2.3e36 8.5 Markwardt et al. (2008); Krimm et al. (2008)
IGR J17379-3747 2014/2 2014/2 2.2e-10 - - 1.9e36 8.5 Negoro et al. (2018a); Bult et al. (2019a)
IGR J17379-3747 2018/3 2018/3 8.0e-10 9e21 1.88 6.9e36 8.5 Negoro et al. (2018a); Bult et al. (2019a)
IGR J17379-3747 2020/9 2020/9 9.2e-11 - 2.1 8e35 8.5 Mereminskiy et al. (2020)
XTE J1739-285 1999/9 1999/9 5.6e-9 1.7e22 - 3.6e37 $<$7.3 Markwardt et al. (1999); Bult et al. (2021b)
XTE J1739-285 2001/5 2001/5 1.0e-9 1.7e22 - 6.4e36 $<$7.3 Kaaret et al. (2007),RXTE/PCA
XTE J1739-285 2003/10 2003/10 1.0e-9 1.7e22 - 6.4e36 $<$7.3 Kaaret et al. (2007),RXTE/PCA
XTE J1739-285 2005/8 2006/3 2.7e-9 1.7e22 - 1.7e37 $<$7.3 Bodaghee et al. (2005)
XTE J1739-285 2012/8 2013/3 1.9e-9 1.7e22 - 1.2e37 $<$7.3 Sanchez-Fernandez et al. (2012); Kuulkers et al. (2013a)
XTE J1739-285 2014/1 2014/12 4.8e-9 1.7e22 - 3.0e37 $<$7.3 MAXI
XTE J1739-285 2019/9 2020/9 4.5e-9 1.7e22 - 2.9e37 $<$7.3 Negoro (2020); Beri et al. (2023), MAXI
XTE J1739-285 2021/5 2021/8 2e-9 1.7e22 2 1.3e37 $<$7.3 Swift/XRT, MAXI
XTE J1739-285 2022/12 2023/2 5.4e-10 1.7e22 - 3.4e36 $<$7.3 Kashyap et al. (2023),MAXI
GRS 1741.9-2853 1990/3 1990/4 1.5e-10 1.5e36 9.0 Sunyaev (1990)
GRS 1741.9-2853 1994/9 1994/9 3.2e-12 1.1e23 2 3.1e34 9.0 Sakano et al. (2002)
GRS 1741.9-2853 1996/9 1996/9 1.8e-10 1.1e23 2.36 1.7e36 9.0 Sakano et al. (2002)
GRS 1741.9-2853 2000/10 2000/10 4.0e-10 9.7e22 1.8 3.9e36 9.0 Muno et al. (2003)
GRS 1741.9-2853 2002/10 2002/10 1.2e-11 1.1e23 1.2e36 9.0 Trap et al. (2009)
GRS 1741.9-2853 2005/4 2005/6 1.1e-10 - - 1.1e36 9.0 Wijnands et al. (2006); Trap et al. (2009)
GRS 1741.9-2853 2006/9 2006/9 1.2e-11 - - 1.2e35 9.0 Degenaar & Wijnands (2009, 2010)
GRS 1741.9-2853 2007/3 2007/4 2.6e-10 - - 2.5e36 9.0 Degenaar & Wijnands (2009, 2010); Trap et al. (2009)
GRS 1741.9-2853 2009/10 2009/10 2.2e-9 - - 2.1e37 9.0 Degenaar & Wijnands (2010)
GRS 1741.9-2853 2010/7 2010/7 2.6e-10 - - 2.5e36 9.0 Degenaar et al. (2014c)
GRS 1741.9-2853 2013/8 2013/8 4.3e-9 - - 4.2e37 9.0 Degenaar et al. (2014c)
GRS 1741.9-2853 2016/3 2016/3 8.7e-11 - - 8.4e35 9.0 Degenaar et al. (2016b)
GRS 1741.9-2853 2017/10 2017/10 1.5e-9 - - 1.5e37 9.0 Degenaar et al. (2017d)
GRS 1741.9-2853 2020/4 2020/10 1.9e-10 - - 1.8e36 9.0 Pike et al. (2021c)
KS 1741-293 1989/8 1989/8 1.1e-9 - - 5.0e36 $<$6.2 in’t Zand et al. (1991)
KS 1741-293 1998/03 1998/9 1.1e-9 - - 5.0e36 $<$6.2 in ’t Zand et al. (1998b)
KS 1741-293 2003/3 2003/3 3.3e-10 - - 1.5e36 $<$6.2 De Cesare et al. (2007)
KS 1741-293 2004/3 2004/3 3.2e-10 3.5e23 2 1.5e36 $<$6.2 De Cesare et al. (2007)
KS 1741-293 2005/7 2005/9 2.5e-10 1.7e23 1.8 1.2e36 $<$6.2 Kuulkers et al. (2007)
KS 1741-293 2007/9 2007/9 3.4e-11 1.7e23 1.8 1.6e35 $<$6.2 Degenaar et al. (2012b)
KS 1741-293 2008/5 2008/8 4.9e-10 1.7e23 1.8 2.3e36 $<$6.2 Degenaar et al. (2012b)
KS 1741-293 2010/2 2010/3 3.4e-10 - - 1.6e36 $<$6.2 Chenevez et al. (2010a)
KS 1741-293 2011/9 2012/2 1.6e-8 - - 7.3e37 $<$6.2 Chenevez et al. (2011a, 2012)
KS 1741-293 2013/2 2013/4 5.7e-10 1.7e23 1.8 2.7e36 $<$6.2 Kuulkers et al. (2013b), Swift/XRT
KS 1741-293 2016/8 2016/8 3.4e-10 - - 1.6e36 $<$6.2 Chenevez et al. (2016)
KS 1741-293 2017/5 2017/5 1.3e-09 1.7e23 1.8 6.0e36 $<$6.2 Bahramian et al. (2021)
KS 1741-293 2019/5 2019/5 7.5e-11 - - 3.4e35 $<$6.2 Heinke et al. (2019a)
1A 1744-361 1976/2 1976/4 1.1e-09 - - 8.3e36 8 Carpenter et al. (1977)
1A 1744-361 1989/08 1989/08 1.3e-09 - - 9.9e36 8 in’t Zand (2004)
1A 1744-361 2003/11 2003/11 2.6e-09 - - 2.0e37 8 Remillard et al. (2003)
1A 1744-361 2004/04 2004/04 3.4e-10 - - 2.6e36 8 Markwardt & Swank (2004b)
1A 1744-361 2005/07 2005/07 5.1e-10 - - 3.9e36 8 Swank & Markwardt (2005)
1A 1744-361 2008/06 2008/6 2.2e-09 - - 1.7e37 8 Remillard & Levine (2008a)
1A 1744-361 2009/11 2009/11 1.4e-09 - - 1.1e37 8 Remillard (2009)
1A 1744-361 2010/03 2010/3 5.8e-10 - - 4.4e36 8 RXTE/PCA
1A 1744-361 2011/09 2011/9 1.7e-09 - - 1.3e37 8 RXTE/PCA
1A 1744-361 2013/08 2013/8 2.1e-09 5.1e21 1.5 1.6e37 8 Bahramian et al. (2013b)
1A 1744-361 2022/5 2022/8 2.4e-09 - - 1.9e37 8 Hughes et al. (2022); Ng et al. (2023)
IGR J17445-2747 2004/2 2004/3 7.4e-11 - 2.2 5.7e35 8 Mereminskiy et al. (2017)
IGR J17445-2747 2017/04 2017/4 9.2e-11 4.9e22 2.2 7.1e35 8 Heinke et al. (2017); Mereminskiy et al. (2017)
IGR J17445-2747 2019/05 2019/5 4.6e-12 - 2.2 3.6e34 8 Heinke et al. (2019a)
IGR J17445-2747 2023/2 2023/3 1.5e-11 5.6e22 2 1.1e35 8 Kashyap et al. (2023)
XMM J174457-2850.3 2001/9 2001/9 6.2e-12 5.9e22 1.9 3.1e34 6.5 Sakano et al. (2002)
XMM J174457-2850.3 2005/6 2005/6 1.0e-10 - - 5.2e35 6.5 Wijnands et al. (2006)
XMM J174457-2850.3 2008/6 2008/7 2.5e-10 1.1e23 1.43 1.3e36 6.5 Degenaar & Wijnands (2010); Degenaar et al. (2014d)
XMM J174457-2850.3 2009/9 2009/10 2.2e-11 7.5e22 2.3 1.1e35 6.5 Degenaar & Wijnands (2010); Degenaar et al. (2014d)
XMM J174457-2850.3 2010/7 2010/8 3.8e-11 7.5e22 2.0 1.9e35 6.5 Degenaar et al. (2014d, 2015)
XMM J174457-2850.3 2012/08 2012/08 1.9e-10 - - 9.8e35 6.5 Degenaar et al. (2014d, 2015)
XMM J174457-2850.3 2016/9 2016/10 3.3e-11 1.1e23 1.6 1.7e35 6.5 Degenaar et al. (2016a)
AX J1745.6-2901 1993/10 1994/10 1.9e-10 2.8e23 2.4 1.5e36 8 Maeda et al. (1996)
AX J1745.6-2901 1997/3 1997/3 2.9e-11 1.3e23 1.44 2.2e35 8 Sakano et al. (2002)
AX J1745.6-2901 2006/2 2006/6 1.2e-10 2.4e23 2.4 9.2e35 8 Degenaar & Wijnands (2010)
AX J1745.6-2901 2007/2 2008/9 8e-10 2.4e23 2.7 6.1e36 8 Degenaar & Wijnands (2010)
AX J1745.6-2901 2010/6 2010/11 8e-11 2.3e23 3.4 6.1e35 8 Degenaar et al. (2010)
AX J1745.6-2901 2013/7 2016/6 3.3e-10 2e23 2.6 6.7e36 8 Degenaar et al. (2014c)
AX J1745.6-2901 2017/4 2017/4 1e-11 2e23 2 8e34 8 Degenaar et al. (2017c)
AX J1745.6-2901 2017/10 2017/10 5.1e-11 3.4e23 1.9 4e35 8 Degenaar et al. (2017e)
AX J1745.6-2901 2019/9 2020/2 3.6e-11 2e23 3.3 3e35 8 Degenaar et al. (2019a, 2020)
AX J1745.6-2901 2021/7 2021/7 1.1e-10 2.2e23 3.5 8.4e35 8 Degenaar et al. (2021)
AX J1745.6-2901 2023/6 2023/11* 3.4e-11 2.2e23 1.0 2.6e35 8 Reynolds et al. (2023)
GRO J1744-28 1995/12 1996/5 3.5e-8 3.5e22 1.2 6.7e37 4 Giles et al. (1996)
GRO J1744-28 1996/12 1997/4 2.1e-8 - - 4.0e37 4 Woods et al. (1999)
GRO J1744-28 2014/1 2014/1 9.7e-9 8.6e22 0.99 1.9e37 4 Degenaar et al. (2014a)
GRO J1744-28 2017/2 2017/2 4.4e-10 7e22 1.2 8.4e35 4 König et al. (2020); Sanna et al. (2017a)
GRO J1744-28 2021/6 2021/ 1.2e-9 9e22 1.1 2.3e36 4 Mereminskiy et al. (2021)
EXO 1745-248 2000/7 2000/8 1.4e-8 - - 6e37 5.9 Heinke et al. (2003a)
EXO 1745-248 2011/10 2011/11 2.1e-9 1.2e22 2.25 8.8e36 5.9 Serino et al. (2012)
EXO 1745-248 2015/3 2015/6 2.0e-8 3.0e22 0.93 8.4e37 5.9 Tetarenko et al. (2016a)
IGR J17480-2446 2010/10 2010/11 1.4e-8 1.2e22 2.1 5.1e37 5.9 Patruno et al. (2012); Miller et al. (2011)
Swift J174805.3-244637 2012/7 2012/8 1.1e-8 - 1.6 4.6e37 5.9 Bahramian et al. (2014)
Swift J174805.3-244637 2023/2 2023/3 2.5e-9 $2$ 1.0e37 5.9 Heinke et al. (2023); Sanna et al. (2023a)
IGR J17473-2721 2005/5 2005/5 1.2e-9 - - 4.3e36 5.5 Markwardt & Swank (2005a)
IGR J17473-2721 2008/3 2008/8 1.3e-8 3.8e22 - 4.7e37 5.5 Altamirano et al. (2008b); Chenevez et al. (2011b)
XMMU J174716.1-281048 2003/3 2012/10 7.4e-12 9e22 2.25 6.3e34 $<$8.4 Degenaar & Wijnands (2007); Kaur et al. (2017)
XMMU J174716.1-281048 2021/4 2021/4 5.9e-12 - - 5e34 $<$8.4 Swift/XRT
EXO 1747-214 1984/4 1985/4 1.6e-9 1.1e22 2.4 2.4e37 $<$11 Parmar et al. (1985); Tomsick et al. (2005)
SAX J1747.0-2853 1976/6 1976/6 7.9e-10 - - 5.3e36 7.5 Proctor et al. (1978); Grebenev et al. (2002)
SAX J1747.0-2853 1991/10 1991/10 4.2e-10 7e22 2.4 2.8e36 7.5 Grebenev et al. (2002)
SAX J1747.0-2853 1998/3 1998/3 5e-10 7e22 2.4 3.4e36 7.5 in ’t Zand et al. (1998a); Sidoli et al. (1998)
SAX J1747.0-2853 1999/3 1999/3 3.4e-9 8.8e22 2.4 2.3e37 7.5 Werner et al. (2004)
SAX J1747.0-2853 2000/2 2000/9 1.1e-8 7e22 2 7.4e37 7.5 Wijnands et al. (2002b); Werner et al. (2004)
SAX J1747.0-2853 2001/9 2001/9 5.4e-9 7e22 2 3.6e37 7.5 Wijnands et al. (2002b); Werner et al. (2004)
SAX J1747.0-2853 2004/3 2004/3 6.4e-9 - 4.3e37 7.5 Markwardt & Swank (2004a)
SAX J1747.0-2853 2005/10 2006/9 3.9e-10 9.5e22 2 2.6e36 7.5 Degenaar et al. (2012b)
SAX J1747.0-2853 2007/10 2007/10 2.2e-9 - - 1.5e37 7.5 Brandt et al. (2007)
SAX J1747.0-2853 2009/3 2009/3 3.4e-11 - - 2.3e35 7.5 Campana et al. (2009)
SAX J1747.0-2853 2011/2 2011/2 6e-9 - - 4e37 7.5 Kennea et al. (2011),Swift/XRT
SAX J1747.0-2853 2012/2 2012/2 7.1e-9 - - 4.8e37 7.5 Chenevez et al. (2012)
SAX J1747.0-2853 2013/4 2013/4 3.5e-9 - - 2.4e37 7.5 Negoro et al. (2013b)
SAX J1747.0-2853 2014/10 2014/10 4.1e-10 - - 2.7e36 7.5 Chenevez et al. (2014)
SAX J1747.0-2853 2016/6 2016/6 4.8e-9 - - 3.2e37 7.5 Clavel et al. (2016)
SAX J1747.0-2853 2017/4 2017/4 5.4e-10 - - 4e36 7.5 Heinke et al. (2017)
SAX J1747.0-2853 2019/2 2019/2 1.3e-8 - - 9e37 7.5 Negoro et al. (2019); Mereminskiy et al. (2019)
SAX J1747.0-2853 2022/4 2022/5 7e-9 - - 5e37 7.5 Iwakiri et al. (2022); Kennea et al. (2022b)
SAX J1747.0-2853 2023/5 2023/6* 1.4e-9 - - 9.4e36 7.5 Li et al. (2023),MAXI,Swift/XRT
SAX J1748.9-2021? 1971/12 1972/1 7.3e-9 5.3e22 2.1 6.3e37 8.5 Markert et al. (1975); Forman et al. (1976)
SAX J1748.9-2021 1998/8 1998/8 1.3e-9 8.2e21 1.44 1.1e37 8.5 in ’t Zand et al. (1999)
SAX J1748.9-2021 2001/8 2001/11 4.3e-9 - - 3.7e37 8.5 in’t Zand et al. (2001)
SAX J1748.9-2021 2005/5 2005/7 5.4e-9 - - 4.7e37 8.5 Markwardt & Swank (2005b); Altamirano et al. (2008a)
SAX J1748.9-2021 2009/12 2010/1 5.4e-9 - - 4.7e37 8.5 Suzuki et al. (2009); Patruno et al. (2010)
SAX J1748.9-2021 2015/2 2015/4 4e-9 5.7e21 2.3 3.4e37 8.5 Pintore et al. (2016)
SAX J1748.9-2021 2017/10 2017/10 1.9e-10 5.1e21 1.6 1.6e36 8.5 Pintore et al. (2018)
SAX J1748.9-2021 2021/11 2021/12 4.9e-10 6e22 1.8 4.2e36 8.5 Pike et al. (2021a, b)
IGR J17494-3030 2012/3 2012/3 1.0e-10 1.8e22 1.9 7.7e35 8 Armas Padilla et al. (2013)
IGR J17494-3030 2020/10 2020/10 1.4e-10 - - 1.0e36 8 Ducci et al. (2020); Ng et al. (2021b)
Swift J1749.4-2807 2006/6 2006/6 9.5e-11 3e22 5e35 6.7 Wijnands et al. (2009)
Swift J1749.4-2807 2010/4 2010/4 5.2e-10 - - 2.8e36 6.7 Ferrigno et al. (2011); Altamirano et al. (2011b)
Swift J1749.4-2807 2021/3 2021/3 3.1e-10 4.5e22 - 1.7e36 6.7 Marino et al. (2022)
IGR J17498-2921 2011/8 2011/9 5.1e-10 1.2e22 1.8 3.5e36 7.6 Falanga et al. (2012); Linares et al. (2011)
IGR J17498-2921 2023/4 2023/5 6.1e-10 1.2e22 1.8 4.2e36 7.6 Sanna et al. (2023b), NICER
SAX J1750.8-2900 1997/3 1997/3 3.5e-9 6e22 2.4 1.9e37 6.79 Natalucci et al. (1999)
SAX J1750.8-2900 2001/3 2001/3 3.5e-9 - - 4.2e36 6.79 Kaaret et al. (2002)
SAX J1750.8-2900 2008/3 2008/10 5.1e-9 2.3e22 1.2 2.8e37 6.79 Markwardt & Swank (2008); Linares et al. (2008)
SAX J1750.8-2900 2011/2 2011/2 3.6e-10 3.2e22 2 2.0e36 6.79 Natalucci et al. (2011)
SAX J1750.8-2900 2015/9 2016/3 2.4e-9 2.3e22 2 1.3e37 6.79 Sanchez-Fernandez et al. (2015), IGR/JEM-X
SAX J1750.8-2900 2018/9 2018/9 4.6e-10 - - 2.5e36 6.79 Bozzo et al. (2018)
SAX J1750.8-2900 2019/4 2019/6 3.8e-10 3e22 2 2.1e36 6.79 Heinke et al. (2019b, a)
XTE J1751-305 1998/6 1998/6 1.7e-9 - 1.8 1.5e37 8.5 Markwardt et al. (2002)
XTE J1751-305 2002/4 2002/4 1.1e-9 - 1.8 9.9e36 8.5 Markwardt et al. (2002)
XTE J1751-305 2005/3 2005/3 1.8e-10 - - 1.6e36 8.5 Swank et al. (2005)
XTE J1751-305 2007/4 2007/4 2.2e-10 - - 1.9e36 8.5 Markwardt & Swank (2007)
XTE J1751-305 2009/10 2009/10 3.2e-10 9.8e21 - 2.8e36 8.5 Chenevez et al. (2009); Miller et al. (2003b)
IGR J17511-3057 2009/9 2009/9 4.5e-10 5e21 1.2 2.6e36 6.9 Bozzo et al. (2010)
IGR J17511-3057 2015/3 2015/4 5e-10 1.2e22 1.5 2.8e36 6.9 Papitto et al. (2016)
SAX J1753.5-2349? 2006/4 2006/4 2.4e-10 - - 1.8e36 8 Turler et al. (2006)
SAX J1753.5-2349 2008/10 2008/10 1.7e-10 1.9e22 2.1 1.3e36 8 Del Santo et al. (2010)
SAX J1753.5-2349 2010/3 2010/3 2.8e-10 - 2.1 2.2e36 8 Chenevez et al. (2010b)
AX J1754.2-2754 1999/10 2008/5 1.1e-11 2.1e22 2.54 5.7e34 6.6 Sakano et al. (2002); Bassa et al. (2008)
AX J1754.2-2754 2008/10 2012/5 1.1e-11 2.3e22 2.9 5.7e34 6.6 Jonker & Keek (2008)
AX J1754.2-2754 2014/2 2015/3 3e-12 2.1e22 2.5 1.6e34 6.6 Swift/XRT
AX J1754.2-2754 2015/8 2015/8 4.6e-12 2.1e22 2.5 2.4e34 6.6 Swift/XRT
AX J1754.2-2754 2015/10 2017/2 1.3e-11 2.1e22 2.5 7e34 6.6 Swift/XRT
Swift J1756.9-2508 2007/6 2007/6 3.7e-10 6.2e22 2.2 2.8e36 8 Krimm et al. (2007b, a)
Swift J1756.9-2508 2009/8 2009/8 5.7e-10 - - 4.4e36 8 Patruno et al. (2009)
Swift J1756.9-2508 2018/4 2018/4 2.5e-10 4.3e22 - 1.9e36 8 Rai & Paul (2019)
Swift J1756.9-2508 2019/6 2019/6 2.6e-10 4.2e22 - 2.0e36 8 Li et al. (2021)
IGR J17591-2342 2018/8 2018/10 3.2e-10 4.4e22 2.4 2.5e36 8 Gusinskaia et al. (2020)
2S 1803-245 1976/5 1976/5 2.2e-8 1.47e22 - 1.4e38 $<$7.3 Cornelisse et al. (2007)
2S 1803-245 1998/4 1998/4 9.1e-9 - - 5.8e37 $<$7.3 Marshall et al. (1998)
1RXS J180408.9-342058 1990 1990 2.0e-12 - 2.3 2.4e34 10.0 Voges et al. (1999)
1RXS J180408.9-342058 2015/1 2015/6 2.5e-9 3.3e21 - 3.0e37 10.0 Degenaar et al. (2016c); Parikh et al. (2017b)
SAX J1806.5-2215 1996/3 1997/10 2.5e-10 - - 1.9e36 $<$8 Cornelisse et al. (2002a)
SAX J1806.5-2215 2011/2 2017/7 1.5e-9 4.2e22 1.7 1.1e37 $<$8 Del Santo et al. (2012); Shaw et al. (2017), Swift/XRT
MAXI J1807+132 2017/3 2017/3 2.2e-10 - - 4e36 $<$12.4 Shidatsu et al. (2017)
MAXI J1807+132 2019/9 2019/9 1.9e-10 1.3e21 2 3.5e36 $<$12.4 Albayati et al. (2021)
MAXI J1807+132 2023/7 2023/10 2.7e-10 1.3e21 2 4.9e36 $<$12.4 Illiano et al. (2023a), Swift/XRT
SAX J1808.4-3658 1996/9 1996/10 2.2e-9 - - 3.2e36 3.5 in ’t Zand et al. (1998c); Revnivtsev (2003)
SAX J1808.4-3658 1998/4 1998/4 1.3e-9 - - 1.9e36 3.5 Chakrabarty et al. (1998); Bult & van der Klis (2015)
SAX J1808.4-3658 2000/1 2000/4 2.8e-10 1.3e21 2.16 4.1e35 3.5 Wijnands et al. (2001a)
SAX J1808.4-3658 2002/10 2002/11 1.8e-9 - - 2.7e36 3.5 Bult & van der Klis (2015)
SAX J1808.4-3658 2005/5 2005/7 1.3e-9 - - 1.9e36 3.5 Bult & van der Klis (2015)
SAX J1808.4-3658 2008/9 2008/11 1.2e-9 - - 1.7e36 3.5 Bult & van der Klis (2015)
SAX J1808.4-3658 2011/11 2011/11 1.7e-9 - - 2.5e36 3.5 Bult & van der Klis (2015)
SAX J1808.4-3658 2015/4 2015/4 1.2e-9 1e21 1.6 1.7e36 3.5 Sanna et al. (2015)
SAX J1808.4-3658 2019/8 2019/8 2.1e-10 2.1e21 2 3.1e35 3.5 Bult et al. (2019b)
SAX J1808.4-3658 2022/8 2022/10 3.7e-10 - - 5.4e35 3.5 Illiano et al. (2023b)
XTE J1807-294 2003/2 2003/6 1.3e-9 5.6e21 - 9.6e36 8 Markwardt et al. (2003); Falanga et al. (2005)
XTE J1810-189 2008/3 2008/6 1.1e-9 3.8e22 1.7 4.0e36 5.5 Chelovekov & Grebenev (2009); Weng et al. (2015)
XTE J1810-189 2009/1 2009/3 1.9e-10 3.8e22 1.7 6.9e35 5.5 Chelovekov & Grebenev (2009),RXTE/PCA
XTE J1810-189 2011/6 2011/6 2.9e-9 4.9e22 2.6 1.0e37 5.5 Rangelov et al. (2014)
XTE J1810-189 2012/12 2013/1 8.6e-10 - - 3.1e36 5.5 Negoro et al. (2013c)
XTE J1810-189 2020/8 2020/11 5.2e-10 6e22 1.8 1.9e36 5.5 Bozzo et al. (2020); Manca et al. (2023)
XTE J1810-189 2023/2 2023/10 1e-9 - - 3.6e36 5.5 Kobayashi et al. (2023), MAXI
SAX J1810.8-2609 1998/3 1998/3 3.1e-10 - - 1.2e36 $<$5.7 Natalucci et al. (2000)
SAX J1810.8-2609 2002/4 2002/4 1.0e-10 5e21 2 3.9e35 $<$5.7 RXTE/PCA,RXTE/ASM
SAX J1810.8-2609 2007/8 2007/10 8e-10 5e21 2 3e36 $<$5.7 Degenaar et al. (2007); Degenaar & Wijnands (2013a)
SAX J1810.8-2609 2012/5 2012/5 5.2e-9 - - 2.0e37 $<$5.7 Degenaar & Wijnands (2013a)
SAX J1810.8-2609 2018/4 2018/4 1.8e-9 - - 7e36 $<$5.7 Negoro et al. (2018b)
SAX J1810.8-2609 2021/5 2021/10 2.1e-9 - - 8.2e36 $<$5.7 Iwakiri et al. (2021); Hughes et al. (2023)
XTE J1812-182 2003/1 2003/3 1.9e-9 1.3e23 2.47 4.5e37 14 Cackett et al. (2006)
XTE J1812-182 2008/8 2008/10 2.2e-9 - - 5.1e37 14 Goodwin et al. (2019)
XTE J1812-182 2020/3 2020/3 2.0e-9 - - 4.7e37 14 Chenevez et al. (2020)
XTE J1814-338 1984/9 1984/9 5e-11 - - 3.8e35 8 Wijnands & Reynolds (2003)
XTE J1814-338 2003/6 2003/7 3.7e-10 1.6e21 2.1 2.8e36 8 Papitto et al. (2007); Krauss et al. (2005)
MAXI J1816-195 2022/6 2022/7 3.2e-9 2.4e22 2.6 1.5e37 $<$6.3 Kennea et al. (2022a); Bult et al. (2022b)
Swift J181723.1-164300 2017/8 2017/8 3.6e-10 8e22 1.5 2.1e36 7 Parikh et al. (2017a)
IGR J18245-2452 2013/3 2013/4 1.2e-9 4e21 1.26 4.3e36 5.5 Papitto et al. (2013); Linares et al. (2014)
SAX J1828.5-1037 1993/4 1993/4 7.2e-12 4.1e22 1.7 3.3e34 $<$6.2 Cornelisse et al. (2002b)
SAX J1828.5-1037 2001/3 2001/3 7.5e-12 5e22 - 6e34 $<$6.2 Hands et al. (2004)
SAX J1828.5-1037 2008/10 2008/11 1.5e-11 4.1e22 1.6 6.9e34 $<$6.2 Campana (2009); Degenaar et al. (2008), Swift/XRT
SAX J1828.5-1037 2011/11 2011/11 2.9e-10 4.1e22 1.6 1.3e36 $<$6.2 Serino et al. (2016), Swift/XRT
Swift J185003.2-005627 2011/5 2011/7 1.2e-10 1.1e22 1.6 2.0e35 $<$3.7 Degenaar et al. (2012a)
Swift J1858.6-0814 2018/10 2020/5 2.1e-9 1.8e23 - 4.1e37 12.8 Negoro et al. (2020); Hare et al. (2020)
HETE J1900.1-2455 2005/6 2015/10 8e-10 2e21 - 2.1e36 4.7 Degenaar et al. (2017a); Kaaret et al. (2006)
1H 1905+000 1974/11 1985/9 1.5e-10 1.9e21 2.1 1.0e36 7.5 Jonker et al. (2006)
Swift J1922.7-1716 2004/11 2006/7 1.3e-10 2.9e21 2.1 3.6e35 4.8 Falanga et al. (2006); Degenaar et al. (2012a)
Swift J1922.7-1716 2011/7 2011/8 2.1e-10 2.7e21 2.0 5.8e35 4.8 Degenaar et al. (2012a)
MAXI J1957+032 2015/5 2015/5 2.0e-10 1.7e21 2 6.0e35 5 Negoro et al. (2015b); Mata Sánchez et al. (2017)
MAXI J1957+032 2015/10 2015/10 3.2e-10 - - 9.5e35 5 Sugimoto et al. (2015); Mata Sánchez et al. (2017)
MAXI J1957+032 2016/1 2016/1 4e-10 - - 1.2e36 5 Tanaka et al. (2016); Beri et al. (2019)
MAXI J1957+032 2016/9 2016/9 6.6e-10 - - 2.0e36 5 Negoro et al. (2016c); Mata Sánchez et al. (2017)
MAXI J1957+032 2022/6 2022/6 8e-10 - - 2.4e36 5 Negoro et al. (2022); Sanna et al. (2022)
XTE J2123-058 1999/6 1999/8 1.7e-9 6e20 2 1.5e37 8.5 Tomsick et al. (1999, 2004)
4U 2129+47 1970 1983/9 4.8e-10 2e21 - 1.8e35 1.75 Forman et al. (1978); Bozzo et al. (2007)
Note. — Published outbursts of verified neutron star LMXBs. Max $F_{X}$ is highest reported flux in each outburst, transformed to 2-10 keV, generally using the power-law spectral index & $N_{H}$ noted. We assume 1 Crab is Fx(2-10)=2e-8 erg/s/cm2. Peak observed intrinsic $L_{X}$ in 2-10 keV, for the assumed distance (listed). A * at the end date indicates that we do not have evidence that the outburst ended–the last recorded observation found it active.
[…]
Notes on individual sources, except the frequent outbursters (which are in Appendix B):
IGR J00291+5934; the distance estimate is 4.2±$\pm 0.5$ kpc, from a photospheric radius expansion (PRE) burst (De Falco et al., 2017).
[…]
EXO 0748-676; Galloway et al. (2008) derive the distance of 7.4±$\pm 0.9$ kpc from a PRE burst.
GS 0836-429; also known as 4U 0836-429, 1M 0836-425, MX 0836-42. Aranzana et al. (2016) place a 9.2 kpc upper limit on the distance from a burst. Note that we use the $N_{H}$ value of $2.7\times 10^{22}$ cm-2 quoted by Chelovekov et al. (2005), which Aranzana et al. (2016) misquote as $2.7\times 10^{21}$ cm-2.
MAXI J0911-655; also known as Swift J0911.9-6452. The 9.5 kpc distance is the estimate to the host globular cluster, NGC 2808 (Watkins et al., 2015). Its long outburst appears to have ended around October 2023 (Heinke & Degenaar, 2023).
XTE J0929-314; Galloway et al. (2002) estimate a lower limit on the distance ($>6$ kpc) from binary evolution considerations.
SAX J1324.5-6313; the distance limit of $<6.2$ kpc is from a burst (Cornelisse et al., 2002b).
MAXI J1421-613; Nobukawa et al. (2020) estimate the distance as 3 kpc from analysis of dust scattering echoes.
Cen X-4 or 4U 1456-32; Kaluzienski et al. (1980) estimate a distance of 1.3±$\pm 0.7$ kpc from the X-ray burst. We reference $N_{H}=9\times 10^{20}$ from Chakrabarty et al. (2014).
MAXI J1621-501; Chenevez et al. (2018) estimate 8.4±$\pm 2$ kpc from a PRE burst.
MAXI J1647-227; the 10 kpc rough distance estimate was made by Onodera et al. (2014).
MXB 1659-298 or MXB 1658-298; the 10 kpc distance estimate is from a PRE burst (Sharma et al., 2018).
IGR J16597-3704; the 9.1 kpc distance estimate is from the globular cluster NGC 6256 (Valenti et al., 2007).
XTE J1701-407; the distance of 6.2±$\pm 0.5$ kpc was estimated from a PRE burst (Falanga et al., 2009; Linares et al., 2009). It is not clear whether XTE J1701-407 has ever reached full quiescence since 2016.
XTE J1701-462; Lin et al. (2009a) measure a distance of 8.8±$\pm 1.3$ kpc via PRE bursts. Swift/XRT observations in March 2023 showed its 2023 outburst declining below $L_{X}<10^{34}$ erg/s.
IGR J17062-6143; Keek et al. (2017) measure a distance of 7.3±$\pm 0.5$ kpc from a PRE burst. We use $N_{H}=2.3\times 10^{21}$ cm-2 from Degenaar et al. (2017b).
XTE J1709-267, or RX J1709.5-2639; Jonker et al. (2003) suggest a distance of 8.5 kpc if it is associated with the nearby globular cluster NGC 6293.
2S 1711-339 or 4U 1711-34; the distance limit of 7.5 kpc is from a burst (Cornelisse et al., 2002b).
1M 1716-315; also known as MXB 1715-32, XB 1715-321, or 1H 1715-321. The distance of 6.9 kpc is estimated from a PRE burst Jonker & Nelemans (2004).
IGR J17191-2821; the distance limit of 11 kpc is from a burst (Markwardt et al., 2007).
XTE J1723-376; the 8 kpc distance limit is from a burst (Marshall et al., 1999).
4U 1730-22; Bult et al. (2022a) measure the distance of 6.9±$\pm 0.2$ kpc from PRE bursts.
KS 1731-260; Muno et al. (2000) estimate the distance from PRE bursts.
XB 1733-30; this quasi-persistent transient in Terzan 1 is also known as SLX 1732-304, XB 1732-304, and 1RXS J173546.9-302859. We use the globular cluster distance estimate of 5.7±$\pm 0.2$ kpc from Baumgardt & Vasiliev (2021).
Swift J1734.5-3027 or IGR J17344-3023; Bozzo et al. (2015) measured a distance of 7.2±$\pm 1.5$ kpc from a PRE burst.
IGR J17379-3747 or IGR J17380-3749; this is assumed to be located in the Galactic bulge, for a distance of roughly 8.5 kpc Chelovekov & Grebenev (2010).
XTE J1739-285; Galloway et al. (2008) find a distance constraint of $<$7.3 kpc from the brightest bursts, assuming a H-rich composition, which appears appropriate for the brightest bursts given their longer rise and decay times (Galloway et al., 2008).
GRS 1741.9-2853; Pike et al. (2021c) measure a distance of 9.0±$\pm 0.5$ kpc from a 2020 May PRE burst. (This is farther than the 7 kpc distance estimate inferred from prior X-ray bursts, which however did not show PRE; Barrière et al. (2015).) The proximity of this source to the Galactic Centre prevents us from using all-sky monitors to track its recurrences.
KS 1741-293; Chelovekov & Grebenev (2011) place an upper limit of $<$6.2 kpc from analyzing bursts. While this source has generally been seen not to exceed a few $10^{36}$ ergs/s (see e.g. Degenaar & Wijnands 2013b), we notice that the 2011-2012 outburst was reported by Chenevez et al. (2012) using INTEGRAL/JEM-X to reach a 3-10 keV flux of 415±$\pm 75$ mCrab, an order of magnitude brighter. We also identify a Swift/XRT observation on May 14, 2017 that reached 5.2 cts/s, or $L_{X}\sim 6.4*10^{36}$ erg/s.
1A 1744-361, or A 1744-361; its proximity to the Galactic center has led many workers to assume an 8 kpc distance, though the only burst constraint is $<$11 kpc (Galloway et al., 2008). The RXTE/PCA Bulge scans show three likely small outbursts, at MJD 52020, 53281, and 54188, all reaching around $F_{X}=5\times 10^{-11}$ ergs/cm2/s. Given the crowded environment, it would be useful to verify these signals. 1A 1744-361 is not clearly resolved by MAXI, so our sensitivity to its outbursts declines after 2011.
IGR J17445-2747; the proximity of this source to the Galactic Centre suggests an 8 kpc distance. Mereminskiy et al. (2017) place limits of $>$5 and $<$12.3 kpc, from analysis of a burst.
XMM J174457-2850.3; Degenaar et al. (2014c) measure a distance of 6.5±$\pm 0.2$ kpc from a PRE burst. We use the long-term Swift XRT monitoring lightcurve (Degenaar et al., 2015) to check for additional outbursts (with signal/error $>$4, $L_{X}>2\times 10^{34}$ erg/s), and find only one (in 2016, Degenaar et al., 2016a) between 2012 and 2022. This appears to be a substantial change from its frequent outbursting behavior before 2012 (Degenaar et al., 2014d).
AX J1745.6-2901; the proximity to the Galactic Centre suggests an 8 kpc distance, which is supported by the peak fluxes of the brightest X-ray bursts (Degenaar & Wijnands, 2009). (Note this is not the same as 1A 1742-289, Kennea & Skinner 1996.)
GRO J1744-28; although the proximity to the Galactic Centre suggests a distance of 8 kpc, several methods of estimating the distance seem to converge on a value around 4 kpc (Sanna et al., 2017c; Gosling et al., 2007; Wang et al., 2007). This adjusts the $L_{X}$ of the brightening episode in Sept. 2008 to $L_{X}=5\times 10^{33}$ erg/s, below our outburst definition. The 2011 MAXI outburst initially identified as GRO 1744-28 (Suwa et al., 2011) was from SAX J1747.0-2853 (Kennea et al., 2011).
EXO 1745-248 (or XB 1745-248), IGR J17480-2446 (or Ter 5 X-2), Swift J17480.5-244637 (or Ter 5 X-3, or MAXI J1747-249): These are all located in the globular cluster Terzan 5. Additional outbursts have been seen in 1981, 1991 (Johnston et al., 1995), 2002 (Wijnands et al., 2002a), and 2004 (Markwardt & Swank, 2004a), but these cannot be confidently identified with one of the identified transients, which require arcsecond-scale localizations in this crowded cluster. We use the distance of 5.9 kpc from Valenti et al. (2007).
IGR J17473-2721 is also known as XTE J1747-274. The distance estimate is from PRE bursts (Chenevez et al., 2011b).
XMMU J174716.1-281048; also known as IGR J17464-2811. The distance estimate is an upper limit from a burst (Degenaar et al., 2011b). Del Santo et al. (2015) report that XMMU J174716.1-281048 was not detected by Swift in 2014 or 2015, but it was not clear if it had returned to full quiescence. Extensive Swift/XRT observations, and several archival XMM-Newton and Chandra observations, allow an accurate flux history of this source. We utilize the Leicester Swift/XRT lightcurves (Evans et al., 2007), archival Chandra observations from 2000, 2001, 2019, and 2022 (ObsIDs 1036, 2271, 21624, and 24222, the first 2 previously reported by Del Santo et al. 2007), and archival XMM-Newton observations from 2012, 2018, and 2020 (ObsIDs 0694641401, 0802410101, 0862471201), along with XMM flux values from 2003 and 2005 from Del Santo et al. (2007), to construct a longer lightcurve for XMMU J174716.1-281048 (Fig. 6. This reveals evidence that XMMU J174716.1-281048 declined significantly in late 2012 and had entered deep quiescence (2-10 keV $L_{X}<10^{33}$ erg/s) by 2018, most likely earlier. The Swift/XRT data also reveal that XMMU J174716.1-281048 experienced a second outburst in April 2021, but limit the period brighter than $L_{X}=3\times 10^{34}$ erg/s to between 2 weeks and 3 months. XMM and Chandra show that it was in deep quiescence in October 2020 and again by April 2022.
Figure 6: The Swift/XRT, Chandra, and XMM-Newton lightcurve of XMMU J174716.1-281048. Red filled triangles are Swift/XRT detections, black downward arrows are Swift/XRT upper limits. Blue filled pentagons, and downward triangles, are Chandra data, while magenta filled pentagons and downward triangles are from XMM-Newton.
EXO 1747-214; the distance upper limit is from the EXOSAT burst (Tomsick et al., 2005).
SAX J1747.0-2853: we use the distance of 7.5±$\pm 1.3$ kpc from PRE bursts of Werner et al. (2004). We find a much smaller flux ($6\times 10^{-9}$ ergs/cm2/s, 2-10 keV unabsorbed) in Feb. 2011 than Kennea et al. (2011). The MAXI page notes that the apparent flare in MAXI data in early 2014 is from a different source. It is unclear whether SAX J1747.0-2853 has been active since May 2023.
SAX J1748.9-2021 and NGC 6440 X-2: these two accreting millisecond X-ray pulsars are both located in the globular cluster NGC 6440 X-2, at d=8.5 kpc (Harris, 1996). They can be distinguished by their spatial positions (using Chandra, Swift/XRT, or radio interferometric imaging), or by detection of X-ray pulsations. SAX J1748.9-2021 has been verified to be the object in outburst in all bright detected outbursts, except the 1971 outburst, but the similarity of that outburst to others verified to be SAX J1748.9-2021 suggest that it was responsible.
NGC 6440 X-2 showed numerous weak, short X-ray outbursts in 2009 and 2010, often with a 1-month recurrence time (Heinke et al., 2010) [^10] Note that Heinke et al. (2010) quote $N_{H}=0.59\times 10^{21}$ in their Table 2, but actually used $N_{H}=0.59\times 10^{22}$. The RXTE PCA bulge scan reveals at least 7 faint outbursts during 2009-2011, but does not show other similar features except one in 2007, and possibly 4 in 1999. It is not certain whether NGC 6440 X-2 only produced outbursts in 2007 and 2009-2011, or whether we missed others (due perhaps to changes in the outburst peak flux). We list NGC 6440 X-2’s outbursts in a separate table (Table 2), to quote their MJD dates.
[…]
NS XRB Year/Month MJD Peak $F_{X}$ Index $L_{X}$ References
NGC 6440 X-2 2007/4 54205 1.4e-10 2 1.2e36 Heinke et al. (2010), RXTE/PCA
NGC 6440 X-2 2009/3 54910 1.9e-10 2 1.6e36 Heinke et al. (2010), RXTE/PCA
NGC 6440 X-2 2009/5-6 54980 1e-10 2 8.6e35 Heinke et al. (2010), RXTE/PCA
NGC 6440 X-2 2009/7 55040 1.7e-10 2 1.5e36 Heinke et al. (2010), RXTE/PCA
NGC 6440 X-2 2009/8-9 55072 1.6e-10 1.79 1.4e36 Heinke et al. (2010); Altamirano et al. (2010d), RXTE/PCA
NGC 6440 X-2 2009/10 55105 5.6e-11 1.79 4.8e35 Heinke et al. (2010)
NGC 6440 X-2 2009/10-11 55132 8.1e-11 1.8 7e35 Heinke et al. (2010)
NGC 6440 X-2 2010/3 55274 1.5e-10 1.9 1.3e36 Altamirano et al. (2010a), RXTE/PCA
NGC 6440 X-2 2010/6 55359 2e-10 - 1.7e36 Patruno & D’Angelo (2013), RXTE/PCA
NGC 6440 X-2 2010/10 55473 1.8e-10 - 1.6e36 Patruno & D’Angelo (2013), RXTE/PCA
NGC 6440 X-2 2011/1 55584 2e-10 - 1.7e36 Patruno & D’Angelo (2013)
NGC 6440 X-2 2011/3 55639 2.8e-10 - 2.4e36 RXTE/PCA
NGC 6440 X-2 2011/9 55819 2.2e-10 - 1.9e36 RXTE/PCA
NGC 6440 X-2 2011/11 55871 2e-10 - 1.7e36 Patruno & D’Angelo (2013)
Table 2: Outbursts of NGC 6440 X-2. $F_{X}$ in 2-10 keV ergs cm-2 s-1, $L_{X}$ in 2-10 keV ergs s-1. $N_{H}$ is about $6\times 10^{21}$ cm-2. D=8.5 kpc (Harris, 1996). RXTE/PCA indicates detection in RXTE/PCA bulge scans.
[…]
IGR J17494-3030; Armas Padilla et al. (2013) assume a distance of 8 kpc, due to its proximity to the Galactic Center.
Swift J1749.4-2807; Wijnands et al. (2009) calculate a distance of 6.7±$\pm 1.3$ kpc, but this is an upper limit as the burst is not clearly PRE.
IGR J17498-2921; Linares et al. (2011) estimate a 7.6 kpc distance from a PRE burst. NICER observations of the 2023 outburst show a rapid drop to quiescence in May 2023.
SAX J1750.8-2900; Galloway et al. (2008) calculate a distance of 6.79±$\pm 0.14$ kpc from PRE bursts.
XTE J1751-305; Papitto et al. (2008) discuss various estimates for the distance, from 6.7 to 11.6 kpc; they and other authors use a default distance of 8.5 kpc, as it is close to the Galactic Center.
IGR J17511-3057; the distance upper limit is 6.9 kpc, from bursts (Altamirano et al., 2010b).
SAX J1753.5-2349; the distance is assumed to be 8 kpc. We include an INTEGRAL-detected outburst in 2006, designated IGR J17536-2339, which may or may not be this object (and thus give it a ? in the table).
AX J1754.2-2754; The distance estimate of 6.6±$\pm 0.4$ kpc, from an INTEGRAL burst reported by Chenevez et al. (2017), assumes a helium composition for the atmosphere.
After its initial ASCA detection in 1999, one short interval of quiescence has been reported (in 2007/2008, Bassa et al. 2008). However, the Swift/XRT lightcurve presented by Shaw et al. (2017) suggests multiple additional quiescent periods. We use Swift/XRT observations to identify additional quiescent episodes (with $L_{X}$(2-10 keV)$<10^{34}$ erg/s, or Swift/XRT count rates $<$0.04 cts/s), in August 2012, May-June 2015, Sept. 2015. Since Feb. 2017, it has been routinely observed by the Swift Bulge Survey (Bahramian et al., 2021), with only one detection in July 2019 at $L_{X}$(2-10 keV)$=7\times 10^{33}$ erg/s. We used a recent 4.2 ks Chandra observation (May 5, 2022, ObsID 24082; PI G. Ponti) to place a strong upper limit of $L_{X}$(2-10 keV)$<1.0\times 10^{32}$ erg/s on AX J1754.2-2754. We show the Swift/XRT data, the new and previously published Chandra data, and the XMM-Newton point of Armas Padilla et al. (2013) in Fig. 7. We conclude that AX J1754.2-2754 spends much of its time at luminosities at least 10-100 times lower than its peak.
Figure 7: The Swift/XRT, Chandra, and XMM-Newton lightcurve of AX J1754.2-2754. Red filled triangles are Swift/XRT detections, black downward arrows are Swift/XRT upper limits. Blue filled pentagons are Chandra data, while the magenta filled pentagon is from XMM-Newton.
Swift J1756.9-2508; the 8 kpc distance estimate is based on its proximity to the Galactic Center. Li et al. (2021) place a lower limit on the distance of 4 kpc.
IGR J17591-2342; Russell et al. (2018) argue that the $N_{H}$ implies a distance $>$6 kpc; since it is near the Galactic Centre, it is typically taken to be 8 kpc away.
2S 1803-245 is also known as XTE J1806-246, and SAX J1806.8-2435. Cornelisse et al. (2007) place an upper limit of $<$7.3 kpc, from a burst in 1998.
1RXS J180408.9-342058; the 10.0±$\pm 1.4$ kpc distance estimate is from PRE bursts (Marino et al., 2019).
SAX J1806.5-2215 showed a long outburst, which started in Feb. 2011 (Altamirano et al., 2011a). This was reported in Atels until 2015 (Sguera et al., 2015), but Swift/XRT observations indicate it continued until July 2017, and was no longer visible to Swift/XRT by January 2018. The $<$8 kpc upper limit comes from an analysis of bursts by (Cornelisse et al., 2002a).
MAXI J1807+132 has an upper limit on its distance of 12.4 kpc, from bursts (Albayati et al., 2021). Its 2023 outburst experienced two re-flares through September and October, visible in the Swift/XRT lightcurve; while it might rebrighten yet again, it appears to be in quiescence in Nov. 2023.
SAX J1808.4-3658; Galloway & Cumming (2006) estimated a distance of 3.5$\pm$0.1 kpc, using PRE bursts and the expected mass transfer rate from gravitational radiation.
XTE J1807-294; the distance is assumed to be 8 kpc (based on its proximity to the Galactic Centre).
XTE J1810-189; Weng et al. (2015) estimate a distance of 5.5±$\pm 1.7$ kpc from a PRE burst.
SAX J1810.8-2609: the distance was constrained by an X-ray burst to $<$5.7 kpc by Fiocchi et al. (2009). We note several likely weak, single-datapoint outbursts in the RXTE/PCA lightcurve, that are not remarked on in the literature. At MJD 52384 we see a 10-sigma detection; also there is a 5-sigma RXTE/ASM detection at 52373. We include this outburst in our list. At MJD 53207, and 54989, we see 3-sigma possible outbursts, without RXTE/ASM detections; we leave these off our list.
XTE J1812-182: also known as XMMU J181227.8-181234. This may be the same as 1H1812-182, detected by HEAO-1 with $F_{X}$(2-10 keV)$=7\times 10^{-10}$ erg/cm2/s in 1977/78 (Wood et al., 1984), but the errorbox is large. Goodwin et al. (2019) estimated the distance as 14$\pm$2 kpc, from analysis of a train of several (non-PRE) bursts.
XTE J1814-338: Wijnands & Reynolds (2003) argued this is the same as EXMS B1810-337. The distance estimate of 8 kpc from X-ray bursts is from Strohmayer et al. (2003).
MAXI J1816-195; we use the $<$6.3 kpc distance constraint from bursts from Chen et al. (2022).
Swift J181723.1-164300; the 7$\pm$1 kpc distance comes from a burst (Parikh et al., 2017a).
IGR J18245-2452; this accreting, bursting transitional millisecond pulsar is in the globular cluster M28, at a distance of 5.5 kpc (Harris, 1996, 2010 revision). Although no other outbursts have been recorded from this cluster, a burst during quiescence has been seen (see Appendix C for AX J1824.5-2451).
SAX J1828.5-1037; the distance upper limit of $<$6.2 kpc is from a burst (Cornelisse et al., 2002b). The flux quoted for the 2008 outburst by Campana (2009) appears to be in error by a factor of ten.
Swift J185003.2-005627; the distance estimate upper limit of 3.7 kpc is from the single observed burst (Degenaar et al., 2012a).
[…]
HETE J1900.1-2455; Galloway et al. (2008) compute a distance of 4.7$\pm$0.6 kpc from a PRE burst.
1H 1905+000 or 4U 1857+01, or 4U 1905+000; Jonker & Nelemans (2004) suggest either 7.5 or 10 kpc based on a PRE burst detected by Chevalier & Ilovaisky (1990), depending if the burst was H-rich or He-rich. As Chevalier & Ilovaisky (1990) identify a moderately long rise time (2.5 s) and decay time ($\tau$=19 s), we suggest the burst was H-rich, and prefer the shorter 7.5 kpc distance.
Swift J1922.7-1716; the distance estimate of 4.8 kpc is from two PRE bursts (Degenaar et al., 2012a).
MAXI 1957+032, or IGR J19566+0326; the distance estimate of 5$\pm$2 kpc is from the hypothesized optical counterpart (which would require a triple system, Ravi 2017).
XTE J2123-058; the distance estimate of 8.5±$\pm 2.5$ kpc combines several constraints from bursts and studies of the optical counterpart in quiescence (Tomsick et al., 2004).
4U 2129+47 or XB 2129+47; Arnason et al. (2021) show that the Gaia distance to the star generally believed to be the counterpart is only 1.75 kpc. This is an accretion-disk corona source, so the $L_{X}$ estimate may be underestimated.
MAXI reported an outburst in December 2019 at 14$\pm$3 mCrab, or $L_{X}=$1e35 erg/s (Kurihara et al., 2019), but it is unclear that this was a real detection; Swift observed within 2 days and only placed (unabsorbed) flux limits $F_{X}<7\times 10^{-13}$ erg/s/cm2, or $L_{X}<2.6\times 10^{32}$ erg/s. We leave this possible outburst off our list.
[…]
Appendix B Frequently outbursting sources
[…]
NS XRB Year/Month Peak $F_{X}$ $L_{X}$ References
2-10 keV 2-10 keV
4U 1608-52 1970/04 2.0e-8 2.4e37 Lochner & Roussel-Dupre (1994)
4U 1608-52 1970/09 1.7e-8 2.1e37 Lochner & Roussel-Dupre (1994)
4U 1608-52 1971/06 1.5e-8 1.8e37 Lochner & Roussel-Dupre (1994)
4U 1608-52 1971/09 1.7e-8 2.1e37 Lochner & Roussel-Dupre (1994)
4U 1608-52 1975/11 1.2e-8 1.5e37 Kaluzienski et al. (1975)
4U 1608-52 1977/07 2.4e-8 2.9e37 Clark & Li (1977)
4U 1608-52 1979/01 6.5e-9 7.9e36 Kaluzienski & Holt (1979a)
4U 1608-52 1979/04 1.3e-8 1.6e37 Oda et al. (1979)
4U 1608-52 1980/4 $<$2.1e-10 $<$2.6e35 Oda (1980)
4U 1608-52 1983/04 1.6e-8 2.0e37 Mitsuda et al. (1984)
4U 1608-52 1984/07 1.0e-9 1.2e36 Gottwald et al. (1987)
4U 1608-52 1986/03 5.8e-10 7.1e35 Penninx et al. (1989)
4U 1608-52 1991/6-12 9.7e-10 1.2e36 Zhang et al. (1996)
4U 1608-52 1996/3 7.1e-9 8.7e36 Méndez et al. (1998)
4U 1608-52 1998/2-4 1.2e-8 1.5e37 Rutledge (2001)
4U 1608-52 2000/01 4.4e-9 5.4e36 Šimon (2004)
4U 1608-52 2000/12 7.4e-9 9.0e36 Šimon (2004)
4U 1608-52 2001/11 1.5e-8 1.8e37 Rutledge (2001)
4U 1608-52 2002/07 1.7e-8 2.1e37 Šimon (2004); Keek et al. (2008)
4U 1608-52 2003/03 6.3e-9 7.7e36 Miller et al. (2003a)
4U 1608-52 2004/03 5.4e-9 6.6e36 Remillard & Miller (2004)
4U 1608-52 2004/7 2.9e-9 3.5e36 Yan & Yu (2015)
4U 1608-52 2005/03 2e-8 2.4e37 Keek et al. (2008)
4U 1608-52 2005/09 6.1e-9 7.4e36 Asai et al. (2012)
4U 1608-52 2006/07 1.9e-9 2.3e36 Markwardt et al. (2006); Greco et al. (2006)
4U 1608-52 2007/06-2009/10 7.8e-8 9.5e37 Palmer et al. (2007); Russell et al. (2009)
4U 1608-52 2010/02 6.3e-9 7.7e36 Armas Padilla et al. (2017)
4U 1608-52 2011/03 1.7e-8 2.1e37 Suguzaki et al. (2011)
4U 1608-52 2012/10 1.2e-8 1.5e37 Sugimoto et al. (2012b)
4U 1608-52 2013/5 3e-9 3.7e36 Zhang & Yu (2013)
4U 1608-52 2014/01 4.1e-9 5.0e36 Zhang & Yu (2014)
4U 1608-52 2014/10 6.4e-9 7.8e36 Negoro et al. (2014)
4U 1608-52 2016/08 5.5e-9 6.7e36 Ono et al. (2016)
4U 1608-52 2017/06 1.1e-8 1.3e37 Jaisawal et al. (2019)
4U 1608-52 2018/06 8.5e-9 1.0e37 van den Eijnden et al. (2020)
4U 1608-52 2019/03 5.7e-9 7.0e36 MAXI
4U 1608-52 2020/05 1.5e-8 1.8e37 Güver et al. (2021)
4U 1608-52 2021/08 2.8e-9 3.4e36 MAXI
4U 1608-52 2022/07 $>$5.7e-9 $>$7.0e36 MAXI
Table 3: Outbursts of 4U 1608-52. The distance is assumed to be 3.2 kpc. $F_{X}$ in ergs cm-2 s-1, $L_{X}$ in ergs s-1. Spectral fits include $N_{H}=2\times 10^{22}$ cm-2, $\Gamma$=1.7 (Mitsuda et al., 1984) and $N_{H}=1.5\times 10^{22}$, $\Gamma$=1.9 (Gottwald et al., 1987).
MXB 1730-335, aka Rapid Burster, 4U 1730-335, or MXB 1730-33; We use a distance of 8.0$\pm$0.3 kpc, as measured to the host globular cluster, Liller 1 (Baumgardt & Vasiliev, 2021). We include a table for its frequent outbursts (Table B). We note that some of these outbursts may be produced by the second known transient in this cluster, CXOU J173324.6-332321, which was discovered in low-level outburst ($L_{X}$[0.5-10]$=5.5\times 10^{35}$ erg/s) in April 2018 between two larger outbursts (Homan et al., 2018), but does not show Type II bursts as reported from many outbursts (e.g. Bagnoli et al., 2015). We quote MJD values for the highest observed countrate of each outburst. Where only intervals are given (e.g. in 1971), we estimate the midpoint of the interval. For spectral fitting, we assume a 1.6 keV blackbody, with $N_{H}=1.8\times 10^{22}$ cm-2 (Marshall et al., 2001; Mahasena et al., 2003). We use RXTE/PCA Bulge Scan values where available, then RXTE/ASM, MAXI, or Swift/XRT. For RXTE/ASM we multiply countrates by 3.4×$\times 10^{-10}$ to get 2-10 keV fluxes, for RXTE/PCA by 2.3×$\times 10^{-12}$. For MAXI we subtract 0.07 cts/cm-2/s (to account for confusion) from the 2-20 keV fluxes, and multiply by 1.1×$\times 10^{-8}$ to get 2-10 keV fluxes. For Swift/XRT data we multiply by 7.5×$\times 10^{-11}$ to get 2-10 keV fluxes, and for INTEGRAL/JEM-X 3-10 keV data we multiply by $1.7\times 10^{-10}$. These different flux estimates are generally in reasonable agreement where they overlap, though the RXTE/ASM has some untrustworthy points when the Sun is very close. We identify 52 outbursts that we do not see specifically identified elsewhere in the literature.
Masetti (2002) identified a change in the Rapid Burster’s outburst interval, from 200 to 100 days, and a factor of two decrease in peak luminosity, in the year 2000. We confirm that the outburst interval has remained generally between 100 and 170 days from 2000 to 2023, while the peak flux has generally averaged $L_{X}=2\times 10^{37}$ erg/s in this time frame. However, in 2023, the average peak flux decreased to $L_{X}=1\times 10^{37}$ erg/s. We see two intervals in the full record that are twice as long as nearby intervals, suggesting the RXTE instruments may have missed outbursts in Dec. 2001 and Dec. 2003 (due to proximity of the Sun). We also see two unusually short intervals, of 28 and 37 days, suggesting the outbursts on MJD 56491 and 57131 (each observed only by MAXI) appear out of sequence, and thus might be due to other transients.
Figure 8: RXTE/PCA (green), RXTE/ASM (blue), MAXI/GSC (red), Swift/XRT (magenta), and INTEGRAL/JEM-X (cyan) lightcurves of MXB 1730-335 (the Rapid Burster). We also include the times of Type II bursts, as observed with the RXTE/PCA (Bagnoli et al., 2015), indicated by short black lines at the top. Upper left, 1996-2002; upper right, 2003-2009; lower left, 2010-2016; lower right, 2017-2023.
[…]
NS XRB Year/Month MJD Peak $F_{X}$ $L_{X}$ References
2-10 keV 2-10 keV
MXB 1730-335 1971/2 41028 Grindlay & Gursky (1977)
MXB 1730-335 1972/4 41448 Grindlay & Gursky (1977)
MXB 1730-335 1976/2 42840 1.7e-9 1.3e37 Lewin et al. (1976b)
MXB 1730-335 1977/3-5 43266 2.2e-9 1.7e37 White et al. (1978)
MXB 1730-335 1977/8 43408 Grindlay & Gursky (1977)
MXB 1730-335 1978/3 43598 Hoffman et al. (1978)
MXB 1730-335 1978/9 43796 Lewin & Joss (1981)
MXB 1730-335 1979/2-3 43936 Inoue et al. (1980); Lewin & Joss (1981)
MXB 1730-335 1979/8-9 44100 Inoue et al. (1980)
MXB 1730-335 1983/4 45437 Kaminker et al. (1990)
MXB 1730-335 1983/8 45564 1.2e-9 9e36 Kunieda et al. (1984); Barr et al. (1987)
MXB 1730-335 1984/7 45886 1.2e-9 9e36 Kawai et al. (1990)
MXB 1730-335 1985/8-9 46305 2.6e-9 2e37 Stella et al. (1988)
MXB 1730-335 1988/8 47385 Dotani et al. (1990)
MXB 1730-335 1989/1 47541 Lewin et al. (1993)
MXB 1730-335 1996/4 50189 6.3e-9 4.8e37 Guerriero et al. (1999),RXTE/ASM
MXB 1730-335 1996/10 50387 7.7e-9 5.8e37 Guerriero et al. (1999),RXTE/ASM
MXB 1730-335 1997/6 50625 9.3e-9 7.0e37 Guerriero et al. (1999),RXTE/ASM
MXB 1730-335 1998/1 50843 6.8e-9 5.2e37 Guerriero et al. (1999); Masetti et al. (2000),RXTE/ASM
MXB 1730-335 1998/8 51045 1.8e-8 1.4e38 Masetti (2002); Mahasena et al. (2003),RXTE/ASM
MXB 1730-335 1999/3 51252 6.9e-9 5.3e37 Masetti (2002); Mahasena et al. (2003),RXTE/PCA, RXTE/ASM
MXB 1730-335 1999/10 51452 6.8e-9 5.2e37 Masetti (2002); Marshall et al. (2001),RXTE/PCA, RXTE/ASM
MXB 1730-335 2000/3 51633 2.2e-9 1.6e37 Masetti (2002),RXTE/ASM, RXTE/PCA
MXB 1730-335 2000/7 51747 3.3e-9 2.5e37 Masetti (2002); Homer et al. (2001),RXTE/PCA, RXTE/ASM
MXB 1730-335 2000/10 51849 2.7e-9 2.0e37 Masetti (2002),RXTE/ASM, RXTE/PCA
MXB 1730-335 2001/2 51954 5.4e-9 4.1e37 Masetti (2002),RXTE/PCA, RXTE/ASM
MXB 1730-335 2001/5 52052 3.9e-9 3.0e37 Masetti (2002),RXTE/PCA, RXTE/ASM
MXB 1730-335 2001/9 52154 3.8e-9 2.9e37 Masetti (2002),RXTE/PCA, RXTE/ASM
MXB 1730-335 2002/4 52387 2.7e-9 2.1e37 Simon (2008),RXTE/PCA, RXTE/ASM
MXB 1730-335 2002/8 52510 1.8e-9 1.4e37 Simon (2008),RXTE/PCA, RXTE/ASM
MXB 1730-335 2003/2 52686 3.6e-9 2.7e37 Falanga et al. (2004); Simon (2008),RXTE/PCA, RXTE/ASM
MXB 1730-335 2003/5 52785 2.7e-9 2.0e37 Simon (2008),RXTE/ASM, RXTE/PCA
MXB 1730-335 2003/9 52886 1.7e-9 1.3e37 Simon (2008),RXTE/PCA, RXTE/ASM
MXB 1730-335 2004/4 53102 5.0e-9 3.8e37 Simon (2008),RXTE/PCA, RXTE/ASM
MXB 1730-335 2004/8 53237 4.5e-9 3.4e37 RXTE/PCA, RXTE/ASM
MXB 1730-335 2004/11 53330 8.0e-9 6.1e37 RXTE/ASM
MXB 1730-335 2005/4 53475 5.0e-9 3.8e37 RXTE/PCA, RXTE/ASM, IGR/JEM-X
MXB 1730-335 2005/8 53602 3.0e-9 2.3e37 RXTE/PCA, RXTE/ASM, IGR/JEM-X
MXB 1730-335 2006/1 53749 5.6e-9 4.3e37 RXTE/ASM
MXB 1730-335 2006/7 53918 3.7e-9 2.8e37 Bagnoli et al. (2015),RXTE/PCA, RXTE/ASM
MXB 1730-335 2006/12 54078 5.6e-9 4.2e37 RXTE/ASM
MXB 1730-335 2007/6 54265 4.8e-9 3.6e37 RXTE/ASM
MXB 1730-335 2007/11 54430 5.0e-9 3.8e37 RXTE/ASM
MXB 1730-335 2008/3 54546 2.2e-9 1.7e37 RXTE/PCA, RXTE/ASM, IGR/JEM-X
MXB 1730-335 2008/6 54620 3.1e-9 2.4e37 RXTE/PCA, RXTE/ASM
MXB 1730-335 2008/9 54727 2.9e-9 2.2e37 RXTE/ASM, RXTE/PCA, IGR/JEM-X
MXB 1730-335 2009/2 54876 1.6e-9 1.1e37 Sala et al. (2012),RXTE/ASM, RXTE/PCA, Swift/XRT, JEM-X
MXB 1730-335 2009/5 54976 3.2e-9 2.4e37 RXTE/ASM, RXTE/PCA
MXB 1730-335 2009/8 55074 4.9e-9 3.7e37 RXTE/ASM, RXTE/PCA, MAXI, IGR/JEM-X
MXB 1730-335 2010/2 55241 4.3e-9 3.3e37 RXTE/PCA, RXTE/ASM, MAXI, IGR/JEM-X
MXB 1730-335 2010/7 55395 4.7e-9 3.6e37 RXTE/ASM, MAXI, RXTE/PCA
MXB 1730-335 2011/1 55562 3.1e-9 2.4e37 MAXI
MXB 1730-335 2011/6 55723 2.3e-9 1.8e37 MAXI, RXTE/PCA
MXB 1730-335 2011/9 55811 3.6e-9 2.8e37 RXTE/PCA, MAXI, Swift/XRT, IGR/JEM-X
MXB 1730-335 2011/12 55902 1.9e-9 1.4e37 MAXI
MXB 1730-335 2012/4 56019 1.1e-9 8.4e36 MAXI, IGR/JEM-X
MXB 1730-335 2012/7 56136 4.1e-9 3.1e37 MAXI
MXB 1730-335 2012/11 56256 2.8e-9 2.1e37 MAXI
MXB 1730-335 2013/3 56352 3.3e-9 2.5e37 MAXI, IGR/JEM-X
MXB 1730-335 2013/6 56463 2.6e-9 2.0e37 MAXI
MXB 1730-335 2013/7 56491 3.7e-9 2.8e37 MAXI
MXB 1730-335 2013/11 56586 2.6e-9 2.0e37 IGR/JEM-X, MAXI
MXB 1730-335 2014/3 56730 2.0e-9 1.5e37 MAXI, IGR/JEM-X
MXB 1730-335 2014/9 56896 5.2e-9 3.9e37 IGR/JEM-X, MAXI
MXB 1730-335 2015/1 57051 5.1e-9 3.9e37 MAXI, IGR/JEM-X
MXB 1730-335 2015/4 57131 1.5e-9 1.1e37 MAXI
MXB 1730-335 2015/5 57168 1.8e-9 1.3e37 Swift/XRT, MAXI
MXB 1730-335 2015/10 57298 3.6e-9 2.8e37 van den Eijnden et al. (2017),MAXI, IGR/JEM-X, Swift/XRT
MXB 1730-335 2016/2 57422 3.5e-9 2.7e37 MAXI, IGR/JEM-X
MXB 1730-335 2016/6 57562 3.2e-9 2.4e37 MAXI
MXB 1730-335 2016/12 57732 2.7e-9 2.1e37 MAXI
MXB 1730-335 2017/3 57833 3.3e-9 2.5e37 MAXI, IGR/JEM-X
MXB 1730-335 2017/8 57975 2.8e-9 2.1e37 Chen et al. (2021), IGR/JEM-X, MAXI, Swift/XRT
MXB 1730-335 2017/11 58071 1.9e-9 1.4e37 MAXI
MXB 1730-335? 2018/3 58178 3.0e-9 2.2e37 Homan et al. (2018), MAXI, IGR/JEM-X
MXB 1730-335 2018/6 58292 1.7e-9 1.3e37 MAXI,Swift/XRT
MXB 1730-335 2018/9 58389 2.7e-9 2.0e37 MAXI
MXB 1730-335 2019/1 58490 1.4e-9 1.0e37 MAXI
MXB 1730-335 2019/3 58571 2.3e-9 1.8e37 MAXI, Swift/XRT
MXB 1730-335 2019/7 58677 2.2e-9 1.7e37 MAXI
MXB 1730-335 2019/11 58794 2.3e-9 1.7e37 MAXI
MXB 1730-335 2020/3 58920 3.5e-9 2.7e37 Swift/XRT, MAXI
MXB 1730-335 2020/8 59083 4.8e-9 3.6e37 MAXI
MXB 1730-335 2021/2 59247 3.3e-9 2.5e37 MAXI
MXB 1730-335 2021/8 59429 6.3e-10 4.8e36 MAXI
MXB 1730-335 2021/12 59553 2.0e-9 1.6e37 MAXI
MXB 1730-335 2022/5 59702 1.8e-9 1.4e37 MAXI
MXB 1730-335 2022/10 59867 1.3e-9 1.0e37 MAXI
MXB 1730-335 2023/1 59973 1.8e-9 1.4e37 MAXI
MXB 1730-335 2023/6 60105 9.9e-10 7.5e36 MAXI
MXB 1730-335 2023/8 60185 1.1e-9 8.6e36 MAXI
Outbursts of MXB 1730-335, the Rapid Burster in the globular cluster Liller 1. $F_{X}$ in ergs cm-2 s-1, $L_{X}$ in ergs s-1. Distance of 8 kpc assumed.
Note. —
GRS 1747-312; We use the distance estimate for the globular cluster Terzan 6 of 7.3±$\pm 0.4$ kpc from Baumgardt & Vasiliev (2021). GRS 1747-312’s outbursts have not been catalogd since 2008, so we turn to all-sky monitors to produce a full list (Table B). We assume $N_{H}=1.4\times 10^{22}$ cm-2 and a photon index of 1.0 (in’t Zand et al., 2000) to transform countrates to fluxes. For RXTE/ASM we multiply countrates by 3.6×$\times 10^{-10}$ to get 2-10 keV fluxes, for RXTE/PCA by $\times 10^{-12}$. For Swift/XRT data we multiply by 9.4×$\times 10^{-11}$ to get 2-10 keV fluxes, for MAXI by $8.7\times 10^{-9}$, and for INTEGRAL/JEM-X by $7.1\times 10^{-11}$. We identify outbursts in the RXTE/ASM, RXTE/PCA, INTEGRAL/JEM-X, and MAXI lightcurves when we see at least two detections (among all the lightcurves) with the countrate exceeding 3 times the error (see Fig. 9). This generally agrees with the identifications in in’t Zand et al. (2003) where they overlap, though the ASM data is noisy in this crowded region. We are also in general agreement with the identifications of Šimon (2009), but are not completely confident of the reality of their epochs 13, 20, and 21 (not included in our list), as the signal-to-noise of the individual detections is not above 3 sigma per day, though the timing of the outbursts appears plausible. The outbursts average every 138 days, with the average varying between 120 and 150 days over time, consistent with previous determinations (in’t Zand et al., 2003; Šimon, 2009). Peak fluxes average $L_{X}=6\times 10^{36}$ erg/s.
GRS 1747-312 is known to show regular eclipses (in’t Zand et al., 2000), which makes it possible to verify that an outburst from Terzan 6 is produced by this transient, given an accurate orbital ephemeris. Saji et al. (2016) noted a burst, and an absence of a predicted X-ray eclipse, during a Suzaku observation in Sept. 2009, when GRS 1747-312 was not detected in RXTE/PCA monitoring.111111van den Berg et al. (2024) point out that a spike is visible in RXTE/PCA scans on MJD 55088, two days earlier, but that pulsations in the scan reveal that this spike was produced by the known accreting millisecond pulsar IGR J17511–3057, then in outburst. More recently, van den Berg et al. (2024) have shown conclusive evidence that a second eclipsing, bursting X-ray binary, Terzan 6 X2, is located 0.7” from GRS 1747-312. We label the 2009 detection by Suzaku as ”Ter 6 X-2” in Table B, and Fig. 9.
Figure 9: RXTE/PCA (green), RXTE/ASM (blue), MAXI/GSC (red), Swift/XRT (magenta), and INTEGRAL/JEM-X (cyan) lightcurves of GRS 1747-312 (in Terzan 6). Upper left, 1996-2002; upper right, 2003-2009; lower left, 2010-2016; lower right, 2017-2023 (slightly different scale, to show one Swift data point).
[…]
NS XRB Year/Month E MJD Peak $F_{X}$ $L_{X}$ References
2-10 keV 2-10 keV
GRS 1747-312 1990/9 48143 8.8e-10 5.5e36 Pavlinsky et al. (1992); in’t Zand et al. (2000)
GRS 1747-312 1996/3 0 50158 1.6e-9 1.0e37 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1996/8 1 50319 1.3e-9 8.2e36 in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1997/1 2 50474 1.2e-9 7.6e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1997/5 3 50598 1.3e-9 8.5e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1997/10 4 50713 1.3e-9 8.2e36 in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1998/2 5 50865 1.1e-9 7.2e36 in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1998/6 6 50985 6.4e-10 4.0e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM
GRS 1747-312* 1998/9 7 51058 1.0e-9 6.4e36 in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM
GRS 1747-312 1999/1 8 51189 3.0e-9 1.9e37 Šimon (2009), RXTE/ASM
GRS 1747-312 1999/5 9 51308 1.1e-9 5.9e36 in’t Zand et al. (2000, 2003); Šimon (2009), PCA, RXTE/ASM
GRS 1747-312* 1999/10 10 51448 8.3e-10 5.3e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2000/2 11 51586 8.0e-10 5.1e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2000/8 12 51760 1.4e-9 8.6e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2001/5 14 52062 1.2e-9 7.6e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2001/10 15 52209 1.0e-9 6.5e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2002/2 16 52331 8.1e-10 5.2e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312* 2002/7 17 52468 8.9e-10 5.7e36 in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2002/10 18 52574 8.8e-10 5.6e36 Šimon (2009), RXTE/ASM
GRS 1747-312 2003/3 19 52713 1.2e-9 7.6e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2004/3 22 53083 7.2e-10 4.5e36 Markwardt & Swank (2004a); Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2004/8 23 53233 8.5e-10 5.4e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2004/11 24 53341 9.7e-10 6.1e36 Šimon (2009), RXTE/ASM
GRS 1747-312 2005/4 25 53482 7.8e-10 4.9e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2005/9 26 53624 5.8e-10 3.7e36 Šimon (2009), RXTE/ASM, PCA, IGR/JEM-X
GRS 1747-312 2006/2 27 53775 8.5e-10 5.4e36 Chenevez et al. (2006); Šimon (2009), RXTE/ASM, PCA, IGR/JEM-X
GRS 1747-312 2006/7 28 53923 9.5e-10 6.0e36 Šimon (2009) RXTE/ASM, PCA
GRS 1747-312 2007/1 29 54129 1.3e-9 8.2e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2007/6 30 54265 1.4e-9 8.7e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2007/11 31 54404 8.0e-10 5.1e36 Šimon (2009), RXTE/ASM, PCA
GRS 1747-312 2008/3 32 54556 6.9e-10 4.4e36 Šimon (2009), RXTE/ASM, PCA, IGR/JEM-X
GRS 1747-312 2008/9 54726 8.3e-10 5.3e36 RXTE/ASM, IGR/JEM-X
GRS 1747-312 2009/2 54869 1.2e-9 7.8e36 RXTE/ASM, PCA
GRS 1747-312 2009/6 55008 1.2e-9 7.7e36 RXTE/ASM, PCA
Ter 6 X-2 2009/9 55090 1.0e-11 6e34 Saji et al. (2016); van den Berg et al. (2024)
GRS 1747-312 2009/11 55145 1.2e-9 7.7e36 RXTE/ASM, MAXI
GRS 1747-312 2010/7 55402 7.3e-10 4.6e36 MAXI, RXTE/PCA, ASM
GRS 1747-312 2010/12 55544 1.0e-9 6.5e36 MAXI
GRS 1747-312 2011/4 55658 8.6e-10 5.5e36 MAXI, RXTE/PCA
GRS 1747-312 2011/7 55749 6.6e-10 4.2e36 MAXI, RXTE/PCA
GRS 1747-312 2011/9 55832 7.7e-10 4.9e36 IGR/JEM-X, RXTE/PCA, MAXI
GRS 1747-312 2011/12 55924 8.9e-10 5.6e36 MAXI
GRS 1747-312 2012/5 56072 6.6e-10 4.2e36 MAXI
GRS 1747-312 2012/10 56221 9.7e-10 6.1e36 MAXI, IGR/JEM-X
GRS 1747-312* 2013/3 56369 7.9e-10 5.0e36 Chenevez et al. (2013); Bahramian et al. (2013a), Swift/XRT, MAXI, JEM-X
GRS 1747-312 2013/7 56476 5.2e-10 3.3e36 MAXI
GRS 1747-312 2013/10 56594 7.0e-10 4.4e36 Chenevez & Sanchez-Fernandez (2013),MAXI, IGR/JEM-X
GRS 1747-312 2014/5 56786 1.2e-9 7.4e36 MAXI
GRS 1747-312 2014/10 56931 1.2e-9 7.7e36 MAXI, IGR/JEM-X
GRS 1747-312 2015/1 57030 1.2e-9 7.4e36 MAXI
GRS 1747-312 2015/3 57107 6.1e-10 3.9e36 MAXI, IGR/JEM-X
GRS 1747-312 2015/7 57227 5.1e-10 3.3e36 MAXI
GRS 1747-312 2015/12 57360 1.5e-9 9.4e36 MAXI
GRS 1747-312* 2016/4 57505 2.2e-9 1.4e37 Bahramian et al. (2016), MAXI, Swift/XRT
GRS 1747-312 2016/9 57649 1.3e-9 8.3e36 MAXI, IGR/JEM-X
GRS 1747-312 2017/1 57768 5.7e-10 3.6e36 MAXI
GRS 1747-312 2017/5 57905 8.0e-10 5.1e36 Swift/XRT, MAXI
GRS 1747-312 2017/10 58035 1.0e-9 6.5e36 Di Gesu et al. (2017); Bozzo & Sanna (2017), MAXI, Swift/XRT, IGR/JEM-X
GRS 1747-312 2018/2 58155 1.3e-9 8.1e36 MAXI, IGR/JEM-X
GRS 1747-312 2018/6 58295 5.7e-10 3.6e36 MAXI
GRS 1747-312 2019/6 58643 1.1e-9 7.2e36 Swift/XRT, MAXI
GRS 1747-312* 2019/10 58758 8.5e-10 5.4e36 Grebenev et al. (2019); Bahramian et al. (2019), Swift/XRT, MAXI
GRS 1747-312 2020/2 58876 5.5e-10 3.5e36 MAXI
GRS 1747-312 2020/5 58997 4.0e-10 2.5e36 MAXI
GRS 1747-312 2020/10 59137 6.6e-10 4.2e36 MAXI
GRS 1747-312 2021/2 59257 4.6e-10 2.9e36 MAXI
GRS 1747-312 2021/6 59383 8.6e-10 5.5e36 Swift/XRT
GRS 1747-312 2021/11 59536 1.4e-9 8.7e36 MAXI
GRS 1747-312 2022/3 59651 7.4e-10 4.7e36 MAXI
GRS 1747-312 2022/7 59774 5.3e-10 3.4e36 MAXI, Swift/XRT
GRS 1747-312 2022/10 59879 1.0e-11 6.6e35 Swift/XRT
GRS 1747-312 2023/1 59958 6.4e-10 4.1e36 MAXI, Swift/XRT
GRS 1747-312 2023/3 60034 6.1e-10 3.9e36 MAXI
GRS 1747-312 2023/6 60120 9.2e-10 5.9e36 MAXI