### v3

# Catalog of outbursts of neutron star LMXBs

Abstract
Many X-ray binaries are transiently accreting. Having statistics on their recurrence times is helpful to address questions related to binary evolution and populations, as well as the physics of binary systems. We compile a catalog of known outbursts of 87 transient neutron star (identified through bursts or pulsations) low-mass X-ray binaries, until mid-2025. Most outbursts are taken from the literature, but we also identify some outbursts from public X-ray monitoring lightcurves. We find 109 outbursts not previously identified in the literature; most are from the frequent transients GRS 1747-312, and the Rapid Burster MXB 1730-335, though we suspect that two outbursts from Liller 1 may be from another transient, besides the Rapid Burster. We also find new outbursts for 10 other systems, and verify substantial quiescent intervals for XMM J174457-2850.3, XMMU J174716.1-281048, and AX J1754.2-2754. Outburst detection has been relatively efficient since 1996 for outbursts above (2-10 keV) ergs cm-2 s-1. While several systems have many known outbursts, 40 of the 87 systems we track have zero or one recorded outburst between 1996 and 2023. This suggests that many faint Galactic Center X-ray binaries may be neutron star X-ray binaries, though we cannot completely rule out the proposition that most neutron star X-ray binaries undergo frequent outbursts below all-sky monitor detection limits.
1 Introduction
The total number of mass-transferring neutron star low-mass X-ray binaries in the Galaxy is thought to be of the order of (Portegies Zwart et al., 1997; Kiel & Hurley, 2006; Jonker et al., 2011; van Haaften et al., 2015). However, only a small fraction of these systems have been detected, typically during outbursts in which their X-ray luminosity increases by a factor of – compared to quiescence. Understanding the full population of these systems, and their binary evolution, is helpful for us to understand the physics of binary systems (e.g. Podsiadlowski et al., 2002; Postnov & Yungelson, 2014; Van et al., 2019), and also to understand the products of their evolution, such as millisecond pulsars (Alpar & Shaham, 1985; Verbunt, 1993; Tauris & Savonije, 1999; Tauris et al., 2012), which may produce the excess gamma-rays seen from the Galactic bulge (Wang et al., 2005; Gonthier et al., 2018). One question of current interest is the nature of X-ray sources of intermediate luminosity ((2-10 keV) erg/s) detected near the Galactic Centre (Muno et al., 2005; Hailey et al., 2018; Mori et al., 2021; Maccarone et al., 2022). The objects in this region that show no outbursts have been argued to be mostly black hole LMXBs, based on an argument that neutron star LMXBs do not have long enough recurrence times. Having detailed statistics on X-ray outbursts from neutron stars can be helpful in beginning to answer questions such as these.
Large numbers of transient X-ray binaries have been detected by a wide range of X-ray surveys. For many of these X-ray binaries, the nature of the accreting object is not confidently known. However, many transient X-ray binaries have secure identifications of the accretor as neutron stars, typically through the identification of thermonuclear X-ray bursts from the neutron star surface (Lewin et al., 1993), or through detection of pulsation from the neutron star spin (Wijnands & van der Klis, 1998). Verification of the opposite case, that the accretor is a black hole, is harder, requiring optical or infrared spectroscopy to measure, through radial velocities, the mass of the unseen companion, and show it is above the generally accepted maximum neutron star mass of 3 M⊙ (Casares & Jonker, 2014).
A number of catalogs of X-ray binaries have been produced over the years. The catalogs of Liu et al. (2001, 2007) have been particularly well-cited, but are now 17 years old. Ritter & Kolb (2003) have an extensive and updated catalog of X-ray binaries and cataclysmic variables with known periods, but do not include objects without known (or suggested) orbital periods. More recent catalogs have been published of black hole X-ray binaries, one focused on compiling detailed lightcurves (Tetarenko et al., 2016b), and one focusing on multiwavelength studies (Corral-Santana et al., 2016). We are interested in the particular problem of understanding the frequency of X-ray outbursts from neutron star LMXBs, but found that we lacked a catalog to answer this question. The catalogs of Avakyan et al. 2023 and Fortin et al. (2024) were released when this work was largely complete, and while they are very useful, they also do not list outbursts. Yan & Yu (2015) and Lin et al. (2019) construct tables of outburst frequency for a number of X-ray binaries with known orbital periods, but these works only include 14 and 19 NS LMXBs respectively. Yan & Yu (2015) do include a detailed listing of the outbursts they include, which is helpful. We identify several more outbursts from 4U 1608-52, XTE J1709-267, MXB 1730-33, XTE J1739-285, 1A 1744–361, SAX J1748.9-2021, and Aquila X-1 in the time range these authors cover, typically fainter outbursts.
§ 2 gives details of the construction of our catalog. In § 3, we make a first use of this catalog to address the question of the nature of intermediate luminosity X-ray sources near the Galactic Centre.
2 Catalog
To construct such a catalog, we used two clear methods to verify the presence of a NS in an LMXB; presence of thermonuclear bursts, and/or presence of coherent millisecond pulsations. We took lists of confirmed neutron star LMXBs from the MINBAR X-ray burst catalog (Galloway et al., 2020), which identified 115 known X-ray bursters using data from BeppoSAX, RXTE, and INTEGRAL (1996-2012), plus the literature.
We omit 41 NS LMXBs in the MINBAR catalog that appear to be persistent. Identifying whether a NS LMXB is persistent can be tricky. Seven known NS transients have had outbursts lasting over a decade (EXO 0748-676, KS 1731-260, XB 1733-30, IGR J17062-6143, HETE J1900.1-2455, 1H 1905+000, and 4U 2129+47), and it is possible that systems we list as persistent will turn off in the future. Other systems are persistently active at a low level, then brighten substantially for weeks or months (e.g. 4U 1705-32, Negoro et al. 2021a; in’t Zand et al. 2007). UW CrB (1E 1603.6+2600) at a distance of kpc (Gaia parallax, Bailer-Jones et al., 2021) appears unusually faint for an actively accreting system, at erg s-1, but this is thought to be due to extreme obscuration of its intrinsic X-ray luminosity (see e.g. Hakala et al., 2005). We list the systems we consider persistent, verified bursting or pulsing NS XRBs in Table 10, in Appendix D.
We omit a few transient bursters; the position of XB 1940-04 (Murakami et al., 1983) is very poorly known. The eccentric X-ray binary Cir X-1 (or 3A 1516-569) has outbursts that typically reflect its orbital period, not disk outbursts (Murdin et al., 1980). During the period from MJD=55000–58500, it was unusually under-luminous, and showed occasional outbursts that were not at exact multiples of its orbital period, suggesting outbursts from a disk. We catalog those outbursts in Appendix B, but due to Cir X-1’s differences in accretion mode compared to other LMXBs, we omit it from most of our compiled analyses. We exclude 1A 1742-289 (or A1742-289), as it is not clear that the burst activity seen from MXB 1743-29 in the Galactic Center in 1976 (Lewin et al., 1976b) can be attributed to this radio-bright (Davies et al., 1976) source (we note that this is definitely not the same object as AX J1745.6-2901; Kennea & Skinner 1996). This leaves 70 sources from the MINBAR catalog (71 when including Cir X-1).
We add to the MINBAR list the known recent bursters111as updated by Jean in ’t Zand; https://burst.sci.monash.edu/wiki/index.php?n=MINBAR.SourceTable 4U 1730-22, MAXI J1816-195, MAXI J0556-332, MAXI J0911-655, Terzan 6 X-2 (see Appendix § B), and SRGA J144459.2-604207 222We include SRGA J144459.2-604207 in our list for completeness, but our statistical calculations include data only through 2023..
We also add known transient accreting millisecond pulsars that are not known bursters, from the Patruno & Watts (2021) catalog and more recent literature, XTE J0929-314, NGC 6440 X-2, XTE J1751-305, Swift J1756.9-2508, XTE J1807-294, IGR J17494-3030, IGR J16597-3704, and MAXI J1957+032. Finally, we include the transient slow (2.14 Hz) accreting pulsar GRO J1744-28. (We also list the two persistent LMXBs that are pulsars in Table 10.) This gives us a total of 86 transient neutron star low-mass X-ray binaries, 87 if including Cir X-1.
We performed literature searches for this list of transients, and identify the known X-ray outbursts from each system. For each outburst, we attempt to determine the peak observed 2-10 keV X-ray luminosity. This will have some significant uncertainty, since not every outburst is caught at its peak, some outbursts are observed only in other energy ranges, and many distances are not well-known. We attempt to use the best-fit spectrum to convert measurements to 2-10 keV fluxes, but we only attempt this with simple power-law spectra (we do not try to convert complex Comptonized models). Where information is lacking, we use the best literature estimate and an assumed photon index of 2.1 to make conversions, or use literature estimates of the 2-10 keV flux in mCrab (assuming 1 Crab = ergs cm-2 s-1 ). A few bursters have recorded bursts without any recorded outbursts (Gotthelf & Kulkarni, 1997; Cornelisse et al., 2002b); we list these in Table 4 in Appendix C.
We present the catalog of known outbursts of neutron star LMXBs in Table 1, for which we provide a sample here. The full table is in (Table A) in Appendix A). 333We also maintain an updated version at https://tinyurl.com/5euncwmh. We include a line for each recorded outburst, with a literature reference for each outburst if available. Several objects with repeated bright outbursts can be easily spotted in the public lightcurves of the RXTE PCA Bulge Scan (Swank & Markwardt, 2001)444https://asd.gsfc.nasa.gov/Craig.Markwardt//galscan/main.html, the RXTE All-Sky Monitor (Levine et al., 1996a)555https://heasarc.gsfc.nasa.gov/docs/xte/asm_products.html and/or in MAXI GSC observations (Matsuoka et al., 2009; Suguzaki et al., 2011)666http://maxi.riken.jp/top/index.html. We have added some obvious outbursts visible in these surveys to the list (e.g. outbursts from XTE J1709-267 in 2005, 2009, 2018/19, and 2022). The line for defining an ‘outburst’ can be unclear; for simplicity we require an increase of a factor of 10 from the quiescent flux (if known), and erg/s.
We separate the objects with many outbursts into separate tables, placed in Appendix B. For the most frequently repeating objects with bright outbursts (of order once per year or more), 4U 1608-52, MXB 1730-33 (the Rapid Burster in Liller 1), Aquila X-1, GRS 1747-312 (the eclipser in Terzan 6), and Cir X-1, we use the three public lightcurves above, as well as Swift/XRT public lightcurves (Evans et al., 2007)777https://www.swift.ac.uk/user_objects/, and the INTEGRAL JEM-X Galactic Bulge Monitoring lightcurves (Kuulkers et al., 2007a)888http://integral.esac.esa.int/BULGE/ to develop the best possible list of outbursts from 1996 to 2023, along with literature identification of outbursts before 1996. For the objects with the most frequent outbursts, we also record the MJD date of the outburst peak (as best we can tell). We place the outbursts of NGC 6440 X-2 in a separate table in Appendix B as well, as their high frequency requires MJD dates to differentiate them.
When identifying additional outbursts, we generally search for multiple points that are above 3 sigma. The MAXI baseline may often vary substantially, so we additionally impose that the potential outburst must rise significantly above the quiescent baseline.
We find additional outbursts, not mentioned in the literature, for 12 systems; XTE J1709-267 (4 outbursts), XTE J1701-407 (1), XTE J1739-285 (2), 1A 1744-361 (2), XMMU J174716.1-281048 (1), AX J1754.2-2754 (3), SAX J1810.8-2609 (1), 4U 1608-52 (3), MXB 1730-335 (aka the Rapid Burster, 52), GRS 1747-312 (37), NGC 6440 X-2 (2), and Aquila X-1 (1), for a total of 109 newly identified outbursts. We also identify 42 additional outbursts from Cir X-1 not mentioned in the literature between 2006 and 2018, though we list these separately as their nature is somewhat ambiguous.
We give comments on each X-ray binary’s outburst history in Appendices A and B, including the reference we use for the distance, and any discrepancies we identify in the literature regarding their outburst histories. Here we highlight a few key points;
XMM J174457-2850.3, which showed frequent low-level outbursting behavior between 2001 and 2012, has seen its activity subside, with only one short outburst since 2012.
We use Chandra, XMM-Newton, and Swift/XRT observations to verify that XMMU J174716.1-281048 declined from a 10-year-long outburst in 2012, reached full quiescence by 2018, showed another (short, 3 months) outburst in 2021, and then returned again to quiescence.
We use Swift/XRT and Chandra observations to show that AX J1754.2-2754 has had (2-10 keV) erg/s for much of the past eight years, in contrast with its quasi-persistent behavior in 1999-2017 (Bassa et al., 2008).
SAX J1806.5-2215’s long outburst, which started in Feb. 2011, ended sometime between July 2017 and January 2018.
We catalog 52 outbursts from MXB 1730-335 (aka the Rapid Burster) that have not previously been identified in the literature. Some of these may not be produced by the Rapid Burster, but by other LMXBs in the host cluster Liller 1; another such outbursting transient has been verified (Homan et al., 2018), and we do identify two unusually short intervals (28 and 37 days) between outbursts that are good candidates for other transients. However, the relatively regular outbursts, and detection of Type II bursts in most cases where instruments are sensitive to them, indicate most outbursts are due to the Rapid Burster. Masetti (2002) noted that the average outburst interval decreased from 200 days to 100 days around 2000, with the peak flux also dropping by a factor of 2. Since 2000, the average outburst interval has varied between 100 and 170 days, and the peak has remained around erg/s.
We catalog 37 previously unidentified outbursts from GRS 1747-312 (in the globular cluster Terzan 6). The outburst interval averages 138 days, with significant scatter, but regular enough to indicate that most outbursts are indeed due to a single source. A second transient bursting source, Terzan 6 X-2, has recently been identified in Terzan 6 (Saji et al., 2016; van den Berg et al., 2024), but appears not to reach such high fluxes as GRS 1747-312.
In §3 we summarize some useful information on monitoring instruments, use the information in these tables to draw some general conclusions about our sensitivity to X-ray outbursts, and apply this information to consideration of interesting X-ray sources near the Galactic Center.
| NS XRB | Begin | End | Peak | Index | d | References | ||
| 2-10 keV | cm-2 | 2-10 keV | kpc | |||||
| IGR J00291+5934 | 1998/11 | 1998/11 | 4.5e-10 | - | - | 9.6e35 | 4.2 | Remillard (2004) |
| IGR J00291+5934 | 2001/9 | 2001/9 | 3.9e-10 | - | - | 8.2e35 | 4.2 | Remillard (2004) |
| IGR J00291+5934 | 2004/12 | 2004/12 | 6.1e-10 | 7e21 | 1.7 | 1.3e36 | 4.2 | Eckert et al. (2004); Galloway et al. (2005) |
| IGR J00291+5934 | 2008/8 | 2008/10 | 6.1e-10 | - | 1.7 | 1.3e36 | 4.2 | Chakrabarty et al. (2008); Lewis et al. (2010) |
| IGR J00291+5934 | 2015/8 | 2015/8 | 5.6e-10 | 1e22 | 1.7 | 1.2e36 | 4.2 | Cummings et al. (2015a); De Falco et al. (2017) |
| MAXI J0556-332 | 2011/1 | 2012/5 | 1.9e-09 | - | - | 4.4e38 | 43.6 | Matsumura et al. (2011); Sugizaki et al. (2013) |
| MAXI J0556-332 | 2012/10 | 2012/12 | 4e-10 | 3.2e20 | - | 9e37 | 43.6 | Sugizaki et al. (2012); Parikh et al. (2017a) |
| MAXI J0556-332 | 2016/1 | 2016/3 | 7.7e-10 | - | - | 1.8e38 | 43.6 | Negoro et al. (2016a); Russell et al. (2016) |
| MAXI J0556-332 | 2020/4 | 2020/11 | 8.6e-10 | - | - | 2.0e38 | 43.6 | Negoro et al. (2020a); Page et al. (2022) |
| EXO 0748-676 | 1984/7 | 2008/8 | 1.1e-9 | 1.2e21 | 1.9 | 7e36 | 7.4 | Parmar et al. (1986); Degenaar et al. (2009) |
3 Discussion
3.1 Monitoring instruments
We first review key information about some relevant monitoring instruments.
The first effective all-sky X-ray monitor was a scintillation 3-12 keV X-ray detector on Vela 5B, which flew from 1969-1979. Its limited collimation and thus high background limited its sensitivity to of order 0.3 Crabs (Priedhorsky & Holt, 1987). Ariel V (1974-1980) had a slightly more sensitive (perhaps 0.2 Crabs) 3-6 keV all-sky monitor, operated as a pinhole camera (Holt, 1976). In 1987-1991, Ginga’s ASM operated from 1-20 keV with sensitivity down to 50 mCrab (Tsunemi et al., 1989).
RXTE/ASM monitoring produced a dramatic increase in detected outbursts, from January 1996 until 2009-2011 (Levine et al., 1996a)999The RXTE/ASM sensitivity decayed from 2009 until RXTE was turned off in 2011.. RXTE/ASM achieved roughly 15 mCrab ( ergs cm-2 s-1, 3-sigma) sensitivity per day, in 2-10 keV, for known sources. The RXTE/ASM team could identify new sources as faint as 50 mCrab with one day’s data, while integrating for a week could delve down to 7 mCrab, in uncrowded regions (Bradt et al., 2001). Indeed, detections from 1996 on did reach , though the completeness limit was higher, perhaps ergs cm-2 s-1, or 50 mCrab. The RXTE/ASM covered most sources daily, except for about 30 days each year when the Sun was nearest for sources in the ecliptic.
From 1996 to 2002, in each spring and fall, BeppoSAX conducted regular monitoring of the Galactic Bulge with its Wide-Field Camera (WFC), with a 40o by 40o field of view, 5’ angular resolution, and sensitivity of about 15 mCrab in 2-28 keV (Heise et al., 1999; in’t Zand et al., 2004). Monitoring of the Galactic Center was performed weekly for two months in the spring and fall (Heise et al., 1999). Jager et al. (1997) estimated the typical observation sensitivity as 10 mCrab at high galactic latitudes, and 30 mCrab at the Galactic Center. BeppoSAX’s WFC discovered a large number of faint ( erg/s) X-ray transients in the Galactic Bulge, and also several X-ray bursts from transients that were too faint for the WFC to detect individually (Cornelisse et al., 2002b).
From 1999 to 2011, RXTE also conducted monitoring of the Galactic Bulge using slews of the PCA instrument (Swank & Markwardt, 2001). These Bulge scans could detect new uncrowded sources down to 0.5 mCrab (Swank & Markwardt, 2001), but in practice, in this crowded region, they reached a sensitivity of, at best, 5 mCrab, or ergs cm-2 s-1 in 2.5-10 keV. Notably, the RXTE/PCA was able to detect several faint (5 mCrab) outbursts of NGC 6440 X-2 in 2009-2011 (Heinke et al., 2010), but these were only barely visible in the Bulge Scan lightcurves. The PCA Bulge scans initially covered about 250 square degrees twice weekly (Markwardt et al., 2002b), extended to 500 square degrees in 2003 (Markwardt, 2006), and for 2009-2011 added the Aquila region, to cover over 700 square degrees (Markwardt & Swank, 2010). These scans were paused for the 2 months when the Sun is closest to the Galactic Center (Swank & Markwardt, 2001).
The INTEGRAL satellite (launched 2002, decommissioned in Feb. 2025) has two relevant coded-aperture instruments: IBIS (Ubertini et al., 2003) has peak sensitivity in the 20-60 keV range with a 20 degree field of view (at half sensitivity) and 12’ angular resolution, while JEM-X (Lund et al., 2003) observes in 3-35 keV, with a 7.5 degree field of view and 3.4’ angular resolution. The INTEGRAL Galactic Bulge hard X-ray monitoring program has been particularly productive, with 5-15 mCrab sensitivity per 1800 s observation (every three days), and summed detections reaching down to 2.4 mCrab (Kuulkers et al., 2007a). INTEGRAL could observe the Bulge only 1/3 of the year, for 2 months in each spring and fall. Thirteen of our transients have principal names starting with IGR (only XTE matches this number), suggesting INTEGRAL’s impact. There appears (Fig. 2) to be a small decline in detections in the past few years, which may be due to declining sensitivity of INTEGRAL’s IBIS/ISGRI instrument.101010The Integral Galactic Bulge Monitoring Program’s website, http://integral.esac.esa.int/BULGE/, presents 18-40 keV IBIS/ISGRI lightcurves for sources, which show increasing noise from 2016 and stop in 2018.
Swift/BAT’s hard X-ray (15-50 keV) monitoring program has operated since 2006, with a daily 16 mCrab detection limit (Krimm et al., 2013). Swift’s XRT instrument (narrow-field imaging in 0.3-10 keV) has also been vital in tracking transients, due to its flexible scheduling, ease of ToO requests, good sensitivity, and simplicity of data analysis (e.g. Evans et al., 2007).
The MAXI instrument (on the International Space Station) has two camera groups, of which the Gas Slit Cameras, or MAXI/GSC (Mihara et al., 2011; Sugizaki et al., 2011) has been significantly more productive. MAXI/GSC has a field of view of 1.5o by 160o, and monitors 85% of the sky every 92 minutes, reaching 95% of the sky each day (Negoro et al., 2016b). MAXI/GSC has monitored the sky since 2009 daily in 2-20 keV, with the best sensitivity in 3-10 keV. Typical sensitivities for 4-day integrations reach 7 mCrab (Negoro et al., 2016b). Positions of isolated sources are typically determined to 0.1o, but MAXI/GSC’s power to resolve the Galactic Center is limited. MAXI’s efficient and automated nova search algorithms (Negoro et al., 2016b) have been highly productive, identifying transients which can then be followed up by other missions to get precise positions.
The SRG spacecraft scans the sky twice per year, and contains two instruments: ART-XC in 4-30 keV, and eROSITA in 0.2-8 keV (Sunyaev et al., 2021). eROSITA completed 4 of 8 planned all-sky surveys (2019-2022), with sensitivity per survey (200 s exposure) reaching ergs cm-2 s-1 in 2-4 keV (Sunyaev et al., 2021), before the Ukraine war terminated Russian/German cooperation. ART-XC continues operating, and reaches 1 mCrab sensitivity in 4-12 keV each survey sweep (Sunyaev et al., 2021). ART-XC detected a new X-ray binary, SRGA J144459.2-604207 (Molkov et al., 2024), during the preparation of this catalog.
Two new instruments capable of transient LMXB identifications began surveying the sky within the last year. The Einstein Probe is a Chinese soft X-ray mission launched on January 9, 2024. Its Wide-Field X-ray Telescope (WXT) monitors the entire night sky on a daily basis, using lobster-eye optics to achieve a 3600 square degree instantaneous field of view, a spatial resolution of 5’, and a nominal 0.5-4 keV sensitivity of 0.8 mCrab in 1000 s (Yuan et al., 2022). The Follow-up X-ray Telescope (FXT) has 30” angular resolution, and sensitivity similar to Swift/XRT, reaching (0.5-2 keV) erg/s in 1500 s (Yuan et al., 2022).
Finally, the Space Variable Objects Monitor (SVOM) is a joint Chinese/French mission launched on 22 June 2024, with a science focus on gamma-ray bursts. Its wide-field hard X-ray (4-150 keV) detector, ECLAIRS, has a sensitivity of 50 mCrab per orbit, with a PSF of 52’ (Godet et al., 2014). It also hosts a narrow-field soft X-ray (0.2-10 keV) detector, MXT, with angular resolution of 4.5’ and a sensitivity reaching to (0.2-10 keV) with 10 ks of integration. SVOM’s mission plan of generally pointing away from the Galactic Plane will tilt its science toward GRBs, but it is likely to also detect LMXB outbursts (Wei et al., 2016).
Both INTEGRAL/IBIS and Swift/BAT are sensitive to harder X-rays than the 2-10 keV range we focus on here. Thus, they will be more likely to detect LMXBs in hard states, vs. soft states at the same 2-10 keV flux. The ratio of Crab-normalized Swift/BAT 15-50 keV vs. MAXI/GSC 2-10 (or 4-10) keV transient NS LMXB fluxes is about 1 in the hard state, and often 0.1 in the soft state (e.g. Asai et al. 2015, Bahramian et al. 2014; cf. Tetarenko et al. 2016b for BH LMXBs, which can reach lower ratios in the soft state). Most transient NS LMXBs observed near the limit of all-sky monitor detection (5 mCrab) at distances of 10 kpc (thus, (2-10) erg/s) are found to be in the hard spectral state (Maccarone & Coppi, 2003; Asai et al., 2015). However, INTEGRAL/IBIS and Swift/BAT would have a 2-10 keV detection threshold roughly 10 times higher, for soft-state transients.
3.2 General trends
In Fig. 1 we show the peak fluxes from all transient neutron star LMXBs (a few lie outside our bounds), with the most frequent outbursters indicated with different colors. In Fig. 2, we show the number of outbursts per year (omitting Cir X-1’s many outbursts in some years, which can be rather distorting).
From Fig. 2, we see that 1996 marks a dramatic change in detected outburst frequency. A significant peak appears around 2009-2011 in all outbursts. Fig. 2 of Tetarenko et al. (2016b), transient outbursts of BH LMXBs, reveals a similar peak in 2007-2012. However, this peak is not significant in brighter NS LMXB outbursts with ergs cm-2 s-1, suggesting that fainter outbursts were more detectable in that time frame (e.g. by the RXTE PCA bulge monitoring program; which detected, e.g., the NGC 6440 X-2 outbursts at ergs cm-2 s-1; Heinke et al. 2010).
We then construct a histogram of fluxes from all outbursts since 1996 (Fig. 3, blue triangles). This is sharply peaked at ergs cm-2 s-1, which might suggest that our data becomes highly incomplete below ergs cm-2 s-1. However, if we exclude the four most frequent outbursters (Aquila X-1, 4U 1608-52, MXB 1730-335, and GRS 1747-312), all of which have relatively long orbital periods, high mass transfer rates, and high outburst fluxes, we see a rather different histogram (Fig. 3, red squares), which does not peak until the bin ergs cm-2 s-1. This suggests that significant incompleteness is happening in the next lower bin, covering ergs cm-2 s-1. Referring to Fig. 1, it seems that the majority of monitors have a limit around ergs cm-2 s-1, so the peak here is likely due to sensitivity limits. Below ergs cm-2 s-1, most detections use pointed instruments (e.g. Swift/XRT, or RXTE in pointing mode).
We can also look at the distribution of peak in detected outbursts (Fig. 4). (We must keep in mind that the majority of these values are likely overestimates, due to the assumed distances being often upper limits, if the observed bursts were not actually Eddington-limited as assumed.) The distribution looks generally similar to the distribution of fluxes, due to the concentration of transient LMXBs in the Galactic Bulge, but due to the variation in distance, the peak is less sharp. An equivalent plot for BH LMXB systems (Fig. 15 in Tetarenko et al. 2016b111111Including a shift down by roughly a factor 3 to move from bolometric to 2-10 keV luminosity.) peaks close to erg/s, with a tail extending down to erg/s. The tendency for BHs to have higher may be attributed to BHs’ bias towards longer orbital periods, which are correlated with higher peak (Wu et al., 2010). At very short periods, even the expected peak luminosities leave black holes in radiatively inefficient states, leading to a strong selection effect against low-luminosity black holes (e.g. Arur & Maccarone 2018). Another selection effect is that verifying a black hole in an LMXB typically requires phase-resolved spectroscopy, which is more difficult in a short-period, thus fainter, system; while verifying a neutron star uses X-ray bursts or pulsations, which aren’t tightly correlated with peak .
Intriguingly, there is a sharper peak in the erg/s bin than appears in the flux histogram. If this peak were due entirely to limited sensitivity below it, we would expect this peak to be less sharp in the histogram than the histogram. It is possible (but not certain) that the number of outbursts falls off below erg/s (”very faint” X-ray binary outbursts). A gap in the (persistent) X-ray luminosity function between and erg/s has previously been identified in globular clusters (Hertz & Grindlay, 1983; Verbunt & Lewin, 2006; Bahramian et al., 2020b), where the sources above, and just below, this gap consist predominantly of neutron star LMXBs. A dropoff below erg/s has also been suggested in the bulge of M31 (Voss & Gilfanov, 2007), in globular clusters in nearby galaxies (Zhang et al., 2011), and possibly (but with inadequate data) for the Milky Way (Sazonov et al., 2006; Bahramian et al., 2021). On the other hand, extensive Swift, XMM-Newton and Chandra monitoring of the Galactic Center identified 28 outbursts from 13 LMXBs, between 2000 and 2014, including 11 between erg/s and erg/s, which argues strongly against any cutoff (Wijnands et al., 2006b; Degenaar et al., 2015). This apparent discrepancy may be reconciled if most persistently-bright systems have erg/s, but transient outbursts are often fainter, so that a given snapshot would find a gap, but a list of peak outburst luminosities would not. Following the standard disk-instability model, the only LMXBs that could remain persistently below erg/s will be a small range of periods for ultracompact systems (e.g. in’t Zand et al., 2007; van Haaften et al., 2013). If this is indeed the case, then more sensitive X-ray monitoring programs (e.g. proposed for STROBE-X, Ray et al. 2019) would find substantially more X-ray outbursts.
3.3 Application to the Galactic Center
Hailey et al. (2018) identified a dozen Chandra X-ray sources within a few parsecs of the Galactic Center that have non-thermal X-ray spectra (best fit with a power law with photon index between 1.5 and 2.9), which along with their values indicates these are quiescent LMXBs. Hailey et al. (2018) and Mori et al. (2021) suggest that these objects must be black hole LMXBs, arguing that all neutron star LMXBs have short recurrence times, generally less than 5 years, while black hole LMXBs can have longer recurrence times. Maccarone et al. (2022) argued that a number of known neutron star LMXB systems do have longer recurrence times, and that the non-thermal quiescent Galactic Center X-ray sources may be dominated by neutron star LMXBs. Mori et al. (2021) and Mori et al. (2022) argued that most neutron star LMXBs undergo frequent outbursts, based on the history of the known NS transients studied in the Galactic Center. This argument assumes that the confirmed neutron star transients are representative of the underlying population, and not subject to biases; for instance, that the more frequent outbursters are easier to confirm as neutron stars (e.g. they produce more X-ray bursts).
Clearer conclusions on the recurrence times of neutron star LMXBs require a thorough compilation of known outbursts, which we performed in this paper, allowing us to comment on the fractions of (observed) neutron star LMXBs with different recurrence times. These conclusions are, however, subject to various systematic uncertainties, due principally to the limited sensitivity and angular resolution of X-ray all-sky monitors.
A key question we can ask is what fraction of transient neutron star LMXBs have a recurrence time greater than 10 years, from the data available. We can answer this by checking for gaps of at least 10 years between 1996 and 2022 (the time frame with the best sensitivity). We identify, for each object, the probability (based on the recorded 1996-2022 outbursts) of not seeing outbursts during a continuous 10-year period. (We allow a continuous time interval to wrap from 2022 to 1996, for this purpose.) For instance, Cen X-4 has a 100% chance of not seeing any outburst in a random 10-year period of this range, while MXB 1659-29 has only a 22% chance of having no outbursts in a 10-year block, and SAX J1808.4-3658 has a 0% chance of seeing no outbursts in a 10-year block. Among 84 systems we study (omitting Terzan 6 X-2, Terzan 5 X-2 and X-3, due to confusion), only 24 do not show any 10-year-long gap, suggesting that the majority have a chance of hiding for 10 years or longer. Randomly selecting a contiguous ten-year period from a random neutron star in our sample gives an average 37% probability of finding no outbursts.
This calculation does not include systems that have been quiescent over the past 30+ years, such as the quiescent LMXBs observed in many globular clusters (e.g. Heinke et al., 2003b; Guillot et al., 2011; Bahramian et al., 2015). As only 15 neutron star LMXBs have been observed to outburst from globular clusters (Bahramian & Degenaar, 2023), but of order 200 quiescent neutron star LMXBs are estimated in Galactic globular clusters (Heinke et al., 2005), this would suggest a neutron star LMXB population over ten times larger than observed. Some evolutionary calculations support the likely existence of such binaries (Meyer-Hofmeister et al., 2024).
The principal caveat is that all-sky monitors have limited sensitivity. We infer that we are missing most outbursts below erg/s; even brighter outbursts for 10% of the year when the Sun is nearby; and some (not easily quantifiable) due to crowding with bright nearby sources. The Swift campaigns on the Galactic Center show that the three closest verified neutron star LMXBs show at least one fainter outburst as well as bright outbursts; Degenaar et al. (2015) list 4 outbursts with peak erg/s, and 11 outbursts with peaks below this. Since 1996, our larger databases give 8 and 22, respectively. Mori et al. (2022) argue that the much larger survey frequency of the Galactic Center gives it the only well-constrained outburst histories, particularly regarding very faint X-ray transient behavior.
However, there are also nine other transients, showing very-faint X-ray outbursts (peak erg/s), that have been detected in the Galactic Center region since 2000, which have not been confirmed as neutron stars through X-ray bursts, and have also not shown previous evidence for being black holes (Degenaar et al., 2015). These 9 very-faint transients have each had only one or two outbursts within the past 23 years, suggesting relatively long recurrence times (20 years), and that a larger population of such transients may exist there. It is generally difficult to attribute X-ray bursts observed near the Galactic Center with RXTE (Galloway et al., 2008), and very faint X-ray binaries may be expected to burst only rarely (Peng et al., 2007; Lin et al., 2019). The fact that known black hole outbursts tend to be brighter than known neutron star outbursts (Yan & Yu, 2015) suggests that some, or most, of these very faint X-ray binaries are neutron stars. However, we have observed few short-period black hole systems, and it is possible that short-period black hole LMXBs are relatively faint; cf. Knevitt et al. 2014.
One empirical test between neutron star and black hole LMXBs, introduced by Wijnands et al. (2015), compares the photon index of a power-law spectrum fitted to a transient LMXB in the very-faint regime ( erg/s), where black hole LMXBs have significantly harder spectra than most neutron star LMXBs. High extinction weakens the power of this test, but it may still be useful. We have plotted the low-extinction neutron star and black hole systems used by Wijnands et al. (2015) to introduce this test in Fig. 6, along with the measurements published for several high-extinction known neutron star systems (AX J1745.6-2901, GRS 1741.9-2853, SAX J1747.0-2853, KS 1741-293, GRO 1744-28, XMMU J1747161-281048, and XMM J174457-2850.3). All but the last of these systems conform to the trend of high photon indices for neutron star systems with between erg/s. Plotting the very-faint X-ray transients, however, we see that their photon indices align with those of black hole LMXBs. Although this is extremely interesting, we follow Wijnands et al. (2015) in pointing out that a number of neutron star LMXBs also show equally hard spectra in this range (e.g. IGR J18245-2452 in M28, Linares et al. 2014; EXO 1745-248, Rivera Sandoval et al. 2018; and XMM J174457-2850.3, Degenaar et al. 2014d). Although some of these are, or may be, transitional millisecond pulsars (Degenaar et al., 2014d), not all such hard-spectrum neutron star LMXBs are, and it seems unlikely that there would be so many transitional millisecond pulsars in active states in such a small region (given that only 2% of quiescent neutron star LMXBs appear as active-state transitional millisecond pulsars, Bahramian et al. 2020b).
For most neutron star transients elsewhere, we cannot constrain their history of fainter outbursts peaking at erg/s. Only for three nearby neutron stars accreting from low-mass companions are we able to rule out outbursts since 1996 above ergs s-1; the LMXB Cen X-4 (at 1.3 kpc), and the two transitional millisecond pulsars PSR J1023+0038 (Archibald et al., 2009) and PSR J12270-4859 (Roy et al., 2015), both at 1.4 kpc. It is unclear if these two transitional millisecond pulsars undergo regular, or even any, major accretion outbursts, as IGR 18245-2452 did. These three objects suggest that most neutron stars accreting from low-mass companions remain in a quiescent state for long periods.
Cen X-4’s X-ray properties are fully consistent with the nonthermal Galactic Center sources described by Mori et al. (2021); Cen X-4’s thermal emission would be completely absorbed by the at the Galactic Center, while its harder, variable power-law tail with photon index 1-1.7 and 2-10 keV (Cackett et al., 2010; Chakrabarty et al., 2014) resembles the properties of the Galactic Center sources. Incidentally, the same is true for many other qLMXBs, including some detected in globular clusters (e.g. EXO 1745-248 and Swift J174805.3-244637 in Terzan 5; Wijnands et al. 2005a; Bahramian et al. 2014). While it is true that most identified globular cluster qLMXBs that have not experienced outbursts lack significant hard spectral tails (Mori et al., 2022), this is likely a selection effect, as globular cluster qLMXBs without detected outbursts are generally selected by their extremely soft spectra (e.g. Heinke et al., 2003b; Guillot et al., 2011). The deepest cluster X-ray observations, of 47 Tuc, indeed identified two strong candidate qLMXBs with significant power-law tails (index 2, erg/s), and no recorded outbursts (Heinke et al., 2005). 47 Tuc is far enough from the ecliptic that it is never blocked from view for all-sky X-ray monitors, and at 4.5 kpc, 20 mCrab comes to erg/s, such that even relatively faint outbursts would have been observed.
Among the other transient neutron star LMXBs with long (10-year) periods between known outbursts, we can identify several more with substantial power-law components in their quiescent spectra; XTE J0929-314 (Wijnands et al., 2005b), SAX J1324.5-6313 (Cornelisse et al., 2002a; Bahramian & Heinke, 2015), XB 1733-30 (CX2 in Terzan 1; Cackett et al. 2006; Bozzo et al. 2020a, GRO J1744-28 (Doroshenko et al., 2020), and 1RXS J180408.9-342058 (Parikh et al., 2017b). XTE J0929-314 and probably XB 1733-30 are ultracompact systems, which would not be dynamically disrupted even in the Galactic Center region.
The XB-NEWS group has conducted frequent optical monitoring of Cen X-4 since 2008 without a proper outburst, although they did observe a slow optical brightening culminating with a failed outburst in late 2020, peaking at erg/s (Baglio et al., 2022); this serves to illustrate that the XB-NEWS optical monitoring (Lewis et al., 2008; Russell et al., 2019b) is effective at identifying outbursts. XB-NEWS monitors 50 systems weekly, including a number of other NS LMXB transients detectable in the optical (such as IGR J00291+59, XTE J2123-058, XTE J0929-314, XTE J1814-338, and HETE J1900.1-2455; Lewis et al. 2008). Careful analysis of the XB-NEWS archive may enable determination of how many fainter outbursts we are missing.
4 Conclusions
We have compiled a list of known neutron star transient LMXBs, from the lists of known bursters and accreting pulsars. We have searched the literature to produce a list of known outbursts from these systems, for each of which we record the date, approximate 2-10 keV peak flux and (inferred, given the best-estimate distance) luminosity, and one (or a few) relevant references. For a few objects, we have identified outbursts not identified in the literature, using public (generally all-sky) lightcurves. The purpose of this list of outbursts is to facilitate more in-depth analyses of the outburst records. Such studies could be done using only the information from these tables, in which case it is important that such investigators bear in mind the non-uniformity and limited sensitivity of these lists, or could use the information and references here as a jumping-off point for more detailed study of one or more objects.
We use the information assembled to make a few new conclusions. For instance, we consider arguments about the distance to 4U 1608-52, and identify periods of prolonged quiescence in AX J1754.2-2754, XMM J174457-2850.3, and XMMU J174716.1-281048 (aka IGR J17464-2811). We then look at the distribution of outbursts with time, and flux, to identify a rough empirical flux completeness limit for X-ray binary outbursts since 1996 of ergs cm-2 s-1 (though a number of outbursts have been identified below this limit, and some above have been missed).
Finally, we look at the distribution of the frequency of recorded outbursts from these transients. While four transients average roughly one to three outbursts per year, nearly half (40 of 87 systems) have zero or one recorded outburst between 1996 and 2023. It is certainly possible that many of these have shown fainter outbursts below our detection level during this period. However, relatively stringent limits on outbursts from nearby transient LMXBs, such as Cen X-4, PSR J1023+0038, and PSR J12270-4859, for several decades suggest that there are transient neutron star LMXBs with properties consistent with the intriguing Galactic Center X-ray binaries identified by Hailey et al. (2018).
Note added in proof: Ni et al. (2025) report a new X-ray transient detected by the Einstein Probe. Wu et al. (2025a) report a refined position, an eclipse, and identify two flares in its X-ray lightcurve as X-ray bursts, indicating that EP J171159.4-333253 is almost certainly a new neutron star X-ray transient.
Acknowledgements
We thank the referee for an exceptionally thoughtful and useful report. Swift/XRT lightcurves supplied by the UK Swift Science Centre at the University of Leicester (Evans et al., 2007). RXTE/ASM results provided by the ASM/RXTE teams at MIT and at the RXTE SOF and GOF at NASA’s GSFC. RXTE/PCA results from NASA Goddard Space Flight Center, https://asd.gsfc.nasa.gov/Craig.Markwardt//galscan/main.html. INTEGRAL/JEM-X lightcurves from Kuulkers et al. (2007a) using the OSA10.1 software, distributed by the ISDC (Courvoisier et al., 2003). MAXI results provided by RIKEN, JAXA and the MAXI team (Matsuoka et al., 2009).
This paper made very extensive use of the Astrophysical Data Service (https://ui.adsabs.harvard.edu), ArXiv (http://arxiv.org), and the Astronomer’s Telegram (http://www.astronomerstelegram.org). CH is supported by NSERC Discovery Grant RGPIN-2023-04264. JZ has been supported by a Mitacs Globalink Research Internship.
This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in the Chandra Data Collection (CDC) 346: https://doi.org/10.25574/cdc.346 (catalog doi:10.25574/cdc.346).
In Appendix A we give the full table (Table A) of data on neutron star LMXB outbursts, except for the frequent outbursters, which are given in Appendix B. Appendix C tabulates information on bursters without recorded outbursts, and Appendix D tabulates the neutron star X-ray binaries that we consider to be persistent systems (not further discussed in this paper).
Appendix A Outburst details
| NS XRB | Begin | End | Peak | Index | d | References | ||
| 2-10 keV | cm-2 | 2-10 keV | kpc | |||||
| IGR J00291+5934 | 1998/11 | 1998/11 | 4.5e-10 | - | - | 9.6e35 | 4.2 | Remillard (2004) |
| IGR J00291+5934 | 2001/9 | 2001/9 | 3.9e-10 | - | - | 8.2e35 | 4.2 | Remillard (2004) |
| IGR J00291+5934 | 2004/12 | 2004/12 | 6.1e-10 | 7e21 | 1.7 | 1.3e36 | 4.2 | Eckert et al. (2004); Galloway et al. (2005) |
| IGR J00291+5934 | 2008/8 | 2008/10 | 6.1e-10 | - | 1.7 | 1.3e36 | 4.2 | Chakrabarty et al. (2008); Lewis et al. (2010) |
| IGR J00291+5934 | 2015/8 | 2015/8 | 5.6e-10 | 1e22 | 1.7 | 1.2e36 | 4.2 | Cummings et al. (2015a); De Falco et al. (2017) |
| MAXI J0556-332 | 2011/1 | 2012/5 | 1.9e-09 | - | - | 4.4e38 | 43.6 | Matsumura et al. (2011); Sugizaki et al. (2013) |
| MAXI J0556-332 | 2012/10 | 2012/12 | 4e-10 | 3.2e20 | - | 9e37 | 43.6 | Sugizaki et al. (2012); Parikh et al. (2017c) |
| MAXI J0556-332 | 2016/1 | 2016/3 | 7.7e-10 | - | - | 1.8e38 | 43.6 | Negoro et al. (2016a); Russell et al. (2016) |
| MAXI J0556-332 | 2020/4 | 2020/11 | 8.6e-10 | - | - | 2.0e38 | 43.6 | Negoro et al. (2020a); Page et al. (2022) |
| EXO 0748-676 | 1984/7 | 2008/8 | 1.1e-9 | 1.2e21 | 1.9 | 7e36 | 7.4 | Parmar et al. (1986); Degenaar et al. (2009) |
| EXO 0748-676 | 2024/5 | 2025/3 | 3.0e-10 | 1.5e21 | 2 | 2.0e36 | 7.4 | D’Elia et al. (2024); Mihara et al. (2024); Degenaar et al. (2025) |
| GS 0836-429 | 1970/12 | 1972/1 | 1.1e-9 | 3.5e22 | 1.1e37 | 9.2 | Markert et al. (1977) | |
| GS 0836-429 | 1990/10 | 1991/5 | 3.8e-10 | 2.2e22 | 1.5 | 3.9e36 | 9.2 | Aoki et al. (1992); Belloni et al. (1993) |
| GS 0836-429 | 2003/1 | 2003/5 | 1.0e-9 | 2.7e22 | 1.5 | 1.0e37 | 9.2 | Rodriguez et al. (2003); Aranzana et al. (2016) |
| GS 0836-429 | 2003/9 | 2004/6 | 1.3e-9 | 2.7e22 | 1.5 | 1.3e37 | 9.2 | Chelovekov et al. (2005); Aranzana et al. (2016) |
| MAXI J0911-655 | 2016/2 | 2023/10 | 1.7e-10 | 2.7e21 | 1.27 | 1.9e36 | 9.5 | Serino et al. (2016b); Sanna et al. (2017b); Ng et al. (2021a) |
| XTE J0929-314 | 2002/4 | 2002/4 | 6.7e-10 | 7.6e20 | 1.7-2 | 2.9e36 | 6 | Remillard (2002); Galloway et al. (2002); Juett et al. (2003) |
| SAX J1324.5-6313 | 2015/8 | 2015/8 | 1.3e-10 | 1.5e22 | - | 1.0e36 | 6.2 | Negoro et al. (2015a); Cornelisse et al. (2002a) |
| MAXI J1421-613 | 2014/1 | 2014/1 | 1.9e-9 | 4.8e22 | 2.1 | 2e36 | 3 | Morooka et al. (2014); Serino et al. (2015); Nobukawa et al. (2020) |
| SRGA J144459.2-604207 | 2022/1 | 2022/1 | 4e-10 | - | - | 3.4e36 | 8.5 | Negoro et al. (2024) |
| SRGA J144459.2-604207 | 2023/12 | 2023/12 | 6e-10 | - | - | 5.2e36 | 8.5 | Negoro et al. (2024); Sguera & Sidoli (2024) |
| SRGA J144459.2-604207 | 2024/2 | 2024/3 | 2.7e-9 | 2.9e22 | 1.8 | 2.3e37 | 8.5 | Mereminskiy et al. (2024); Molkov et al. (2024); Ng et al. (2024) |
| Cen X-4 | 1969/07 | 1969/7 | 6e-7 | - | - | 1.2e38 | 1.3 | Evans et al. (1970); Kuulkers et al. (2009) |
| Cen X-4 | 1979/5 | 1979/5 | 9.6e-8 | 9e20 | - | 1.9e37 | 1.3 | Kaluzienski et al. (1980); Kuulkers et al. (2009) |
| MAXI J1621-501 | 2017/9 | 2019/1 | 3.1e-9 | 6.1e22 | 2.7 | 2.6e37 | 8.4 | Hashimoto et al. (2017); Gorgone et al. (2019); Chenevez et al. (2018) |
| MAXI J1647-227 | 2012/6 | 2012/6 | 9.2e-10 | 4.3e21 | 1.5 | 1.1e37 | 10 | Negoro et al. (2012); Onodera et al. (2014); Kennea et al. (2012) |
| MXB 1659-298 | 1976/10 | 1979/7 | 1.9e-9 | - | - | 2.3e37 | 10 | Lewin et al. (1976a); Cominsky et al. (1983) |
| MXB 1659-298 | 1999/9 | 1999/9 | 6.5e-10 | - | - | 7.8e36 | 10 | in ’t Zand et al. (1999a); Heise et al. (1999); Oosterbroek et al. (2001) |
| MXB 1659-298 | 2015/9 | 2015/9 | 5.5e-10 | 2.7e21 | - | 6.6e36 | 10 | Negoro et al. (2015a); Sharma et al. (2018) |
| IGR J16597-3704 | 2017/10 | 2017/10 | 2.2e-10 | 1.5e22 | 1.46 | 2.2e36 | 9.1 | Bozzo et al. (2017); Sanna et al. (2018) |
| XTE J1701-407 | 2008/6 | 2011/7 | 7.3e-10 | 3.4e22 | 2.0 | 3.3e36 | 6.2 | Markwardt et al. (2008c); Falanga et al. (2009); Linares et al. (2009a) |
| XTE J1701-407 | 2011/9 | 2011/9 | 3.8e-10 | 3e22 | 1.4 | 1.8e36 | 6.2 | Degenaar et al. (2011a) |
| XTE J1701-407 | 2016/8 | 2017/4 | 3.5e-10 | - | - | 1.6e36 | 6.2 | Swift/XRT |
| XTE J1701-407 | 2018/3 | 2018/3 | 3.5e-10 | - | - | 1.6e36 | 6.2 | Barthelmy et al. (2018),Swift/XRT |
| XTE J1701-407 | 2019/4 | 2019/4 | 4e-10 | 3.2e22 | 1.9 | 1.8e36 | 6.2 | Iwakiri et al. (2019) |
| XTE J1701-407 | 2020/1 | 2020/3 | 1.8e-10 | - | - | 8.2e35 | 6.2 | Onori et al. (2020); Palmer et al. (2020) |
| XTE J1701-462 | 2005/12 | 2007/8 | 4.3e-8 | 2.7e22 | 1.9 | 4e38 | 8.8 | Remillard et al. (2006); Lin et al. (2009b); Kennea et al. (2006) |
| XTE J1701-462 | 2022/9 | 2023/3 | 1.4e-8 | 3e22 | - | 1.3e38 | 8.8 | Iwakiri et al. (2022a); Cocchi et al. (2023),Swift/XRT |
| IGR J17062-6143 | 2005/12 | 2022/9* | 1.4e-10 | 2.3e21 | - | 8.7e35 | 7.3 | Churazov et al. (2007); Remillard & Levine (2008b); Bult et al. (2021a) |
| XTE J1709-267 | 1990/3 | 1990/8 | 4.3e-9 | - | - | 3.7e37 | 8.5 | Yamauchi & Koyama (1990); Yamauchi (2010) |
| XTE J1709-267 | 1997/1 | 1997/4 | 6.7e-9 | - | - | 5.8e37 | 8.5 | Marshall et al. (1997); Cocchi et al. (1998) |
| XTE J1709-267 | 2002/3 | 2002/4 | 7.3e-9 | 4.4e21 | - | 6.3e37 | 8.5 | Jonker et al. (2003) |
| XTE J1709-267 | 2004/5 | 2004/5 | 7.7e-10 | - | - | 6.6e36 | 8.5 | Markwardt et al. (2004); Jonker et al. (2004) |
| XTE J1709-267 | 2005/5 | 2005/5 | 2e-10 | - | - | 1.7e36 | 8.5 | RXTE PCA Scan |
| XTE J1709-267 | 2007/11 | 2008/9 | 1.9e-9 | - | - | 1.6e37 | 8.5 | Remillard (2007), RXTE PCA Scan |
| XTE J1709-267 | 2009/3 | 2009/5 | 2.5e-10 | - | - | 2.2e36 | 8.5 | RXTE PCA Scan |
| XTE J1709-267 | 2010/7 | 2010/7 | 1.9e-9 | - | - | 1.6e37 | 8.5 | Negoro et al. (2010a) |
| XTE J1709-267 | 2012/8 | 2012/8 | 3.7e-9 | - | - | 3.2e37 | 8.5 | Sanchez-Fernandez et al. (2012); Degenaar et al. (2013) |
| XTE J1709-267 | 2013/8 | 2013/8 | 1e-9 | - | - | 8.9e36 | 8.5 | Negoro et al. (2013a) |
| XTE J1709-267 | 2016/6 | 2016/6 | 4.5e-9 | - | - | 3.9e37 | 8.5 | Nakahira et al. (2016); Ludlam et al. (2017) |
| XTE J1709-267 | 2017/8 | 2017/8 | 3.9e-10 | - | - | 3.4e36 | 8.5 | Negoro et al. (2017a) |
| XTE J1709-267 | 2018/12 | 2019/1 | 2.2e-9 | - | - | 1.9e37 | 8.5 | MAXI |
| XTE J1709-267 | 2020/9 | 2020/12 | 2.2e-9 | - | - | 1.9e37 | 8.5 | Adachi et al. (2020) |
| XTE J1709-267 | 2022/6 | 2022/6 | 1.3e-9 | - | - | 1.1e37 | 8.5 | MAXI |
| XTE J1709-267 | 2023/10 | 2023/11 | 1.4e-9 | - | - | 1.2e37 | 8.5 | Negoro et al. (2023a),MAXI |
| XTE J1709-267 | 2025/5 | 2025/6* | 3.5e-10 | 3.9e21 | 2.0 | 3.0e36 | 8.5 | Wu et al. (2025b) |
| 2S 1711-339 | 1975/2 | 1976/9 | 3.2e-9 | 1.5e22 | 2.2 | 2.2e37 | 7.5 | Carpenter et al. (1977); Cornelisse et al. (2002b) |
| 2S 1711-339 | 1998/7 | 1999/5 | 1.1e-9 | - | - | 7.2e36 | 7.5 | Remillard (1998) |
| 2S 1711-339 | 2004/3 | 2004/3 | 4.5e-10 | - | - | 3.0e36 | 7.5 | Markwardt & Swank (2004a) |
| 2S 1711-339 | 2010/2 | 2010/2 | 4.3e-10 | - | - | 2.9e36 | 7.5 | Negoro et al. (2010b) |
| 2S 1711-339 | 2016/11 | 2016/11 | 1.1e-9 | - | - | 7.7e36 | 7.5 | Negoro et al. (2016c) |
| 1M 1716-315 | 1971/9 | 1984/6 | 1.1e-9 | 2.1e21 | 2 | 6.2e36 | 6.9 | Markert et al. (1976); Makishima et al. (1981a); Jonker et al. (2007) |
| IGR J17191-2821 | 2007/3 | 2007/5 | 1.8e-9 | 3.4e21 | 1.75 | 2.7e37 | 11 | Turler et al. (2007); Altamirano et al. (2010c); Markwardt et al. (2007) |
| IGR J17191-2821 | 2012/6 | 2012/6 | 9.3e-10 | - | - | 1.3e37 | 11 | Sugimoto et al. (2012a) |
| XTE J1723-376 | 1999/2 | 1999/3 | 1.9e-9 | 4e22 | - | 1.4e37 | 8 | Marshall & Markwardt (1999); Marshall et al. (1999) |
| 4U 1730-22 | 1972/8 | 1972/8 | 2.9e-9 | - | - | 1.7e37 | 6.9 | Cominsky et al. (1978); Chen et al. (1997) |
| 4U 1730-22 | 2021/6 | 2024/3 | 2.4e-9 | 3.9e21 | - | 1.4e37 | 6.9 | Kobayashi et al. (2021); Chen et al. (2023a); Bult et al. (2022a), MAXI, Swift/XRT |
| KS 1731-260 | 1988/10 | 2001/2 | 5.2e-9 | 1.1e22 | - | 3.0e37 | 7 | Syunyaev et al. (1990); Wijnands et al. (2001b) |
| XB 1733-30 | 1980/7 | 1999/4 | 3.3e-9 | 1.8e22 | 2.1 | 1.3e37 | 5.7 | Makishima et al. (1981b); Guainazzi et al. (1999) |
| XB 1733-30 | 2020/4 | 2025/5* | 2e-9 | 2.2e22 | 2 | 7.7e36 | 5.7 | Mereminskiy et al. (2020b); Negoro et al. (2023b); Homan et al. (2023) |
| Swift J1734.5-3027 | 2013/5 | 2013/6 | 3e-10 | - | - | 2e36 | 7.2 | La Parola et al. (2013); Bozzo et al. (2015a) |
| Swift J1734.5-3027 | 2013/9 | 2013/9 | 3e-10 | 1e22 | 1.7 | 1.9e36 | 7.2 | Kennea et al. (2013); Bozzo et al. (2015a) |
| Swift J1734.5-3027 | 2013/10 | 2014/4 | 3.6e-10 | 1e22 | 2.36 | 2.2e36 | 7.2 | Bozzo et al. (2015a) |
| IGR J17379-3747 | 2004/2 | 2004/2 | 1.3e-10 | - | - | 1.1e36 | 8.5 | Markwardt et al. (2008a); Chelovekov & Grebenev (2010) |
| IGR J17379-3747 | 2005/3 | 2005/3 | 2.2e-10 | - | - | 1.9e36 | 8.5 | Bult et al. (2019a) |
| IGR J17379-3747 | 2008/9 | 2008/9 | 2.7e-10 | - | - | 2.3e36 | 8.5 | Markwardt et al. (2008a); Krimm et al. (2008) |
| IGR J17379-3747 | 2014/2 | 2014/2 | 2.2e-10 | - | - | 1.9e36 | 8.5 | Negoro et al. (2018a); Bult et al. (2019a) |
| IGR J17379-3747 | 2018/3 | 2018/3 | 8.0e-10 | 9e21 | 1.88 | 6.9e36 | 8.5 | Negoro et al. (2018a); Bult et al. (2019a) |
| IGR J17379-3747 | 2020/9 | 2020/9 | 9.2e-11 | - | 2.1 | 8e35 | 8.5 | Mereminskiy et al. (2020a) |
| XTE J1739-285 | 1999/9 | 1999/9 | 5.6e-9 | 1.7e22 | - | 3.6e37 | 7.3 | Markwardt et al. (1999); Bult et al. (2021b) |
| XTE J1739-285 | 2001/5 | 2001/5 | 1.0e-9 | 1.7e22 | - | 6.4e36 | 7.3 | Kaaret et al. (2007),RXTE/PCA |
| XTE J1739-285 | 2003/10 | 2003/10 | 1.0e-9 | 1.7e22 | - | 6.4e36 | 7.3 | Kaaret et al. (2007),RXTE/PCA |
| XTE J1739-285 | 2005/8 | 2006/3 | 2.7e-9 | 1.7e22 | - | 1.7e37 | 7.3 | Bodaghee et al. (2005) |
| XTE J1739-285 | 2012/8 | 2013/3 | 1.9e-9 | 1.7e22 | - | 1.2e37 | 7.3 | Sanchez-Fernandez et al. (2012); Kuulkers et al. (2013a) |
| XTE J1739-285 | 2014/1 | 2014/12 | 4.8e-9 | 1.7e22 | - | 3.0e37 | 7.3 | MAXI |
| XTE J1739-285 | 2019/9 | 2020/9 | 4.5e-9 | 1.7e22 | - | 2.9e37 | 7.3 | Mereminskiy & Grebenev (2019); Negoro (2020); Beri et al. (2023), MAXI |
| XTE J1739-285 | 2021/5 | 2021/8 | 2e-9 | 1.7e22 | 2 | 1.3e37 | 7.3 | Swift/XRT, MAXI |
| XTE J1739-285 | 2022/12 | 2023/2 | 5.4e-10 | 1.7e22 | - | 3.4e36 | 7.3 | Kashyap et al. (2023),MAXI |
| GRS 1741.9-2853 | 1990/3 | 1990/4 | 1.5e-10 | 1.5e36 | 9.0 | Sunyaev (1990); Pavlinsky et al. (1994) | ||
| GRS 1741.9-2853 | 1994/9 | 1994/9 | 3.2e-12 | 1.1e23 | 2 | 3.1e34 | 9.0 | Sakano et al. (2002) |
| GRS 1741.9-2853 | 1996/9 | 1996/9 | 1.8e-10 | 1.1e23 | 2.36 | 1.7e36 | 9.0 | Sakano et al. (2002) |
| GRS 1741.9-2853 | 2000/10 | 2000/10 | 4.0e-10 | 9.7e22 | 1.8 | 3.9e36 | 9.0 | Muno et al. (2003) |
| GRS 1741.9-2853 | 2002/10 | 2002/10 | 1.2e-11 | 1.1e23 | 1.2e36 | 9.0 | Trap et al. (2009) | |
| GRS 1741.9-2853 | 2005/4 | 2005/6 | 1.1e-10 | - | - | 1.1e36 | 9.0 | Wijnands et al. (2005c, 2006b); Trap et al. (2009) |
| GRS 1741.9-2853 | 2006/9 | 2006/9 | 1.2e-11 | - | - | 1.2e35 | 9.0 | Degenaar & Wijnands (2009, 2010) |
| GRS 1741.9-2853 | 2007/3 | 2007/4 | 2.6e-10 | - | - | 2.5e36 | 9.0 | Kuulkers et al. (2007b); Degenaar & Wijnands (2009); Trap et al. (2009) |
| GRS 1741.9-2853 | 2009/10 | 2009/10 | 2.2e-9 | - | - | 2.1e37 | 9.0 | Chenevez et al. (2009a); Degenaar & Wijnands (2010) |
| GRS 1741.9-2853 | 2010/7 | 2010/7 | 2.6e-10 | - | - | 2.5e36 | 9.0 | Degenaar et al. (2014c) |
| GRS 1741.9-2853 | 2013/8 | 2013/8 | 4.3e-9 | - | - | 4.2e37 | 9.0 | Degenaar et al. (2014c) |
| GRS 1741.9-2853 | 2016/3 | 2016/3 | 8.7e-11 | - | - | 8.4e35 | 9.0 | Degenaar et al. (2016b) |
| GRS 1741.9-2853 | 2017/10 | 2017/10 | 1.5e-9 | - | - | 1.5e37 | 9.0 | Degenaar et al. (2017d) |
| GRS 1741.9-2853 | 2020/4 | 2020/10 | 1.9e-10 | - | - | 1.8e36 | 9.0 | Degenaar et al. (2020a); Pike et al. (2021c) |
| KS 1741-293 | 1989/8 | 1989/8 | 1.1e-9 | - | - | 5.0e36 | 6.2 | in’t Zand et al. (1991) |
| KS 1741-293 | 1998/3 | 1998/9 | 1.1e-9 | - | - | 5.0e36 | 6.2 | in ’t Zand et al. (1998c); Sakano et al. (2002) |
| KS 1741-293 | 2003/3 | 2003/3 | 3.3e-10 | - | - | 1.5e36 | 6.2 | De Cesare et al. (2007) |
| KS 1741-293 | 2004/3 | 2004/3 | 3.2e-10 | 3.5e23 | 2 | 1.5e36 | 6.2 | De Cesare et al. (2007) |
| KS 1741-293 | 2005/7 | 2005/9 | 2.5e-10 | 1.7e23 | 1.8 | 1.2e36 | 6.2 | Kuulkers et al. (2007a) |
| KS 1741-293 | 2007/9 | 2007/9 | 3.4e-11 | 1.7e23 | 1.8 | 1.6e35 | 6.2 | Degenaar et al. (2012b) |
| KS 1741-293 | 2008/5 | 2008/8 | 4.9e-10 | 1.7e23 | 1.8 | 2.3e36 | 6.2 | Degenaar et al. (2008, 2012b) |
| KS 1741-293 | 2010/2 | 2010/3 | 3.4e-10 | - | - | 1.6e36 | 6.2 | Chenevez et al. (2010a) |
| KS 1741-293 | 2011/9 | 2012/2 | 1.6e-8 | - | - | 7.3e37 | 6.2 | Linares et al. (2011a); Chenevez et al. (2011a, 2012) |
| KS 1741-293 | 2013/2 | 2013/4 | 5.7e-10 | 1.7e23 | 1.8 | 2.7e36 | 6.2 | Kuulkers et al. (2013b), Swift/XRT |
| KS 1741-293 | 2016/8 | 2016/8 | 3.4e-10 | - | - | 1.6e36 | 6.2 | Chenevez et al. (2016) |
| KS 1741-293 | 2017/5 | 2017/5 | 1.3e-09 | 1.7e23 | 1.8 | 6.0e36 | 6.2 | Bahramian et al. (2021) |
| KS 1741-293 | 2019/5 | 2019/5 | 7.5e-11 | - | - | 3.4e35 | 6.2 | Heinke et al. (2019a) |
| 1A 1744-361 | 1976/2 | 1976/4 | 1.1e-09 | - | - | 8.3e36 | 8 | Carpenter et al. (1977) |
| 1A 1744-361 | 1989/8 | 1989/8 | 1.3e-09 | - | - | 9.9e36 | 8 | in’t Zand (2004) |
| 1A 1744-361 | 2003/11 | 2003/11 | 2.6e-09 | - | - | 2.0e37 | 8 | Remillard et al. (2003) |
| 1A 1744-361 | 2004/4 | 2004/4 | 3.4e-10 | - | - | 2.6e36 | 8 | Markwardt & Swank (2004b) |
| 1A 1744-361 | 2005/7 | 2005/7 | 5.1e-10 | - | - | 3.9e36 | 8 | Swank & Markwardt (2005) |
| 1A 1744-361 | 2008/6 | 2008/6 | 2.2e-09 | - | - | 1.7e37 | 8 | Remillard & Levine (2008a) |
| 1A 1744-361 | 2009/11 | 2009/11 | 1.4e-09 | - | - | 1.1e37 | 8 | Remillard (2009) |
| 1A 1744-361 | 2010/3 | 2010/3 | 5.8e-10 | - | - | 4.4e36 | 8 | RXTE/PCA |
| 1A 1744-361 | 2011/9 | 2011/9 | 1.7e-09 | - | - | 1.3e37 | 8 | RXTE/PCA |
| 1A 1744-361 | 2013/8 | 2013/8 | 2.1e-09 | 5.1e21 | 1.5 | 1.6e37 | 8 | Bahramian et al. (2013b) |
| 1A 1744-361 | 2022/5 | 2022/8 | 2.4e-09 | - | - | 1.9e37 | 8 | Kobayashi et al. (2022); Hughes et al. (2022); Ng et al. (2023) |
| 1A 1744-361 | 2024/7 | 2025/2* | 5.9e-9 | 5.7e21 | 1.4 | 4.5e37 | 8 | Xu et al. (2024); Cheng et al. (2024) |
| IGR J17445-2747 | 2004/2 | 2004/3 | 7.4e-11 | - | 2.2 | 5.7e35 | 8 | Bird et al. (2006); Mereminskiy et al. (2017b) |
| IGR J17445-2747 | 2017/4 | 2017/4 | 9.2e-11 | 4.9e22 | 2.2 | 7.1e35 | 8 | Heinke et al. (2017); Mereminskiy et al. (2017b) |
| IGR J17445-2747 | 2019/5 | 2019/5 | 4.6e-12 | - | 2.2 | 3.6e34 | 8 | Heinke et al. (2019a) |
| IGR J17445-2747 | 2023/2 | 2023/3 | 1.5e-11 | 5.6e22 | 2 | 1.1e35 | 8 | Kashyap et al. (2023) |
| XMM J174457-2850.3 | 2001/9 | 2001/9 | 6.2e-12 | 5.9e22 | 1.9 | 3.1e34 | 6.5 | Sakano et al. (2002) |
| XMM J174457-2850.3 | 2005/6 | 2005/6 | 1.0e-10 | - | - | 5.2e35 | 6.5 | Wijnands et al. (2005c, 2006b) |
| XMM J174457-2850.3 | 2008/6 | 2008/7 | 2.5e-10 | 1.1e23 | 1.43 | 1.3e36 | 6.5 | Degenaar & Wijnands (2010); Degenaar et al. (2014d) |
| XMM J174457-2850.3 | 2009/9 | 2009/10 | 2.2e-11 | 7.5e22 | 2.3 | 1.1e35 | 6.5 | Degenaar & Wijnands (2010); Degenaar et al. (2014d) |
| XMM J174457-2850.3 | 2010/7 | 2010/8 | 3.8e-11 | 7.5e22 | 2.0 | 1.9e35 | 6.5 | Degenaar et al. (2014d, 2015) |
| XMM J174457-2850.3 | 2012/8 | 2012/8 | 1.9e-10 | - | - | 9.8e35 | 6.5 | Degenaar et al. (2014d, 2015) |
| XMM J174457-2850.3 | 2016/9 | 2016/10 | 3.3e-11 | 1.1e23 | 1.6 | 1.7e35 | 6.5 | Degenaar et al. (2016a) |
| AX J1745.6-2901 | 1993/10 | 1994/10 | 1.9e-10 | 2.8e23 | 2.4 | 1.5e36 | 8 | Maeda et al. (1996) |
| AX J1745.6-2901 | 1997/3 | 1997/3 | 2.9e-11 | 1.3e23 | 1.44 | 2.2e35 | 8 | Sakano et al. (2002) |
| AX J1745.6-2901 | 2006/2 | 2006/6 | 1.2e-10 | 2.4e23 | 2.4 | 9.2e35 | 8 | Degenaar & Wijnands (2010) |
| AX J1745.6-2901 | 2007/2 | 2008/9 | 8e-10 | 2.4e23 | 2.7 | 6.1e36 | 8 | Degenaar & Wijnands (2010) |
| AX J1745.6-2901 | 2010/6 | 2010/11 | 8e-11 | 2.3e23 | 3.4 | 6.1e35 | 8 | Degenaar et al. (2010) |
| AX J1745.6-2901 | 2013/7 | 2016/6 | 3.3e-10 | 2e23 | 2.6 | 6.7e36 | 8 | Degenaar et al. (2014c) |
| AX J1745.6-2901 | 2017/4 | 2017/4 | 1e-11 | 2e23 | 2 | 8e34 | 8 | Degenaar et al. (2017c) |
| AX J1745.6-2901 | 2017/10 | 2017/10 | 5.1e-11 | 3.4e23 | 1.9 | 4e35 | 8 | Degenaar et al. (2017e) |
| AX J1745.6-2901 | 2019/9 | 2020/2 | 3.6e-11 | 2e23 | 3.3 | 3e35 | 8 | Degenaar et al. (2019a, 2020b) |
| AX J1745.6-2901 | 2021/7 | 2021/7 | 1.1e-10 | 2.2e23 | 3.5 | 8.4e35 | 8 | Degenaar et al. (2021) |
| AX J1745.6-2901 | 2023/6 | 2025/5* | 3.4e-11 | 2.2e23 | 1.0 | 2.6e35 | 8 | Reynolds et al. (2023); Degenaar et al. (2025) |
| GRO J1744-28 | 1995/12 | 1996/5 | 3.5e-8 | 3.5e22 | 1.2 | 6.7e37 | 4 | Fishman et al. (1995); Giles et al. (1996) |
| GRO J1744-28 | 1996/12 | 1997/4 | 2.1e-8 | - | - | 4.0e37 | 4 | Kouveliotou et al. (1997); Woods et al. (1999) |
| GRO J1744-28 | 2014/1 | 2014/1 | 9.7e-9 | 8.6e22 | 0.99 | 1.9e37 | 4 | Negoro et al. (2014a); Degenaar et al. (2014a) |
| GRO J1744-28 | 2017/2 | 2017/2 | 4.4e-10 | 7e22 | 1.2 | 8.4e35 | 4 | Mereminskiy et al. (2017a); König et al. (2020); Sanna et al. (2017a) |
| GRO J1744-28 | 2021/6 | 2021/6 | 1.2e-9 | 9e22 | 1.1 | 2.3e36 | 4 | Mereminskiy et al. (2021a) |
| EXO 1745-248 | 2000/7 | 2000/8 | 1.4e-8 | - | - | 6e37 | 5.9 | Markwardt & Swank (2000); Heinke et al. (2003a) |
| EXO 1745-248 | 2011/10 | 2011/11 | 2.1e-9 | 1.2e22 | 2.25 | 8.8e36 | 5.9 | Altamirano et al. (2011a); Serino et al. (2012) |
| EXO 1745-248 | 2015/3 | 2015/6 | 2.0e-8 | 3.0e22 | 0.93 | 8.4e37 | 5.9 | Altamirano et al. (2015); Tetarenko et al. (2016a) |
| IGR J17480-2446 | 2010/10 | 2010/11 | 1.4e-8 | 1.2e22 | 2.1 | 5.1e37 | 5.9 | Bordas et al. (2010); Patruno et al. (2012); Miller et al. (2011) |
| Swift J174805.3-244637 | 2012/7 | 2012/8 | 1.1e-8 | - | 1.6 | 4.6e37 | 5.9 | Wijnands et al. (2012); Bahramian et al. (2014) |
| Swift J174805.3-244637 | 2023/2 | 2023/3 | 2.5e-9 | 2.2e22 | - | 1.0e37 | 5.9 | Negoro et al. (2023c); Heinke et al. (2023); Sanna et al. (2023a) |
| IGR J17473-2721 | 2005/5 | 2005/5 | 1.2e-9 | - | - | 4.3e36 | 5.5 | Grebenev et al. (2005b); Markwardt & Swank (2005a) |
| IGR J17473-2721 | 2008/3 | 2008/8 | 1.3e-8 | 3.8e22 | - | 4.7e37 | 5.5 | Del Monte et al. (2008); Altamirano et al. (2008b); Chenevez et al. (2011b) |
| XMMU J174716.1-281048 | 2003/3 | 2012/10 | 7.4e-12 | 9e22 | 2.25 | 6.3e34 | 8.4 | Sidoli & Mereghetti (2003); Degenaar & Wijnands (2007); Kaur et al. (2017) |
| XMMU J174716.1-281048 | 2021/4 | 2021/4 | 5.9e-12 | - | - | 5e34 | 8.4 | Swift/XRT |
| EXO 1747-214 | 1984/4 | 1985/4 | 1.6e-9 | 1.1e22 | 2.4 | 2.4e37 | 11 | Parmar et al. (1985); Tomsick et al. (2005) |
| SAX J1747.0-2853 | 1976/6 | 1976/6 | 7.9e-10 | - | - | 5.3e36 | 7.5 | Proctor et al. (1978); Grebenev et al. (2002) |
| SAX J1747.0-2853 | 1991/10 | 1991/10 | 4.2e-10 | 7e22 | 2.4 | 2.8e36 | 7.5 | Grebenev et al. (2002) |
| SAX J1747.0-2853 | 1998/3 | 1998/3 | 5e-10 | 7e22 | 2.4 | 3.4e36 | 7.5 | in ’t Zand et al. (1998a); Sidoli et al. (1998) |
| SAX J1747.0-2853 | 1999/3 | 1999/3 | 3.4e-9 | 8.8e22 | 2.4 | 2.3e37 | 7.5 | Werner et al. (2004) |
| SAX J1747.0-2853 | 2000/2 | 2000/9 | 1.1e-8 | 7e22 | 2 | 7.4e37 | 7.5 | Markwardt et al. (2000); Wijnands et al. (2002b); Werner et al. (2004) |
| SAX J1747.0-2853 | 2001/9 | 2001/9 | 5.4e-9 | 7e22 | 2 | 3.6e37 | 7.5 | Wijnands et al. (2002b); Werner et al. (2004) |
| SAX J1747.0-2853 | 2004/3 | 2004/3 | 6.4e-9 | - | 4.3e37 | 7.5 | Markwardt & Swank (2004a) | |
| SAX J1747.0-2853 | 2005/10 | 2006/9 | 3.9e-10 | 9.5e22 | 2 | 2.6e36 | 7.5 | Wijnands et al. (2005d); Degenaar et al. (2012b) |
| SAX J1747.0-2853 | 2007/10 | 2007/10 | 2.2e-9 | - | - | 1.5e37 | 7.5 | Brandt et al. (2007) |
| SAX J1747.0-2853 | 2009/3 | 2009/3 | 3.4e-11 | - | - | 2.3e35 | 7.5 | Chenevez et al. (2009b); Campana et al. (2009) |
| SAX J1747.0-2853 | 2011/2 | 2011/2 | 6e-9 | - | - | 4e37 | 7.5 | Kennea et al. (2011),Swift/XRT |
| SAX J1747.0-2853 | 2012/2 | 2012/2 | 7.1e-9 | - | - | 4.8e37 | 7.5 | Chenevez et al. (2012) |
| SAX J1747.0-2853 | 2013/4 | 2013/4 | 3.5e-9 | - | - | 2.4e37 | 7.5 | Negoro et al. (2013b) |
| SAX J1747.0-2853 | 2014/10 | 2014/10 | 4.1e-10 | - | - | 2.7e36 | 7.5 | Chenevez et al. (2014) |
| SAX J1747.0-2853 | 2016/6 | 2016/6 | 4.8e-9 | - | - | 3.2e37 | 7.5 | Clavel et al. (2016) |
| SAX J1747.0-2853 | 2017/4 | 2017/10 | 5.4e-10 | - | - | 4e36 | 7.5 | Heinke et al. (2017),Swift/XRT |
| SAX J1747.0-2853 | 2019/2 | 2019/9 | 1.3e-8 | - | - | 9e37 | 7.5 | Negoro et al. (2019); Mereminskiy et al. (2019),Swift/XRT |
| SAX J1747.0-2853 | 2022/4 | 2022/8 | 7e-9 | - | - | 5e37 | 7.5 | Iwakiri et al. (2022b); Kennea et al. (2022b),Swift/XRT |
| SAX J1747.0-2853 | 2023/5 | 2023/6 | 1.4e-9 | - | - | 9.4e36 | 7.5 | Negoro et al. (2023d); Li et al. (2023),MAXI,Swift/XRT |
| SAX J1748.9-2021? | 1971/12 | 1972/1 | 7.3e-9 | 5.3e22 | 2.1 | 6.3e37 | 8.5 | Markert et al. (1975); Forman et al. (1976a) |
| SAX J1748.9-2021 | 1998/8 | 1998/8 | 1.3e-9 | 8.2e21 | 1.44 | 1.1e37 | 8.5 | in ’t Zand et al. (1998b, 1999b) |
| SAX J1748.9-2021 | 2001/8 | 2001/11 | 4.3e-9 | - | - | 3.7e37 | 8.5 | in’t Zand et al. (2001) |
| SAX J1748.9-2021 | 2005/5 | 2005/7 | 5.4e-9 | - | - | 4.7e37 | 8.5 | Markwardt & Swank (2005b); Altamirano et al. (2008a) |
| SAX J1748.9-2021 | 2009/12 | 2010/1 | 5.4e-9 | - | - | 4.7e37 | 8.5 | Suzuki et al. (2009); Patruno et al. (2010) |
| SAX J1748.9-2021 | 2015/2 | 2015/4 | 4e-9 | 5.7e21 | 2.3 | 3.4e37 | 8.5 | Kuulkers et al. (2015); Pintore et al. (2016) |
| SAX J1748.9-2021 | 2017/10 | 2017/10 | 1.9e-10 | 5.1e21 | 1.6 | 1.6e36 | 8.5 | Negoro et al. (2017b); Pintore et al. (2018) |
| SAX J1748.9-2021 | 2021/11 | 2021/12 | 4.9e-10 | 6e22 | 1.8 | 4.2e36 | 8.5 | Negoro et al. (2021b); Pike et al. (2021a, b) |
| IGR J17494-3030 | 2012/3 | 2012/3 | 1.0e-10 | 1.8e22 | 1.9 | 7.7e35 | 8 | Boissay et al. (2012); Armas Padilla et al. (2013) |
| IGR J17494-3030 | 2020/10 | 2020/10 | 1.4e-10 | - | - | 1.0e36 | 8 | Ducci et al. (2020); Ng et al. (2021b) |
| Swift J1749.4-2807 | 2006/6 | 2006/6 | 9.5e-11 | 3e22 | 5e35 | 6.7 | Schady et al. (2006); Wijnands et al. (2009) | |
| Swift J1749.4-2807 | 2010/4 | 2010/4 | 5.2e-10 | - | - | 2.8e36 | 6.7 | Pavan et al. (2010); Ferrigno et al. (2011); Altamirano et al. (2011c) |
| Swift J1749.4-2807 | 2021/3 | 2021/3 | 3.1e-10 | 4.5e22 | - | 1.7e36 | 6.7 | Mereminskiy et al. (2021b); Marino et al. (2022) |
| IGR J17498-2921 | 2011/8 | 2011/9 | 5.1e-10 | 1.2e22 | 1.8 | 2.0e36 | 5.7 | Gibaud et al. (2011); Falanga et al. (2012); Linares et al. (2011b) |
| IGR J17498-2921 | 2023/4 | 2023/5 | 6.1e-10 | 1.2e22 | 1.8 | 2.4e36 | 5.7 | Grebenev et al. (2023); Sanna et al. (2023b), NICER |
| SAX J1750.8-2900 | 1997/3 | 1997/3 | 3.5e-9 | 6e22 | 2.4 | 1.9e37 | 6.79 | Bazzano et al. (1997); Natalucci et al. (1999) |
| SAX J1750.8-2900 | 2001/3 | 2001/3 | 3.5e-9 | - | - | 4.2e36 | 6.79 | Kaaret et al. (2002) |
| SAX J1750.8-2900 | 2008/3 | 2008/10 | 5.1e-9 | 2.3e22 | 1.2 | 2.8e37 | 6.79 | Markwardt & Swank (2008a); Linares et al. (2008) |
| SAX J1750.8-2900 | 2011/2 | 2011/2 | 3.6e-10 | 3.2e22 | 2 | 2.0e36 | 6.79 | Fiocchi et al. (2011); Natalucci et al. (2011) |
| SAX J1750.8-2900 | 2015/9 | 2016/3 | 2.4e-9 | 2.3e22 | 2 | 1.3e37 | 6.79 | Cummings et al. (2015b); Sanchez-Fernandez et al. (2015), IGR/JEM-X |
| SAX J1750.8-2900 | 2018/9 | 2018/9 | 4.6e-10 | - | - | 2.5e36 | 6.79 | Bozzo et al. (2018) |
| SAX J1750.8-2900 | 2019/4 | 2019/6 | 3.8e-10 | 3e22 | 2 | 2.1e36 | 6.79 | Heinke et al. (2019b, a) |
| XTE J1751-305 | 1998/6 | 1998/6 | 1.7e-9 | - | 1.8 | 1.5e37 | 8.5 | Markwardt et al. (2002b) |
| XTE J1751-305 | 2002/4 | 2002/4 | 1.1e-9 | - | 1.8 | 9.9e36 | 8.5 | Markwardt & Swank (2002); Markwardt et al. (2002b) |
| XTE J1751-305 | 2005/3 | 2005/3 | 1.8e-10 | - | - | 1.6e36 | 8.5 | Grebenev et al. (2005c); Swank et al. (2005) |
| XTE J1751-305 | 2007/4 | 2007/4 | 2.2e-10 | - | - | 1.9e36 | 8.5 | Markwardt & Swank (2007) |
| XTE J1751-305 | 2009/10 | 2009/10 | 3.2e-10 | 9.8e21 | - | 2.8e36 | 8.5 | Chenevez et al. (2009a); Miller et al. (2003b) |
| IGR J17511-3057 | 2009/9 | 2009/9 | 4.5e-10 | 5e21 | 1.2 | 2.6e36 | 6.9 | Baldovin et al. (2009); Bozzo et al. (2010) |
| IGR J17511-3057 | 2015/3 | 2015/4 | 5e-10 | 1.2e22 | 1.5 | 2.8e36 | 6.9 | Bozzo et al. (2015b); Papitto et al. (2016) |
| IGR J17511-3057 | 2025/2 | 2025/2* | 3e-10 | - | - | 2e36 | 6.9 | Sguera & Sidoli (2025); Sguera (2025) |
| SAX J1753.5-2349? | 2006/4 | 2006/4 | 2.4e-10 | - | - | 1.8e36 | 8 | Turler et al. (2006) |
| SAX J1753.5-2349 | 2008/10 | 2008/10 | 1.7e-10 | 1.9e22 | 2.1 | 1.3e36 | 8 | Markwardt et al. (2008b); Del Santo et al. (2010) |
| SAX J1753.5-2349 | 2010/3 | 2010/3 | 2.8e-10 | - | 2.1 | 2.2e36 | 8 | Chenevez et al. (2010b) |
| AX J1754.2-2754 | 1999/10 | 2008/5 | 1.1e-11 | 2.1e22 | 2.54 | 5.7e34 | 6.6 | Sakano et al. (2002); Bassa et al. (2008) |
| AX J1754.2-2754 | 2008/10 | 2012/5 | 1.1e-11 | 2.3e22 | 2.9 | 5.7e34 | 6.6 | Jonker & Keek (2008) |
| AX J1754.2-2754 | 2014/2 | 2015/3 | 3e-12 | 2.1e22 | 2.5 | 1.6e34 | 6.6 | Swift/XRT |
| AX J1754.2-2754 | 2015/8 | 2015/8 | 4.6e-12 | 2.1e22 | 2.5 | 2.4e34 | 6.6 | Swift/XRT |
| AX J1754.2-2754 | 2015/10 | 2017/3 | 1.3e-11 | 2.1e22 | 2.5 | 7e34 | 6.6 | Chenevez et al. (2017), Swift/XRT |
| Swift J1756.9-2508 | 2007/6 | 2007/6 | 3.7e-10 | 6.2e22 | 2.2 | 2.8e36 | 8 | Krimm et al. (2007b, a) |
| Swift J1756.9-2508 | 2009/8 | 2009/8 | 5.7e-10 | - | - | 4.4e36 | 8 | Markwardt et al. (2009); Patruno et al. (2009) |
| Swift J1756.9-2508 | 2018/4 | 2018/4 | 2.5e-10 | 4.3e22 | - | 1.9e36 | 8 | Mereminskiy et al. (2018); Rai & Paul (2019) |
| Swift J1756.9-2508 | 2019/6 | 2019/6 | 2.6e-10 | 4.2e22 | - | 2.0e36 | 8 | Sanna et al. (2019); Li et al. (2021) |
| IGR J17591-2342 | 2018/8 | 2018/10 | 3.2e-10 | 4.4e22 | 2.4 | 2.5e36 | 8 | Ducci et al. (2018); Gusinskaia et al. (2020) |
| IGR J17597-2201 | 2001/2 | 2008/3 | 4.6e-10 | 4.5e22 | 1.7 | 3.5e36 | 8 | Lutovinov et al. (2003); Markwardt & Swank (2003a); Bodaghee et al. (2012) |
| 2S 1803-245 | 1976/5 | 1976/5 | 2.2e-8 | 1.47e22 | - | 1.4e38 | 7.3 | Jernigan (1976); Cornelisse et al. (2007) |
| 2S 1803-245 | 1998/4 | 1998/4 | 9.1e-9 | - | - | 5.8e37 | 7.3 | Muller et al. (1998); Marshall et al. (1998a) |
| 1RXS J180408.9-342058 | 1990 | 1990 | 2.0e-12 | - | 2.3 | 2.4e34 | 10.0 | Voges et al. (1999) |
| 1RXS J180408.9-342058 | 2015/1 | 2015/6 | 2.5e-9 | 3.3e21 | - | 3.0e37 | 10.0 | Krimm et al. (2015); Degenaar et al. (2016c); Parikh et al. (2017b) |
| SAX J1806.5-2215 | 1996/3 | 1997/10 | 2.5e-10 | - | - | 1.9e36 | 8 | Cornelisse et al. (2002a) |
| SAX J1806.5-2215 | 2011/2 | 2017/7 | 1.5e-9 | 4.2e22 | 1.7 | 1.1e37 | 8 | Altamirano et al. (2011b); Del Santo et al. (2012); Shaw et al. (2017), Swift/XRT |
| MAXI J1807+132 | 2017/3 | 2017/3 | 2.2e-10 | - | - | 4e36 | 12.4 | Negoro et al. (2017c); Shidatsu et al. (2017) |
| MAXI J1807+132 | 2019/9 | 2019/9 | 1.9e-10 | 1.3e21 | 2 | 3.5e36 | 12.4 | Shidatsu et al. (2019); Albayati et al. (2021) |
| MAXI J1807+132 | 2023/7 | 2023/10 | 2.7e-10 | 1.3e21 | 2 | 4.9e36 | 12.4 | Saikia et al. (2023); Illiano et al. (2023a); Rout et al. (2025), Swift/XRT |
| SAX J1808.4-3658 | 1996/9 | 1996/10 | 2.2e-9 | - | - | 1.9e36 | 2.7 | in ’t Zand et al. (1998d); Revnivtsev (2003) |
| SAX J1808.4-3658 | 1998/4 | 1998/4 | 1.3e-9 | - | - | 1.1e36 | 2.7 | Marshall et al. (1998b); Chakrabarty et al. (1998); Bult & van der Klis (2015) |
| SAX J1808.4-3658 | 2000/1 | 2000/4 | 2.8e-10 | 1.3e21 | 2.16 | 2.4e35 | 2.7 | van der Klis et al. (2000); Wijnands et al. (2001a) |
| SAX J1808.4-3658 | 2002/10 | 2002/11 | 1.8e-9 | - | - | 1.6e36 | 2.7 | Markwardt et al. (2002a); Bult & van der Klis (2015) |
| SAX J1808.4-3658 | 2005/5 | 2005/7 | 1.3e-9 | - | - | 1.1e36 | 2.7 | Markwardt et al. (2005); Bult & van der Klis (2015) |
| SAX J1808.4-3658 | 2008/9 | 2008/11 | 1.2e-9 | - | - | 1.0e36 | 2.7 | Markwardt & Swank (2008b); Bult & van der Klis (2015) |
| SAX J1808.4-3658 | 2011/11 | 2011/11 | 1.7e-9 | - | - | 1.5e36 | 2.7 | Markwardt et al. (2011); Bult & van der Klis (2015) |
| SAX J1808.4-3658 | 2015/4 | 2015/4 | 1.2e-9 | 1e21 | 1.6 | 1.0e36 | 2.7 | Sanna et al. (2015a, b) |
| SAX J1808.4-3658 | 2019/8 | 2019/8 | 2.1e-10 | 2.1e21 | 2 | 1.8e35 | 2.7 | Goodwin et al. (2019b); Russell et al. (2019a); Bult et al. (2019b) |
| SAX J1808.4-3658 | 2022/8 | 2022/10 | 3.7e-10 | - | - | 3.2e35 | 2.7 | Sanna et al. (2022a); Imai et al. (2022); Illiano et al. (2023b) |
| XTE J1807-294 | 2003/2 | 2003/6 | 1.3e-9 | 5.6e21 | - | 9.6e36 | 8 | Markwardt et al. (2003); Falanga et al. (2005) |
| XTE J1810-189 | 2008/3 | 2008/6 | 1.1e-9 | 3.8e22 | 1.7 | 4.0e36 | 5.5 | Markwardt & Swank (2008c); Chelovekov & Grebenev (2009); Weng et al. (2015) |
| XTE J1810-189 | 2009/1 | 2009/3 | 1.9e-10 | 3.8e22 | 1.7 | 6.9e35 | 5.5 | Chelovekov & Grebenev (2009),RXTE/PCA |
| XTE J1810-189 | 2011/6 | 2011/6 | 2.9e-9 | 4.9e22 | 2.6 | 1.0e37 | 5.5 | Rangelov et al. (2014) |
| XTE J1810-189 | 2012/12 | 2013/1 | 8.6e-10 | - | - | 3.1e36 | 5.5 | Negoro et al. (2013c) |
| XTE J1810-189 | 2020/8 | 2020/11 | 5.2e-10 | 6e22 | 1.8 | 1.9e36 | 5.5 | Bozzo et al. (2020b, a); Manca et al. (2023) |
| XTE J1810-189 | 2023/2 | 2023/10 | 1e-9 | - | - | 3.6e36 | 5.5 | Kobayashi et al. (2023), MAXI |
| SAX J1810.8-2609 | 1998/3 | 1998/3 | 3.1e-10 | - | - | 1.2e36 | 5.7 | Ubertini et al. (1998); Natalucci et al. (2000) |
| SAX J1810.8-2609 | 2002/4 | 2002/4 | 1.0e-10 | 5e21 | 2 | 3.9e35 | 5.7 | RXTE/PCA,RXTE/ASM |
| SAX J1810.8-2609 | 2007/8 | 2007/10 | 8e-10 | 5e21 | 2 | 3e36 | 5.7 | Parsons et al. (2007); Degenaar et al. (2007); Degenaar & Wijnands (2013a) |
| SAX J1810.8-2609 | 2012/5 | 2012/5 | 5.2e-9 | - | - | 2.0e37 | 5.7 | Degenaar & Wijnands (2013a) |
| SAX J1810.8-2609 | 2018/4 | 2018/4 | 1.8e-9 | - | - | 7e36 | 5.7 | Negoro et al. (2018b) |
| SAX J1810.8-2609 | 2021/5 | 2021/10 | 2.1e-9 | - | - | 8.2e36 | 5.7 | Iwakiri et al. (2021); Hughes et al. (2023) |
| XTE J1812-182 | 2003/1 | 2003/3 | 1.9e-9 | 1.3e23 | 2.47 | 4.5e37 | 14 | Cackett et al. (2006) |
| XTE J1812-182 | 2008/8 | 2008/10 | 2.2e-9 | - | - | 5.1e37 | 14 | Markwardt et al. (2008d); Goodwin et al. (2019a) |
| XTE J1812-182 | 2020/3 | 2020/3 | 2.0e-9 | - | - | 4.7e37 | 14 | Chenevez et al. (2020) |
| XTE J1814-338 | 1984/9 | 1984/9 | 5e-11 | - | - | 3.8e35 | 8 | Wijnands & Reynolds (2003) |
| XTE J1814-338 | 2003/6 | 2003/7 | 3.7e-10 | 1.6e21 | 2.1 | 2.8e36 | 8 | Markwardt & Swank (2003b); Papitto et al. (2007); Krauss et al. (2005) |
| MAXI J1816-195 | 2022/6 | 2022/7 | 3.2e-9 | 2.4e22 | 2.6 | 1.5e37 | 6.3 | Negoro et al. (2022a); Kennea et al. (2022a); Bult et al. (2022b) |
| Swift J181723.1-164300 | 2017/8 | 2017/8 | 3.6e-10 | 8e22 | 1.5 | 2.1e36 | 7 | Barthelmy et al. (2017); Parikh et al. (2017a) |
| IGR J18245-2452 | 2013/3 | 2013/4 | 1.2e-9 | 4e21 | 1.26 | 4.3e36 | 5.5 | Eckert et al. (2013); Papitto et al. (2013); Linares et al. (2014) |
| SAX J1828.5-1037 | 1993/4 | 1993/4 | 7.2e-12 | 4.1e22 | 1.7 | 3.3e34 | 6.2 | Cornelisse et al. (2002b) |
| SAX J1828.5-1037 | 2001/3 | 2001/3 | 7.5e-12 | 5e22 | - | 6e34 | 6.2 | Hands et al. (2004) |
| SAX J1828.5-1037 | 2008/10 | 2008/11 | 1.5e-11 | 4.1e22 | 1.6 | 6.9e34 | 6.2 | Campana (2009); Degenaar & Wijnands (2008), Swift/XRT |
| SAX J1828.5-1037 | 2011/11 | 2011/11 | 2.9e-10 | 4.1e22 | 1.6 | 1.3e36 | 6.2 | Asada et al. (2011); Serino et al. (2016a), Swift/XRT |
| Swift J185003.2-005627 | 2011/5 | 2011/7 | 1.2e-10 | 1.1e22 | 1.6 | 2.0e35 | 3.7 | Beardmore et al. (2011); Degenaar et al. (2012a) |
| Swift J1858.6-0814 | 2018/10 | 2020/5 | 2.1e-9 | 1.8e23 | - | 4.1e37 | 12.8 | Krimm et al. (2018); Negoro et al. (2020b); Hare et al. (2020) |
| HETE J1900.1-2455 | 2005/6 | 2015/10 | 8e-10 | 2e21 | - | 2.1e36 | 4.7 | Vanderspek et al. (2005); Degenaar et al. (2017a); Kaaret et al. (2006) |
| 1H 1905+000 | 1974/11 | 1985/9 | 1.5e-10 | 1.9e21 | 2.1 | 1.0e36 | 7.5 | Seward et al. (1976); Jonker et al. (2006a) |
| Swift J1922.7-1716 | 2004/11 | 2006/7 | 1.3e-10 | 2.9e21 | 2.1 | 3.6e35 | 4.8 | Falanga et al. (2006); Degenaar et al. (2012a) |
| Swift J1922.7-1716 | 2011/7 | 2011/8 | 2.1e-10 | 2.7e21 | 2.0 | 5.8e35 | 4.8 | Nakahira et al. (2011); Degenaar et al. (2012a) |
| Swift J1922.7-1716 | 2024/2 | 2024/5 | 4.2e-10 | - | 2.2 | 1.2e36 | 4.8 | Nakajima et al. (2024),Swift/XRT |
| MAXI J1957+032 | 2015/5 | 2015/5 | 2.0e-10 | 1.7e21 | 2 | 6.0e35 | 5 | Negoro et al. (2015b); Mata Sánchez et al. (2017) |
| MAXI J1957+032 | 2015/10 | 2015/10 | 3.2e-10 | - | - | 9.5e35 | 5 | Sugimoto et al. (2015); Mata Sánchez et al. (2017) |
| MAXI J1957+032 | 2016/1 | 2016/1 | 4e-10 | - | - | 1.2e36 | 5 | Tanaka et al. (2016); Beri et al. (2019) |
| MAXI J1957+032 | 2016/9 | 2016/9 | 6.6e-10 | - | - | 2.0e36 | 5 | Negoro et al. (2016d); Mata Sánchez et al. (2017) |
| MAXI J1957+032 | 2022/6 | 2022/6 | 8e-10 | - | - | 2.4e36 | 5 | Negoro et al. (2022b); Sanna et al. (2022b) |
| MAXI J1957+032 | 2025/5 | 2025/5 | 9e-10 | - | - | 3e36 | 5 | Sun et al. (2025); Negoro et al. (2025); Illiano et al. (2025) |
| XTE J2123-058 | 1998/6 | 1998/8 | 1.7e-9 | 6e20 | 2 | 1.5e37 | 8.5 | Levine et al. (1998); Tomsick et al. (1999, 2004) |
| 4U 2129+47 | 1970 | 1983/9 | 4.8e-10 | 2e21 | - | 1.8e35 | 1.75 | Forman et al. (1978); Bozzo et al. (2007) |
Note. — Published outbursts of verified neutron star LMXBs. Max is highest reported flux in each outburst, transformed to 2-10 keV, generally using the power-law spectral index & noted. We assume 1 Crab is (2-10)= erg cm-2 s-1. Peak observed intrinsic in 2-10 keV, for the assumed distance (listed). A * at the end date indicates that we do not have evidence that the outburst ended–the last recorded observation found it active. Reference to an instrument (e.g. MAXI) indicates we have made use of that instrument’s archive to generate some information for that line.
Notes on individual sources, except the frequent outbursters (which are in Appendix B):
IGR J00291+5934; the distance estimate is 4.2 kpc, from a photospheric radius expansion (PRE) burst (De Falco et al., 2017).
MAXI J0556-332; the distance of 43.6 kpc is from modeling the quiescent X-ray spectrum Parikh et al. (2017a), and is supported by comparison of its outburst behaviour (Homan et al., 2014).
EXO 0748-676; Galloway et al. (2008) derive the distance of 7.4 kpc from a PRE burst.
GS 0836-429; also known as 4U 0836-429, 1M 0836-425, MX 0836-42. Aranzana et al. (2016) place a 9.2 kpc upper limit on the distance from a burst. Note that we use the value of cm-2 quoted by Chelovekov et al. (2005), which Aranzana et al. (2016) misquote as cm-2.
MAXI J0911-655; also known as Swift J0911.9-6452. The 9.5 kpc distance is the estimate to the host globular cluster, NGC 2808 (Watkins et al., 2015). Its long outburst appears to have ended around October 2023 (Heinke & Degenaar, 2023).
XTE J0929-314; Galloway et al. (2002) estimate a lower limit on the distance ( kpc) from binary evolution considerations.
SAX J1324.5-6313; the distance limit of kpc is from a burst (Cornelisse et al., 2002b).
MAXI J1421-613; Nobukawa et al. (2020) estimate the distance as 3 kpc from analysis of dust scattering echoes.
Cen X-4 or 4U 1456-32; Kaluzienski et al. (1980) estimate a distance of 1.3 kpc from the X-ray burst. We reference from Chakrabarty et al. (2014).
MAXI J1621-501; Chenevez et al. (2018) estimate 8.4 kpc from a PRE burst.
MAXI J1647-227; the 10 kpc rough distance estimate was made by Onodera et al. (2014).
MXB 1659-298 or MXB 1658-298, or MXB 1659-29, or H 1658-298; the 10 kpc distance estimate is from a PRE burst (Sharma et al., 2018).
IGR J16597-3704; the 9.1 kpc distance estimate is from the globular cluster NGC 6256 (Valenti et al., 2007).
XTE J1701-407; the distance of 6.2 kpc was estimated from a PRE burst (Falanga et al., 2009; Linares et al., 2009a). It is not clear whether XTE J1701-407 has ever reached full quiescence since 2016.
XTE J1701-462; Lin et al. (2009a) measure a distance of 8.8 kpc via PRE bursts. Swift/XRT observations in March 2023 showed its 2023 outburst declining below erg/s.
IGR J17062-6143; Keek et al. (2017) measure a distance of 7.3 kpc from a PRE burst. We use cm-2 from Degenaar et al. (2017b).
XTE J1709-267, or RX J1709.5-2639; Jonker et al. (2003) suggest a distance of 8.5 kpc if it is associated with the nearby globular cluster NGC 6293.
2S 1711-339 or 4U 1711-34; the distance limit of 7.5 kpc is from a burst (Cornelisse et al., 2002b).
1M 1716-315; also known as MXB 1715-32, XB 1715-321, or 1H 1715-321. The distance of 6.9 kpc is estimated from a PRE burst Jonker & Nelemans (2004).
IGR J17191-2821; the distance limit of 11 kpc is from a burst (Markwardt et al., 2007).
XTE J1723-376; the 8 kpc distance limit is from a burst (Marshall et al., 1999).
4U 1730-22; Bult et al. (2022a) measure the distance of 6.9 kpc from PRE bursts. The end of its second outburst is uncertain, but it was clearly in quiescence in a Swift/XRT observation in May 2025.
KS 1731-260; Muno et al. (2000) estimate the distance from PRE bursts.
XB 1733-30; this quasi-persistent transient in Terzan 1 is also known as SLX 1732-304, XB 1732-304, and 1RXS J173546.9-302859. We use the globular cluster distance estimate of 5.7 kpc from Baumgardt & Vasiliev (2021).
Swift J1734.5-3027 or IGR J17344-3023; Bozzo et al. (2015a) measured a distance of 7.2 kpc from a PRE burst.
IGR J17379-3747 or IGR J17380-3749; this is assumed to be located in the Galactic bulge, for a distance of roughly 8.5 kpc Chelovekov & Grebenev (2010).
XTE J1739-285; Galloway et al. (2008) find a distance constraint of 7.3 kpc from the brightest bursts, assuming a H-rich composition, which appears appropriate for the brightest bursts given their longer rise and decay times (Galloway et al., 2008).
GRS 1741.9-2853; Pike et al. (2021c) measure a distance of 9.0 kpc from a 2020 May PRE burst. (This is farther than the 7 kpc distance estimate inferred from prior X-ray bursts, which however did not show PRE; Barrière et al. (2015).) The proximity of this source to the Galactic Centre prevents us from using all-sky monitors to track its recurrences.
KS 1741-293; Chelovekov & Grebenev (2011) place an upper limit of 6.2 kpc from analyzing bursts. While this source has generally been seen not to exceed a few ergs s-1 (see e.g. Degenaar & Wijnands 2013b), we notice that the 2011-2012 outburst was reported by Chenevez et al. (2012) using INTEGRAL/JEM-X to reach a 3-10 keV flux of 415 mCrab, an order of magnitude brighter. We also identify a Swift/XRT observation on May 14, 2017 that reached 5.2 cts/s, or erg/s.
1A 1744-361, or A 1744-361; its proximity to the Galactic center has led many workers to assume an 8 kpc distance, though the only burst constraint is 11 kpc (Galloway et al., 2008). The RXTE/PCA Bulge scans show three likely small outbursts, at MJD 52020, 53281, and 54188, all reaching around ergs cm-2 s-1. Given the crowded environment, it would be useful to verify these signals. 1A 1744-361 is not clearly resolved by MAXI, so our sensitivity to its outbursts declines after 2011.
IGR J17445-2747; the proximity of this source to the Galactic Centre suggests an 8 kpc distance. Mereminskiy et al. (2017b) place limits of 5 and 12.3 kpc, from analysis of a burst.
XMM J174457-2850.3; Degenaar et al. (2014c) measure a distance of 6.5 kpc from a PRE burst. We use the long-term Swift XRT monitoring lightcurve (Degenaar et al., 2015) to check for additional outbursts (with signal/error 4, erg/s), and find only one (in 2016, Degenaar et al., 2016a) between 2012 and 2022. This appears to be a substantial change from its frequent outbursting behavior before 2012 (Degenaar et al., 2014d).
AX J1745.6-2901; the proximity to the Galactic Centre suggests an 8 kpc distance, which is supported by the peak fluxes of the brightest X-ray bursts (Degenaar & Wijnands, 2009). (Note this is not the same as 1A 1742-289, Kennea & Skinner 1996.)
GRO J1744-28; although the proximity to the Galactic Centre suggests a distance of 8 kpc, several methods of estimating the distance seem to converge on a value around 4 kpc (Sanna et al., 2017c; Gosling et al., 2007; Wang et al., 2007). This adjusts the of the brightening episode in Sept. 2008 to erg/s, below our outburst definition. The 2011 MAXI outburst initially identified as GRO 1744-28 (Suwa et al., 2011) was from SAX J1747.0-2853 (Kennea et al., 2011).
EXO 1745-248 (or XB 1745-248), IGR J17480-2446 (or Ter 5 X-2), Swift J17480.5-244637 (or Ter 5 X-3, or MAXI J1747-249): These are all located in the globular cluster Terzan 5. Additional outbursts have been seen in 1981, 1991 (Johnston et al., 1995), 2002 (Wijnands et al., 2002a), and 2004 (Markwardt & Swank, 2004a), but these cannot be confidently identified with one of the identified transients, which require arcsecond-scale localizations in this crowded cluster. We use the distance of 5.9 kpc from Valenti et al. (2007).
IGR J17473-2721 is also known as XTE J1747-274. The distance estimate is from PRE bursts (Chenevez et al., 2011b).
XMMU J174716.1-281048; also known as IGR J17464-2811. The distance estimate is an upper limit from a burst (Degenaar et al., 2011b). Del Santo et al. (2015) report that XMMU J174716.1-281048 was not detected by Swift in 2014 or 2015, but it was not clear if it had returned to full quiescence. Extensive Swift/XRT observations, and several archival XMM-Newton and Chandra observations, allow an accurate flux history of this source. We utilize the Leicester Swift/XRT lightcurves (Evans et al., 2007), archival Chandra observations from 2000, 2001, 2019, and 2022 (ObsIDs 1036, 2271, 21624, and 24222, the first 2 previously reported by Del Santo et al. 2007), and archival XMM-Newton observations from 2012, 2018, and 2020 (ObsIDs 0694641401, 0802410101, 0862471201), along with XMM flux values from 2003 and 2005 from Del Santo et al. (2007), to construct a longer lightcurve for XMMU J174716.1-281048 (Fig. 7. This reveals evidence that XMMU J174716.1-281048 declined significantly in late 2012 and had entered deep quiescence (2-10 keV erg/s) by 2018, most likely earlier. The Swift/XRT data also reveal that XMMU J174716.1-281048 experienced a second outburst in April 2021, but limit the period brighter than erg/s to between 2 weeks and 3 months. XMM and Chandra show that it was in deep quiescence in October 2020 and again by April 2022.
EXO 1747-214; the distance upper limit is from the EXOSAT burst (Tomsick et al., 2005).
SAX J1747.0-2853: we use the distance of 7.5 kpc from PRE bursts of Werner et al. (2004). We find a much smaller flux ( ergs cm-2 s-1, 2-10 keV unabsorbed) in Feb. 2011 than Kennea et al. (2011). The MAXI page notes that the apparent flare in MAXI data in early 2014 is from a different source.
SAX J1748.9-2021 and NGC 6440 X-2: these two accreting millisecond X-ray pulsars are both located in the globular cluster NGC 6440 X-2, at d=8.5 kpc (Harris, 1996). They can be distinguished by their spatial positions (using Chandra, Swift/XRT, or radio interferometric imaging), or by detection of X-ray pulsations. SAX J1748.9-2021 has been verified to be the object in outburst in all bright detected outbursts, except the 1971 outburst, but the similarity of that outburst to others verified to be SAX J1748.9-2021 suggest that it was responsible.
NGC 6440 X-2 showed numerous weak, short X-ray outbursts in 2009 and 2010, often with a 1-month recurrence time (Heinke et al., 2010)121212Note that Heinke et al. (2010) quote cm-2 in their Table 2, but actually used cm-2.. The RXTE PCA bulge scan reveals at least 7 faint outbursts during 2009-2011, but does not show other similar features except one in 2007, and possibly 4 in 1999. It is not certain whether NGC 6440 X-2 only produced outbursts in 2007 and 2009-2011, or whether we missed others (due perhaps to changes in the outburst peak flux). We list NGC 6440 X-2’s outbursts in a separate table (Table 2), to quote their MJD dates.
| NS XRB | Year/Month | MJD | Peak | Index | References | |
| NGC 6440 X-2 | 2007/4 | 54205 | 1.4e-10 | 2 | 1.2e36 | Heinke et al. (2010), RXTE/PCA |
| NGC 6440 X-2 | 2009/3 | 54910 | 1.9e-10 | 2 | 1.6e36 | Heinke et al. (2010), RXTE/PCA |
| NGC 6440 X-2 | 2009/5-6 | 54980 | 1e-10 | 2 | 8.6e35 | Heinke et al. (2010), RXTE/PCA |
| NGC 6440 X-2 | 2009/7 | 55040 | 1.7e-10 | 2 | 1.5e36 | Heinke & Budac (2009); Heinke et al. (2010), RXTE/PCA |
| NGC 6440 X-2 | 2009/8-9 | 55072 | 1.6e-10 | 1.79 | 1.4e36 | Heinke et al. (2009, 2010); Altamirano et al. (2010d), RXTE/PCA |
| NGC 6440 X-2 | 2009/10 | 55105 | 5.6e-11 | 1.79 | 4.8e35 | Heinke et al. (2010) |
| NGC 6440 X-2 | 2009/10-11 | 55132 | 8.1e-11 | 1.8 | 7e35 | Heinke et al. (2010) |
| NGC 6440 X-2 | 2010/3 | 55274 | 1.5e-10 | 1.9 | 1.3e36 | Altamirano et al. (2010a), RXTE/PCA |
| NGC 6440 X-2 | 2010/6 | 55359 | 2e-10 | - | 1.7e36 | Patruno et al. (2010); Patruno & D’Angelo (2013), RXTE/PCA |
| NGC 6440 X-2 | 2010/10 | 55473 | 1.8e-10 | - | 1.6e36 | Patruno & D’Angelo (2013), RXTE/PCA |
| NGC 6440 X-2 | 2011/1 | 55584 | 2e-10 | - | 1.7e36 | Patruno & D’Angelo (2013) |
| NGC 6440 X-2 | 2011/3 | 55639 | 2.8e-10 | - | 2.4e36 | RXTE/PCA |
| NGC 6440 X-2 | 2011/9 | 55819 | 2.2e-10 | - | 1.9e36 | RXTE/PCA |
| NGC 6440 X-2 | 2011/11 | 55871 | 2e-10 | - | 1.7e36 | Patruno & D’Angelo (2013) |
IGR J17494-3030; Armas Padilla et al. (2013) assume a distance of 8 kpc, due to its proximity to the Galactic Center.
Swift J1749.4-2807; Wijnands et al. (2009) calculate a distance of 6.7 kpc, but this is an upper limit as the burst is not clearly PRE.
IGR J17498-2921; Linares et al. (2011b) estimate a 7.6 kpc distance from a PRE burst, but Galloway et al. (2024) infer kpc from more detailed modeling, which we adopt. NICER observations of the 2023 outburst show a rapid drop to quiescence in May 2023.
SAX J1750.8-2900; Galloway et al. (2008) calculate a distance of 6.79 kpc from PRE bursts.
XTE J1751-305; Papitto et al. (2008) discuss various estimates for the distance, from 6.7 to 11.6 kpc; they and other authors use a default distance of 8.5 kpc, as it is close to the Galactic Center.
IGR J17511-3057; the distance upper limit is 6.9 kpc, from bursts (Altamirano et al., 2010b).
SAX J1753.5-2349; the distance is assumed to be 8 kpc. We include an INTEGRAL-detected outburst in 2006, designated IGR J17536-2339, which may or may not be this object (and thus give it a ? in the table).
AX J1754.2-2754; The distance estimate of 6.6 kpc, from an INTEGRAL burst reported by Chenevez et al. (2017), assumes a helium composition for the atmosphere.
After its initial ASCA detection in 1999, one short interval of quiescence has been reported (in 2007/2008, Bassa et al. 2008). However, the Swift/XRT lightcurve presented by Shaw et al. (2017) suggests multiple additional quiescent periods. We use Swift/XRT observations to identify additional quiescent episodes (with (2-10 keV) erg/s, or Swift/XRT count rates 0.04 cts/s), in August 2012, May-June 2015, Sept. 2015. Since Feb. 2017, it has been routinely observed by the Swift Bulge Survey (Bahramian et al., 2021), with only one detection in July 2019 at (2-10 keV) erg/s. We used a recent 4.2 ks Chandra observation (May 5, 2022, ObsID 24082; PI G. Ponti) to place a strong upper limit of (2-10 keV) erg/s on AX J1754.2-2754. We show the Swift/XRT data, the new and previously published Chandra data, and the XMM-Newton point of Armas Padilla et al. (2013) in Fig. 8. We conclude that AX J1754.2-2754 spends much of its time at luminosities at least 10-100 times lower than its peak.
Swift J1756.9-2508; the 8 kpc distance estimate is based on its proximity to the Galactic Center. Li et al. (2021) place a lower limit on the distance of 4 kpc.
IGR J17591-2342; Russell et al. (2018) argue that the implies a distance 6 kpc; since it is near the Galactic Centre, it is typically taken to be 8 kpc away.
IGR J17597-2201, or XTE J1759-220; the 8 kpc distance estimate assumes it is near the Galactic Center, which is consistent with the (unconstraining, kpc) X-ray burst constraint (Galloway et al., 2008).
2S 1803-245 is also known as XTE J1806-246, and SAX J1806.8-2435. Cornelisse et al. (2007) place an upper limit of 7.3 kpc, from a burst in 1998.
1RXS J180408.9-342058; the 10.0 kpc distance estimate is from PRE bursts (Marino et al., 2019).
SAX J1806.5-2215 showed a long outburst, which started in Feb. 2011 (Altamirano et al., 2011b). This was reported in Atels until 2015 (Sguera et al., 2015), but Swift/XRT observations indicate it continued until July 2017, and was no longer visible to Swift/XRT by January 2018. The 8 kpc upper limit comes from an analysis of bursts by (Cornelisse et al., 2002a).
MAXI J1807+132 has an upper limit on its distance of 12.4 kpc, from bursts (Albayati et al., 2021). Its 2023 outburst experienced two re-flares through September and October, visible in the Swift/XRT lightcurve; while it might rebrighten yet again, it appears to be in quiescence in Nov. 2023.
SAX J1808.4-3658; Galloway & Cumming (2006) estimated a distance of 3.50.1 kpc, using PRE bursts; this was revised to kpc by Galloway et al. (2024).
XTE J1807-294; the distance is assumed to be 8 kpc (based on its proximity to the Galactic Centre).
XTE J1810-189; Weng et al. (2015) estimate a distance of 5.5 kpc from a PRE burst.
SAX J1810.8-2609: the distance was constrained by an X-ray burst to 5.7 kpc by Fiocchi et al. (2009). We note several likely weak, single-datapoint outbursts in the RXTE/PCA lightcurve, that are not remarked on in the literature. At MJD 52384 we see a 10-sigma detection; also there is a 5-sigma RXTE/ASM detection at 52373. We include this outburst in our list. At MJD 53207, and 54989, we see 3-sigma possible outbursts, without RXTE/ASM detections; we leave these off our list.
XTE J1812-182: also known as XMMU J181227.8-181234. This may be the same as 1H1812-182, detected by HEAO-1 with (2-10 keV) erg/cm2/s in 1977/78 (Wood et al., 1984), but the errorbox is large. Goodwin et al. (2019a) estimated the distance as 142 kpc, from analysis of a train of several (non-PRE) bursts.
XTE J1814-338: Wijnands & Reynolds (2003) argued this is the same as EXMS B1810-337. The distance estimate of 8 kpc from X-ray bursts is from Strohmayer et al. (2003).
MAXI J1816-195; we use the 6.3 kpc distance constraint from bursts from Chen et al. (2022).
Swift J181723.1-164300; the 71 kpc distance comes from a burst (Parikh et al., 2017a).
IGR J18245-2452; this accreting, bursting transitional millisecond pulsar is in the globular cluster M28, at a distance of 5.5 kpc (Harris, 1996, 2010 revision). Although no other outbursts have been recorded from this cluster, a burst during quiescence has been seen (see Appendix C for AX J1824.5-2451).
SAX J1828.5-1037; the distance upper limit of 6.2 kpc is from a burst (Cornelisse et al., 2002b). The flux quoted for the 2008 outburst by Campana (2009) appears to be in error by a factor of ten.
Swift J185003.2-005627; the distance estimate upper limit of 3.7 kpc is from the single observed burst (Degenaar et al., 2012a).
Swift J1858.6–0814; Buisson et al. (2020) use PRE bursts to estimate a distance of 12.8 kpc, though the large and variable ( cm-2, Hare et al. 2020) implies systematic uncertainties in this measurement.
HETE J1900.1-2455; Galloway et al. (2008) compute a distance of 4.70.6 kpc from a PRE burst.
1H 1905+000 or 4U 1857+01, or 4U 1905+000; Jonker & Nelemans (2004) suggest either 7.5 or 10 kpc based on a PRE burst detected by Chevalier & Ilovaisky (1990), depending if the burst was H-rich or He-rich. As Chevalier & Ilovaisky (1990) identify a moderately long rise time (2.5 s) and decay time (=19 s), we suggest the burst was H-rich, and prefer the shorter 7.5 kpc distance.
Swift J1922.7-1716; the distance estimate of 4.8 kpc is from two PRE bursts (Degenaar et al., 2012a).
MAXI 1957+032, or IGR J19566+0326; the distance estimate of 52 kpc is from the hypothesized optical counterpart (which would require a triple system, Ravi 2017).
XTE J2123-058; the distance estimate of 8.5 kpc combines several constraints from bursts and studies of the optical counterpart in quiescence (Tomsick et al., 2004).
4U 2129+47 or XB 2129+47; Arnason et al. (2021) show that the Gaia distance to the star generally believed to be the counterpart is only 1.75 kpc. This is an accretion-disk corona source, so the estimate may be underestimated.
MAXI reported an outburst in December 2019 at 143 mCrab, or 1e35 erg/s (Kurihara et al., 2019), but it is unclear that this was a real detection; Swift observed within 2 days and only placed (unabsorbed) flux limits erg/s/cm2, or erg/s. We leave this possible outburst off our list.
Appendix B Frequently outbursting sources
4U 1608-52, or 4U 1608-522, X1608-52, QX Nor, in Table 3. We assume a distance of 3.20.3 kpc (Galloway et al., 2008). This is reasonably consistent with the estimate of Özel et al. (2016) of kpc. 4U 1608-52 has remained continually in low-level outburst for several years at several points in its history, most notably 1999-2000 (Wachter et al., 2002) and 2007-2009 (Šimon, 2020), but also probably in the 1970s (Šimon, 2020). We see 4U 1608-52 to be in a continuous outburst from July 2022 through July 2023, with a normal, bright, short outburst superimposed in November 2022.
| NS XRB | Year/Month | Peak | References | |
| 2-10 keV | 2-10 keV | |||
| 4U 1608-52 | 1970/4 | 2.0e-8 | 2.4e37 | Lochner & Roussel-Dupre (1994) |
| 4U 1608-52 | 1970/9 | 1.7e-8 | 2.1e37 | Lochner & Roussel-Dupre (1994) |
| 4U 1608-52 | 1971/6 | 1.5e-8 | 1.8e37 | Matilsky et al. (1972); Lochner & Roussel-Dupre (1994) |
| 4U 1608-52 | 1971/9 | 1.7e-8 | 2.1e37 | Lochner & Roussel-Dupre (1994); Tananbaum et al. (1976) |
| 4U 1608-52 | 1975/11 | 1.2e-8 | 1.5e37 | Kaluzienski et al. (1975) |
| 4U 1608-52 | 1977/7 | 2.4e-8 | 2.9e37 | Clark & Li (1977) |
| 4U 1608-52 | 1979/1 | 6.5e-9 | 7.9e36 | Kaluzienski & Holt (1979a) |
| 4U 1608-52 | 1979/4 | 1.3e-8 | 1.6e37 | Oda et al. (1979) |
| 4U 1608-52 | 1980/4 | 2.1e-10 | 2.6e35 | Oda (1980) |
| 4U 1608-52 | 1983/4 | 1.6e-8 | 2.0e37 | Mitsuda et al. (1984) |
| 4U 1608-52 | 1984/7 | 1.0e-9 | 1.2e36 | Gottwald et al. (1987) |
| 4U 1608-52 | 1986/3 | 5.8e-10 | 7.1e35 | Penninx et al. (1989) |
| 4U 1608-52 | 1991/6-12 | 9.7e-10 | 1.2e36 | Zhang et al. (1996) |
| 4U 1608-52 | 1996/3 | 7.1e-9 | 8.7e36 | Marshall & Angelini (1996); Méndez et al. (1998) |
| 4U 1608-52 | 1998/2-4 | 1.2e-8 | 1.5e37 | Cui & Swank (1998); Rutledge (2001) |
| 4U 1608-52 | 2000/1 | 4.4e-9 | 5.4e36 | Šimon (2004) |
| 4U 1608-52 | 2000/12 | 7.4e-9 | 9.0e36 | Šimon (2004) |
| 4U 1608-52 | 2001/11 | 1.5e-8 | 1.8e37 | Rutledge (2001) |
| 4U 1608-52 | 2002/7 | 1.7e-8 | 2.1e37 | Šimon (2004); Keek et al. (2008) |
| 4U 1608-52 | 2003/3 | 6.3e-9 | 7.7e36 | Miller et al. (2003a) |
| 4U 1608-52 | 2004/3 | 5.4e-9 | 6.6e36 | Remillard & Miller (2004) |
| 4U 1608-52 | 2004/7 | 2.9e-9 | 3.5e36 | Yan & Yu (2015) |
| 4U 1608-52 | 2005/3 | 2e-8 | 2.4e37 | Keek et al. (2008) |
| 4U 1608-52 | 2005/9 | 6.1e-9 | 7.4e36 | Asai et al. (2012) |
| 4U 1608-52 | 2006/7 | 1.9e-9 | 2.3e36 | Markwardt et al. (2006); Greco et al. (2006) |
| 4U 1608-52 | 2007/6-2009/10 | 7.8e-8 | 9.5e37 | Palmer et al. (2007); Russell et al. (2009) |
| 4U 1608-52 | 2010/2 | 6.3e-9 | 7.7e36 | Morii (2010); Armas Padilla et al. (2017) |
| 4U 1608-52 | 2011/3 | 1.7e-8 | 2.1e37 | Suguzaki et al. (2011); Asai et al. (2013) |
| 4U 1608-52 | 2012/10 | 1.2e-8 | 1.5e37 | Sugimoto et al. (2012b) |
| 4U 1608-52 | 2013/5 | 3e-9 | 3.7e36 | Zhang & Yu (2013) |
| 4U 1608-52 | 2014/1 | 4.1e-9 | 5.0e36 | Zhang & Yu (2014) |
| 4U 1608-52 | 2014/10 | 6.4e-9 | 7.8e36 | Negoro et al. (2014b) |
| 4U 1608-52 | 2016/8 | 5.5e-9 | 6.7e36 | Ono et al. (2016) |
| 4U 1608-52 | 2017/6 | 1.1e-8 | 1.3e37 | Jaisawal et al. (2019) |
| 4U 1608-52 | 2018/6 | 8.5e-9 | 1.0e37 | van den Eijnden et al. (2020) |
| 4U 1608-52 | 2019/3 | 5.7e-9 | 7.0e36 | MAXI |
| 4U 1608-52 | 2020/5 | 1.5e-8 | 1.8e37 | Güver et al. (2021) |
| 4U 1608-52 | 2021/8 | 2.8e-9 | 3.4e36 | MAXI |
| 4U 1608-52 | 2022/7-2023/7 | 2.1e-8 | 2.6e37 | Chen et al. (2023b, 2024), MAXI |
| 4U 1608-52 | 2024/2 | 4e-9 | 5e36 | MAXI |
| 4U 1608-52 | 2025/3 | 1.8e-8 | 2.2e37 | Wang et al. (2025),MAXI |
MXB 1730-335, aka Rapid Burster, 4U 1730-335, or MXB 1730-33. We use a distance of 8.00.3 kpc, as measured to the host globular cluster, Liller 1 (Baumgardt & Vasiliev, 2021). We include a table for its frequent outbursts (Table B). We note that some of these outbursts may be produced by the second known transient in this cluster, CXOU J173324.6-332321, which was discovered in low-level outburst ([0.5-10] erg/s) in April 2018 between two larger outbursts (Homan et al., 2018), but likely does not show Type II bursts as reported from many outbursts (e.g. Bagnoli et al., 2015). We quote MJD values for the highest observed countrate of each outburst. Where only intervals are given (e.g. in 1971), we estimate the midpoint of the interval. For spectral fitting, we assume a 1.6 keV blackbody, with cm-2 (Marshall et al., 2001; Mahasena et al., 2003). We use RXTE/PCA Bulge Scan values where available, then RXTE/ASM, MAXI, or Swift/XRT. For RXTE/ASM we multiply countrates by 3.4 to get 2-10 keV fluxes, for RXTE/PCA by 2.3. For MAXI we subtract 0.07 cts/cm-2/s (to account for confusion) from the 2-20 keV fluxes, and multiply by 1.1 to get 2-10 keV fluxes. For Swift/XRT data we multiply by 7.5 to get 2-10 keV fluxes, and for INTEGRAL/JEM-X 3-10 keV data we multiply by . These different flux estimates are generally in reasonable agreement where they overlap, though the RXTE/ASM data has some untrustworthy points when the Sun is very close. We identify ******* 52 outbursts that we do not see specifically identified elsewhere in the literature.
Masetti (2002) identified a change in the Rapid Burster’s outburst interval, from 200 to 100 days, and a factor of two decrease in peak luminosity, in the year 2000. We confirm that the outburst interval has remained generally between 100 and 170 days from 2000 to 2023, while the peak flux has generally averaged erg/s in this time frame. However, in 2023, the average peak flux decreased to erg/s. We see two intervals in the full record that are twice as long as nearby intervals, suggesting the RXTE instruments may have missed outbursts in Dec. 2001 and Dec. 2003 (due to proximity of the Sun). We also see two unusually short intervals, of 28 and 37 days, suggesting the outbursts on MJD 56491 and 57131 (each observed only by MAXI) appear out of sequence, and thus might be due to other transients.
| NS XRB | Year/Month | MJD | Peak | References | |
| 2-10 keV | 2-10 keV | ||||
| MXB 1730-335 | 1971/2 | 41028 | Grindlay & Gursky (1977) | ||
| MXB 1730-335 | 1972/4 | 41448 | Grindlay & Gursky (1977) | ||
| MXB 1730-335 | 1976/2 | 42840 | 1.7e-9 | 1.3e37 | Lewin et al. (1976c) |
| MXB 1730-335 | 1977/3-5 | 43266 | 2.2e-9 | 1.7e37 | White et al. (1978) |
| MXB 1730-335 | 1977/8 | 43408 | Grindlay & Gursky (1977) | ||
| MXB 1730-335 | 1978/3 | 43598 | Hoffman et al. (1978) | ||
| MXB 1730-335 | 1978/9 | 43796 | Lewin & Joss (1981) | ||
| MXB 1730-335 | 1979/2-3 | 43936 | Inoue et al. (1980); Lewin & Joss (1981) | ||
| MXB 1730-335 | 1979/8-9 | 44100 | Inoue et al. (1980) | ||
| MXB 1730-335 | 1983/4 | 45437 | Kaminker et al. (1990) | ||
| MXB 1730-335 | 1983/8 | 45564 | 1.2e-9 | 9e36 | Tanaka (1983); Kunieda et al. (1984); Barr et al. (1987) |
| MXB 1730-335 | 1984/7 | 45886 | 1.2e-9 | 9e36 | Makino (1984); Kawai et al. (1990) |
| MXB 1730-335 | 1985/8-9 | 46305 | 2.6e-9 | 2e37 | Stella et al. (1985, 1988) |
| MXB 1730-335 | 1988/8 | 47385 | Dotani et al. (1990) | ||
| MXB 1730-335 | 1989/1 | 47541 | Lewin et al. (1993) | ||
| MXB 1730-335 | 1996/4 | 50189 | 6.3e-9 | 4.8e37 | Levine et al. (1996b); Guerriero et al. (1999),RXTE/ASM |
| MXB 1730-335 | 1996/10 | 50387 | 7.7e-9 | 5.8e37 | Remillard (1996); Guerriero et al. (1999),RXTE/ASM |
| MXB 1730-335 | 1997/6 | 50625 | 9.3e-9 | 7.0e37 | Guerriero et al. (1997, 1999),RXTE/ASM |
| MXB 1730-335 | 1998/1 | 50843 | 6.8e-9 | 5.2e37 | Guerriero et al. (1998, 1999); Masetti et al. (2000),RXTE/ASM |
| MXB 1730-335 | 1998/8 | 51045 | 1.8e-8 | 1.4e38 | Fox et al. (1998); Masetti (2002); Mahasena et al. (2003),RXTE/ASM |
| MXB 1730-335 | 1999/3 | 51252 | 6.9e-9 | 5.3e37 | Masetti (2002); Mahasena et al. (2003),RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 1999/10 | 51452 | 6.8e-9 | 5.2e37 | Fox & Lewin (1999); Masetti (2002); Marshall et al. (2001),RXTE/PCA,ASM |
| MXB 1730-335 | 2000/3 | 51633 | 2.2e-9 | 1.6e37 | Fox & Lewin (2000); Masetti (2002),RXTE/ASM, RXTE/PCA |
| MXB 1730-335 | 2000/7 | 51747 | 3.3e-9 | 2.5e37 | Masetti (2002); Homer et al. (2001),RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 2000/10 | 51849 | 2.7e-9 | 2.0e37 | Masetti (2002),RXTE/ASM, RXTE/PCA |
| MXB 1730-335 | 2001/2 | 51954 | 5.4e-9 | 4.1e37 | Masetti (2002),RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 2001/5 | 52052 | 3.9e-9 | 3.0e37 | Masetti (2002),RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 2001/9 | 52154 | 3.8e-9 | 2.9e37 | Masetti (2002),RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 2002/4 | 52387 | 2.7e-9 | 2.1e37 | Simon (2008),RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 2002/8 | 52510 | 1.8e-9 | 1.4e37 | Simon (2008),RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 2003/2 | 52686 | 3.6e-9 | 2.7e37 | Falanga et al. (2004); Simon (2008),RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 2003/5 | 52785 | 2.7e-9 | 2.0e37 | Simon (2008),RXTE/ASM, RXTE/PCA |
| MXB 1730-335 | 2003/9 | 52886 | 1.7e-9 | 1.3e37 | Simon (2008),RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 2004/4 | 53102 | 5.0e-9 | 3.8e37 | Simon (2008),RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 2004/8 | 53237 | 4.5e-9 | 3.4e37 | RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 2004/11 | 53330 | 8.0e-9 | 6.1e37 | RXTE/ASM |
| MXB 1730-335 | 2005/4 | 53475 | 5.0e-9 | 3.8e37 | RXTE/PCA, RXTE/ASM, IGR/JEM-X |
| MXB 1730-335 | 2005/8 | 53602 | 3.0e-9 | 2.3e37 | Kretschmar et al. (2005), RXTE/PCA, RXTE/ASM, IGR/JEM-X |
| MXB 1730-335 | 2006/1 | 53749 | 5.6e-9 | 4.3e37 | RXTE/ASM |
| MXB 1730-335 | 2006/7 | 53918 | 3.7e-9 | 2.8e37 | Bagnoli et al. (2015),RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 2006/12 | 54078 | 5.6e-9 | 4.2e37 | RXTE/ASM |
| MXB 1730-335 | 2007/6 | 54265 | 4.8e-9 | 3.6e37 | RXTE/ASM |
| MXB 1730-335 | 2007/11 | 54430 | 5.0e-9 | 3.8e37 | RXTE/ASM |
| MXB 1730-335 | 2008/2 | 54546 | 2.2e-9 | 1.7e37 | Kuulkers et al. (2008), RXTE/PCA, RXTE/ASM, IGR/JEM-X |
| MXB 1730-335 | 2008/6 | 54620 | 3.1e-9 | 2.4e37 | RXTE/PCA, RXTE/ASM |
| MXB 1730-335 | 2008/9 | 54727 | 2.9e-9 | 2.2e37 | RXTE/ASM, RXTE/PCA, IGR/JEM-X |
| MXB 1730-335 | 2009/2 | 54876 | 1.6e-9 | 1.1e37 | Sala et al. (2012), RXTE/ASM, RXTE/PCA, Swift/XRT, JEM-X |
| MXB 1730-335 | 2009/5 | 54976 | 3.2e-9 | 2.4e37 | RXTE/ASM, RXTE/PCA |
| MXB 1730-335 | 2009/8 | 55074 | 4.9e-9 | 3.7e37 | RXTE/ASM, RXTE/PCA, MAXI, IGR/JEM-X |
| MXB 1730-335 | 2010/2 | 55241 | 4.3e-9 | 3.3e37 | RXTE/PCA, RXTE/ASM, MAXI, IGR/JEM-X |
| MXB 1730-335 | 2010/7 | 55395 | 4.7e-9 | 3.6e37 | RXTE/ASM, MAXI, RXTE/PCA |
| MXB 1730-335 | 2011/1 | 55562 | 3.1e-9 | 2.4e37 | MAXI |
| MXB 1730-335 | 2011/6 | 55723 | 2.3e-9 | 1.8e37 | MAXI, RXTE/PCA |
| MXB 1730-335 | 2011/9 | 55811 | 3.6e-9 | 2.8e37 | Chenevez et al. (2011a), RXTE/PCA, MAXI, Swift/XRT, IGR/JEM-X |
| MXB 1730-335 | 2011/12 | 55902 | 1.9e-9 | 1.4e37 | MAXI |
| MXB 1730-335 | 2012/4 | 56019 | 1.1e-9 | 8.4e36 | MAXI, IGR/JEM-X |
| MXB 1730-335 | 2012/7 | 56136 | 4.1e-9 | 3.1e37 | MAXI |
| MXB 1730-335 | 2012/11 | 56256 | 2.8e-9 | 2.1e37 | MAXI |
| MXB 1730-335 | 2013/3 | 56352 | 3.3e-9 | 2.5e37 | Chenevez et al. (2013a), MAXI, IGR/JEM-X |
| MXB 1730-335 | 2013/6 | 56463 | 2.6e-9 | 2.0e37 | MAXI |
| MXB 1730-335 | 2013/7 | 56491 | 3.7e-9 | 2.8e37 | MAXI |
| MXB 1730-335 | 2013/10 | 56586 | 2.6e-9 | 2.0e37 | Chenevez & Sanchez-Fernandez (2013), IGR/JEM-X, MAXI |
| MXB 1730-335 | 2014/3 | 56730 | 2.0e-9 | 1.5e37 | MAXI, IGR/JEM-X |
| MXB 1730-335 | 2014/9 | 56896 | 5.2e-9 | 3.9e37 | IGR/JEM-X, MAXI |
| MXB 1730-335 | 2015/1 | 57051 | 5.1e-9 | 3.9e37 | MAXI, IGR/JEM-X |
| MXB 1730-335 | 2015/4 | 57131 | 1.5e-9 | 1.1e37 | MAXI |
| MXB 1730-335 | 2015/5 | 57168 | 1.8e-9 | 1.3e37 | Swift/XRT, MAXI |
| MXB 1730-335 | 2015/10 | 57298 | 3.6e-9 | 2.8e37 | van den Eijnden et al. (2017),MAXI, IGR/JEM-X, Swift/XRT |
| MXB 1730-335 | 2016/2 | 57422 | 3.5e-9 | 2.7e37 | MAXI, IGR/JEM-X |
| MXB 1730-335 | 2016/6 | 57562 | 3.2e-9 | 2.4e37 | MAXI |
| MXB 1730-335 | 2016/12 | 57732 | 2.7e-9 | 2.1e37 | MAXI |
| MXB 1730-335 | 2017/3 | 57833 | 3.3e-9 | 2.5e37 | MAXI, IGR/JEM-X |
| MXB 1730-335 | 2017/8 | 57975 | 2.8e-9 | 2.1e37 | Chen et al. (2021), IGR/JEM-X, MAXI, Swift/XRT |
| MXB 1730-335 | 2017/11 | 58071 | 1.9e-9 | 1.4e37 | MAXI |
| MXB 1730-335? | 2018/3 | 58178 | 3.0e-9 | 2.2e37 | Homan et al. (2018), MAXI, IGR/JEM-X |
| MXB 1730-335 | 2018/6 | 58292 | 1.7e-9 | 1.3e37 | Bahramian et al. (2018), MAXI,Swift/XRT |
| MXB 1730-335 | 2018/9 | 58389 | 2.7e-9 | 2.0e37 | MAXI |
| MXB 1730-335 | 2019/1 | 58490 | 1.4e-9 | 1.0e37 | MAXI |
| MXB 1730-335 | 2019/3 | 58571 | 2.3e-9 | 1.8e37 | MAXI, Swift/XRT |
| MXB 1730-335 | 2019/7 | 58677 | 2.2e-9 | 1.7e37 | MAXI |
| MXB 1730-335 | 2019/11 | 58794 | 2.3e-9 | 1.7e37 | MAXI |
| MXB 1730-335 | 2020/3 | 58920 | 3.5e-9 | 2.7e37 | van den Eijnden et al. (2024),Swift/XRT, MAXI |
| MXB 1730-335 | 2020/8 | 59083 | 4.8e-9 | 3.6e37 | MAXI |
| MXB 1730-335 | 2021/2 | 59247 | 3.3e-9 | 2.5e37 | MAXI |
| MXB 1730-335 | 2021/8 | 59429 | 6.3e-10 | 4.8e36 | MAXI |
| MXB 1730-335 | 2021/12 | 59553 | 2.0e-9 | 1.6e37 | MAXI |
| MXB 1730-335 | 2022/5 | 59702 | 1.8e-9 | 1.4e37 | MAXI |
| MXB 1730-335 | 2022/10 | 59867 | 1.3e-9 | 1.0e37 | MAXI |
| MXB 1730-335 | 2023/1 | 59973 | 1.8e-9 | 1.4e37 | MAXI |
| MXB 1730-335 | 2023/6 | 60105 | 9.9e-10 | 7.5e36 | MAXI |
| MXB 1730-335 | 2023/8 | 60185 | 1.1e-9 | 8.6e36 | MAXI |
| MXB 1730-335 | 2024/1 | 60329 | 1.9e-9 | 1.4e37 | MAXI |
| MXB 1730-335 | 2024/6 | 60490 | 2.6e-9 | 2.0e37 | MAXI |
| MXB 1730-335 | 2024/10 | 60594 | 2.1e-9 | 1.6e37 | MAXI |
| MXB 1730-335 | 2025/1 | 60689 | 1.3e-9 | 9.5e36 | MAXI |
| MXB 1730-335 | 2025/4 | 60791 | 1.0e-9 | 7.5e36 | MAXI |
Note. — Outbursts of MXB 1730-335, the Rapid Burster in the globular cluster Liller 1. in ergs cm-2 s-1, in ergs s-1. Distance of 8 kpc assumed. Reference to an instrument (e.g. MAXI) indicates we have made use of that instrument’s archive to generate some information for that line.
GRS 1747-312; We use the distance estimate for the globular cluster Terzan 6 of 7.3 kpc from Baumgardt & Vasiliev (2021). GRS 1747-312’s outbursts have not been cataloged since 2008, so we turn to all-sky monitors to produce a full list (Table B). We assume cm-2 and a photon index of 1.0 (in’t Zand et al., 2000) to transform countrates to fluxes. For RXTE/ASM we multiply countrates by 3.6 to get 2-10 keV fluxes, for RXTE/PCA by 2.1. For Swift/XRT data we multiply by 9.4 to get 2-10 keV fluxes, for MAXI by , and for INTEGRAL/JEM-X by . We average RXTE/ASM data over a 3-day window (weighting by the error), to reduce the noise. We identify outbursts in the RXTE/ASM, RXTE/PCA, INTEGRAL/JEM-X, and MAXI lightcurves when we see at least two detections (among all the lightcurves) with the countrate exceeding 3 times the error above the average quiescent rate of 0.5 ct/s (see Fig. 10). This generally agrees with the identifications in in’t Zand et al. (2003) where they overlap, though the ASM data is noisy in this crowded region. We are also in good agreement with the identifications of Šimon (2009), which use a more complicated smoothing algorithm to construct their lightcurves. We note that the last two outbursts detected by the RXTE/ASM (late 2010 to early 2011) are especially noisy; we do not use these data, though we plot them in Fig. 10. The outbursts average every 138 days, with the average varying between 120 and 150 days over time, consistent with previous determinations (in’t Zand et al., 2003; Šimon, 2009). Peak fluxes average erg/s.
GRS 1747-312 is known to show regular eclipses (in’t Zand et al., 2000), which makes it possible to verify that an outburst from Terzan 6 is produced by this transient, given an accurate orbital ephemeris. Saji et al. (2016) noted a burst, and an absence of a predicted X-ray eclipse, during a Suzaku observation in Sept. 2009, when GRS 1747-312 was not detected in RXTE/PCA monitoring.131313van den Berg et al. (2024) point out that a spike is visible in RXTE/PCA scans on MJD 55088, two days earlier, but that pulsations in the scan reveal that this spike was produced by the known accreting millisecond pulsar IGR J17511–3057, then in outburst. More recently, van den Berg et al. (2024) have shown conclusive evidence that a second eclipsing, bursting X-ray binary, Terzan 6 X2, is located 0.7” from GRS 1747-312. We label the 2009 detection by Suzaku as ”Ter 6 X2” in Table B.
| NS XRB | Year/Month | E | MJD | Peak | References | |
| 2-10 keV | 2-10 keV | |||||
| GRS 1747-312 | 1990/9 | 48143 | 8.8e-10 | 5.5e36 | Pavlinsky et al. (1992); in’t Zand et al. (2000) | |
| GRS 1747-312 | 1996/3 | 0 | 50157 | 1.3e-9 | 8.2e36 | in’t Zand et al. (2003); Šimon (2009), RXTE/ASM |
| GRS 1747-312 | 1996/8 | 1 | 50318 | 1.1e-9 | 6.8e36 | in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM |
| GRS 1747-312 | 1997/1 | 2 | 50475 | 9.4e-10 | 6.0e36 | in’t Zand et al. (2003); Šimon (2009), RXTE/ASM |
| GRS 1747-312 | 1997/5 | 3 | 50598 | 1.2e-9 | 7.4e36 | in’t Zand et al. (2003); Šimon (2009), RXTE/ASM |
| GRS 1747-312 | 1997/10 | 4 | 50714 | 1.1e-9 | 7.2e36 | in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM |
| GRS 1747-312 | 1998/2 | 5 | 50867 | 1.2e-9 | 7.5e36 | in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM |
| GRS 1747-312 | 1998/6 | 6 | 50986 | 3.9e-10 | 2.5e36 | in’t Zand et al. (2003); Šimon (2009), RXTE/ASM |
| GRS 1747-312* | 1998/9 | 7 | 51057 | 9.1e-10 | 5.8e36 | in’t Zand et al. (2000, 2003); Šimon (2009), RXTE/ASM |
| GRS 1747-312 | 1999/1 | (8) | 51189 | 1.7e-9 | 1.1e37 | Šimon (2009), RXTE/ASM |
| GRS 1747-312 | 1999/5 | 9 | 51308 | 1.0e-9 | 6.3e36 | in’t Zand et al. (2000, 2003); Šimon (2009), PCA, RXTE/ASM |
| GRS 1747-312* | 1999/10 | 10 | 51451 | 9.9e-10 | 6.3e36 | in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312* | 2000/2 | 11 | 51585 | 1.3e-9 | 8.2e36 | in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312* | 2000/8 | 12 | 51759 | 1.1e-9 | 6.9e36 | in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312* | 2001/5 | 14 | 52063 | 9.5e-10 | 6.0e36 | in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312* | 2001/10 | 15 | 52207 | 1.0e-9 | 6.7e36 | in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312* | 2002/2 | 16 | 52337 | 8.8e-10 | 5.6e36 | in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312* | 2002/7 | 17 | 52470 | 9.6e-10 | 6.1e36 | in’t Zand et al. (2003); Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312 | 2003/3 | 19 | 52713 | 8.7e-10 | 5.5e36 | Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312 | 2004/3 | 22 | 53100 | 1.0e-9 | 6.5e36 | Markwardt & Swank (2004a); Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312 | 2004/8 | 23 | 53233 | 1.1e-9 | 6.7e36 | Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312 | 2004/11 | 24 | 53341 | 1.0e-9 | 6.4e36 | Šimon (2009), RXTE/ASM |
| GRS 1747-312 | 2005/4 | 25 | 53489 | 9.1e-10 | 5.8e36 | Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312 | 2005/9 | 26 | 53633 | 8.8e-10 | 5.6e36 | Šimon (2009), RXTE/ASM, PCA, IGR/JEM-X |
| GRS 1747-312 | 2006/2 | 27 | 53782 | 1.1e-9 | 6.9e36 | Chenevez et al. (2006); Šimon (2009), RXTE/ASM, PCA, IGR/JEM-X |
| GRS 1747-312 | 2006/7 | 28 | 53926 | 1.4e-9 | 8.8e36 | Šimon (2009) RXTE/ASM, PCA |
| GRS 1747-312 | 2007/1 | 29 | 54128 | 1.4e-9 | 8.9e36 | Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312 | 2007/6 | 30 | 54264 | 1.2e-9 | 7.9e36 | Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312 | 2007/11 | 31 | 54406 | 9.1e-10 | 5.8e36 | Šimon (2009), RXTE/ASM, PCA |
| GRS 1747-312 | 2008/3 | 32 | 54552 | 7.5e-10 | 4.8e36 | Šimon (2009), RXTE/ASM, PCA, IGR/JEM-X |
| GRS 1747-312 | 2008/9 | 54726 | 5.9e-10 | 3.7e36 | IGR/JEM-X | |
| GRS 1747-312 | 2009/2 | 54870 | 1.1e-9 | 7.1e36 | RXTE/ASM, PCA | |
| GRS 1747-312 | 2009/6 | 55008 | 1.1e-9 | 7.0e36 | RXTE/ASM, PCA | |
| Ter 6 X2 | 2009/9 | 55090 | 1.0e-11 | 6e34 | Saji et al. (2016); van den Berg et al. (2024) | |
| GRS 1747-312 | 2009/11 | 55147 | 1.1e-9 | 7.0e36 | RXTE/ASM, MAXI | |
| GRS 1747-312 | 2010/7 | 55402 | 7.3e-10 | 4.6e36 | MAXI, RXTE/PCA, ASM | |
| GRS 1747-312 | 2010/12 | 55544 | 1.0e-9 | 6.5e36 | MAXI | |
| GRS 1747-312 | 2011/4 | 55658 | 8.6e-10 | 5.5e36 | MAXI, RXTE/PCA | |
| GRS 1747-312 | 2011/7 | 55749 | 6.6e-10 | 4.2e36 | MAXI, RXTE/PCA | |
| GRS 1747-312 | 2011/9 | 55832 | 7.7e-10 | 4.9e36 | IGR/JEM-X, RXTE/PCA, MAXI | |
| GRS 1747-312 | 2011/12 | 55924 | 8.9e-10 | 5.6e36 | MAXI | |
| GRS 1747-312 | 2012/5 | 56072 | 6.6e-10 | 4.2e36 | MAXI | |
| GRS 1747-312 | 2012/10 | 56221 | 9.7e-10 | 6.1e36 | MAXI, IGR/JEM-X | |
| GRS 1747-312* | 2013/3 | 56369 | 7.9e-10 | 5.0e36 | Chenevez et al. (2013b); Bahramian et al. (2013a), Swift/XRT, MAXI, JEM-X | |
| GRS 1747-312 | 2013/7 | 56476 | 5.2e-10 | 3.3e36 | MAXI | |
| GRS 1747-312 | 2013/10 | 56594 | 7.0e-10 | 4.4e36 | Chenevez & Sanchez-Fernandez (2013),MAXI, IGR/JEM-X | |
| GRS 1747-312 | 2014/5 | 56786 | 1.2e-9 | 7.4e36 | MAXI | |
| GRS 1747-312 | 2014/10 | 56931 | 1.2e-9 | 7.7e36 | MAXI, IGR/JEM-X | |
| GRS 1747-312 | 2015/1 | 57030 | 1.2e-9 | 7.4e36 | MAXI | |
| GRS 1747-312 | 2015/3 | 57107 | 6.1e-10 | 3.9e36 | MAXI, IGR/JEM-X | |
| GRS 1747-312 | 2015/7 | 57227 | 5.1e-10 | 3.3e36 | MAXI | |
| GRS 1747-312 | 2015/12 | 57360 | 1.5e-9 | 9.4e36 | MAXI | |
| GRS 1747-312* | 2016/4 | 57505 | 2.2e-9 | 1.4e37 | Bahramian et al. (2016), MAXI, Swift/XRT | |
| GRS 1747-312 | 2016/9 | 57649 | 1.3e-9 | 8.3e36 | MAXI, IGR/JEM-X | |
| GRS 1747-312 | 2017/1 | 57768 | 5.7e-10 | 3.6e36 | MAXI | |
| GRS 1747-312 | 2017/5 | 57905 | 8.0e-10 | 5.1e36 | Swift/XRT, MAXI | |
| GRS 1747-312 | 2017/10 | 58035 | 1.0e-9 | 6.5e36 | Di Gesu et al. (2017); Bozzo & Sanna (2017), MAXI, Swift/XRT, IGR/JEM-X | |
| GRS 1747-312 | 2018/2 | 58155 | 1.3e-9 | 8.1e36 | MAXI, IGR/JEM-X | |
| GRS 1747-312 | 2018/6 | 58295 | 5.7e-10 | 3.6e36 | MAXI | |
| GRS 1747-312 | 2019/6 | 58643 | 1.1e-9 | 7.2e36 | van den Berg et al. (2024), Swift/XRT, MAXI | |
| GRS 1747-312* | 2019/10 | 58758 | 8.5e-10 | 5.4e36 | Grebenev et al. (2019); Bahramian et al. (2019), Swift/XRT, MAXI | |
| GRS 1747-312 | 2020/2 | 58876 | 5.5e-10 | 3.5e36 | MAXI | |
| GRS 1747-312 | 2020/5 | 58997 | 4.0e-10 | 2.5e36 | MAXI | |
| GRS 1747-312 | 2020/10 | 59137 | 6.6e-10 | 4.2e36 | MAXI | |
| GRS 1747-312 | 2021/2 | 59257 | 4.6e-10 | 2.9e36 | MAXI | |
| GRS 1747-312 | 2021/6 | 59383 | 8.6e-10 | 5.5e36 | van den Berg et al. (2024), Swift/XRT | |
| GRS 1747-312 | 2021/11 | 59536 | 1.4e-9 | 8.7e36 | MAXI | |
| GRS 1747-312 | 2022/3 | 59651 | 7.4e-10 | 4.7e36 | MAXI | |
| GRS 1747-312 | 2022/7 | 59774 | 5.3e-10 | 3.4e36 | MAXI, Swift/XRT | |
| GRS 1747-312 | 2022/10 | 59879 | 1.0e-11 | 6.6e35 | Swift/XRT | |
| GRS 1747-312 | 2023/1 | 59958 | 6.4e-10 | 4.1e36 | MAXI, Swift/XRT | |
| GRS 1747-312 | 2023/3 | 60034 | 6.1e-10 | 3.9e36 | MAXI | |
| GRS 1747-312 | 2023/6 | 60120 | 8.4e-10 | 5.3e36 | Yang et al. (2023), MAXI | |
| GRS 1747-312 | 2024/7 | 60508 | 8.4e-10 | 5.3e36 | MAXI | |
| GRS 1747-312 | 2025/3 | 60747 | 5.0e-10 | 3.2e36 | MAXI |
Outbursts of GRS 1747-312, the eclipsing burster in the globular cluster Terzan 6. in 2-10 keV ergs cm-2 s-1, in 2-10 keV ergs s-1. is about cm-2. D=7.3 kpc (Baumgardt & Vasiliev, 2021). E is epoch number of Šimon (2009). Saji et al. (2016) identify an outburst as likely produced by a different X-ray transient, labeled ”Ter 6 X-2?”. * indicates a reported eclipse during a pointed X-ray observation.) Reference to an instrument (e.g. MAXI) indicates we have made use of that instrument’s archive to generate some information for that line.
Note. —
Aquila X-1, 3U 1908+00 or 4U 1908+005. The distance is estimated at 4.5 kpc (Galloway et al., 2008). is typically estimated between 3e21 and 6e21 (e.g. Rutledge et al. 2002, Maccarone & Coppi 2003, Sakurai et al. 2012, Campana et al. 2014); we take =4e21 cm-2. We assume a power-law photon index of 2 for energy range conversions. We find conversions of 5.6e-9 from MAXI 2-20 keV countrate to 2-10 keV flux, 2.7e-10 from RXTE/ASM, 2.3e-12 from RXTE/PCA (for 5 units), 2.8e-11 for Swift/XRT. Aquila X-1’s outbursts have been studied in great detail in the literature, with significant compilations of outburst details by Kitamoto et al. (1993); Lin et al. (2007); Campana et al. (2013); Güngör et al. (2014); Ootes et al. (2018) among others. We list all known X-ray outbursts from Aquila X-1 in Table B. We identify one recent X-ray outburst from Aquila X-1 (in May 2020) that appears not to have been reported in the literature.
| NS XRB | Begin | End | Peak | References | |
| 2-10 keV | 2-10 keV | ||||
| Aql X-1 | 1965/4 | 1965/4 | 1.2e-9 | 2.9e36 | Friedman et al. (1967) (rocket) |
| Aql X-1 | 1969/7 | 1969/8 | 2e-9 | 5e36 | Kitamoto et al. (1993) (Vela5b) |
| Aql X-1 | 1969/12 | 1970/1 | 1.0e-8 | 2.4e37 | Kitamoto et al. (1993) (Vela5b) |
| Aql X-1 | 1970/8 | 1970/10 | 6.4e-9 | 1.5e37 | Price et al. (1972); Kitamoto et al. (1993) (Vela5b,rocket) |
| Aql X-1 | 1971/3 | 1971/4 | 5.2e-9 | 1.3e37 | Cominsky et al. (1978); Kitamoto et al. (1993) (Uhuru,Vela5b) |
| Aql X-1 | 1971/9 | 1971/11 | 1.1e-8 | 2.7e37 | Kitamoto et al. (1993) (Vela5b, OSO-7) |
| Aql X-1 | 1972/4 | 1972/5 | 6.4e-9 | 1.5e37 | Kitamoto et al. (1993) (Vela5b) |
| Aql X-1 | 1972/12 | 1973/3 | 1.2e-8 | 3e37 | Kitamoto et al. (1993) (Vela5b, OSO-7) |
| Aql X-1 | 1974/4 | 1974/6 | 2e-9 | 5e36 | Kitamoto et al. (1993) (Vela5b) |
| Aql X-1 | 1975/6 | 1975/7 | 2.0e-8 | 5e37 | Buff et al. (1977); Kaluzienski et al. (1977a); Kitamoto et al. (1993) (Vela5b,SAS-3,Ariel-5) |
| Aql X-1 | 1976/6 | 1976/7 | 2.0e-8 | 5e37 | Wilson et al. (1976); Kaluzienski et al. (1977a) (Ariel-5) |
| Aql X-1 | 1977/1 | 1977/1 | 6e-9 | 1.4e37 | Holt et al. (1977) (Ariel-5) |
| Aql X-1 | 1978/5 | 1978/8 | 2.6e-8 | 6.3e37 | Charles et al. (1980); Kitamoto et al. (1993); Chen et al. (1997) (Vela5b,Ariel-5) |
| Aql X-1 | 1979/3 | 1979/4 | 1e-8 | 2.4e37 | Kaluzienski & Holt (1979b); Czerny et al. (1987); Kitamoto et al. (1993) (Vela-5b,Ariel-5,Einstein) |
| Aql X-1 | 1980/5 | 1980/6 | 2.4e-9 | 6e36 | Koyama et al. (1981); Oda & Margon (1980) (Hakucho) |
| Aql X-1 | 1981/8 | 1981/8 | 1.4e-8 | 3.4e37 | Paradijs et al. (1981) (Hakucho) |
| Aql X-1 | 1987/3 | 1987/3 | 3.1e-8 | 7.5e37 | Kitamoto et al. (1993) (Ginga) |
| Aql X-1 | 1988/11 | 1988/11 | 1.1e-8 | 2.6e37 | Makino & GINGA Team (1988) (Ginga) |
| Aql X-1 | 1989/9 | 1989/10 | 1.2e-8 | 2.9e37 | Kitamoto et al. (1993) (Ginga) |
| Aql X-1 | 1990/6 | 1990/10 | - | - | Chevalier & Ilovaisky (1991); Hjellming et al. (1990) (optical, radio) |
| Aql X-1 | 1991/5 | 1991/10 | - | - | Ilovaisky & Chevalier (1991); Harmon et al. (1996) (optical,BATSE) |
| Aql X-1 | 1992/4 | 1992/5 | - | - | Ilovaisky & Chevalier (1992); Harmon et al. (1996) (optical,BATSE) |
| Aql X-1 | 1993/7 | 1993/8 | - | - | Ilovaisky & Chevalier (1993); Harmon et al. (1996) (optical,BATSE) |
| Aql X-1 | 1994/4 | 1994/7 | 5e-9 | 1.2e37 | Garcia et al. (1994); Rubin et al. (2000) (ASCA,BATSE) |
| Aql X-1 | 1995/7 | 1995/9 | - | - | Ilovaisky & Chevalier (1995, 1996) (optical) |
| Aql X-1 | 1996/2 | 1996/2 | 1.7e-9 | 4e36 | Campana et al. (2013) (RXTE/ASM) |
| Aql X-1 | 1996/5 | 1996/8 | 7.6e-10 | 1.8e36 | Ilovaisky & Chevalier (1996); Campana et al. (2013) (RXTE/ASM) |
| Aql X-1 | 1997/1 | 1997/4 | 7.6e-9 | 1.8e37 | Levine et al. (1997); Campana et al. (1998) (RXTE/ASM, PCA) |
| Aql X-1 | 1997/8 | 1997/9 | 4.4e-9 | 1.1e37 | Charles et al. (1997); Campana et al. (2013) (RXTE/ASM, PCA) |
| Aql X-1 | 1998/2 | 1998/5 | 9.6e-9 | 2.3e37 | Swank et al. (1998); Cui et al. (1998); Campana et al. (2013) (RXTE/ASM, PCA) |
| Aql X-1 | 1999/5 | 1999/11 | 7.5e-9 | 1.8e37 | Jain et al. (1999); Yu et al. (2003); Campana et al. (2013) (RXTE/ASM, PCA) |
| Aql X-1 | 2000/9 | 2000/11 | 1.3e-8 | 3.0e37 | Rutledge (2000); Yu et al. (2003); Lin et al. (2007); Campana et al. (2013) (RXTE/ASM, PCA) |
| Aql X-1 | 2001/6 | 2001/8 | 1.3e-9 | 3.1e36 | Bailyn et al. (2001); Yu & Dolence (2007); Campana et al. (2013) (RXTE/ASM, PCA) |
| Aql X-1 | 2002/2 | 2002/4 | 4.2e-9 | 1.0e37 | Swank & Smith (2002); Tudose et al. (2009); Campana et al. (2013) (RXTE/ASM, PCA) |
| Aql X-1 | 2003/2 | 2003/4 | 1.3e-8 | 3.1e37 | Swank et al. (2003); Campana et al. (2013) (RXTE/ASM, PCA) |
| Aql X-1 | 2004/2 | 2004/6 | 5.2e-9 | 1.2e37 | Molkov et al. (2004); Yu & Dolence (2007); Tudose et al. (2009); Campana et al. (2013) (RXTE/ASM, PCA) |
| Aql X-1 | 2005/3 | 2005/5 | 1.3e-9 | 3.2e36 | Grebenev et al. (2005a); Rodriguez et al. (2006); Tudose et al. (2009); Campana et al. (2013) (RXTE/ASM, PCA, IGR) |
| Aql X-1 | 2005/11 | 2005/11 | 1.4e-9 | 3.3e36 | Rodriguez & Shaw (2005); Campana et al. (2013) (RXTE/ASM, PCA) |
| Aql X-1 | 2006/7 | 2006/8 | 1.1e-9 | 2.7e36 | Wijnands et al. (2006a); Campana et al. (2013) (RXTE/ASM, Swift/XRT) |
| Aql X-1 | 2007/5 | 2007/6 | 1.3e-9 | 3.0e36 | Rodriguez et al. (2007); Campana et al. (2013) (RXTE/ASM, Swift/XRT) |
| Aql X-1 | 2007/9 | 2007/10 | 5e-9 | 1.2e37 | Corbet et al. (2007); Sakurai et al. (2012); Campana et al. (2013) (RXTE/ASM, Suzaku) |
| Aql X-1 | 2008/5 | 2008/6 | 1.1e-9 | 2.6e36 | Palmer et al. (2008); Campana et al. (2013) (RXTE/ASM) |
| Aql X-1 | 2009/3 | 2009/5 | 8.8e-10 | 2.1e36 | Russell & Lewis (2009); Campana et al. (2013) (RXTE/ASM) |
| Aql X-1 | 2009/11 | 2009/12 | 6.2e-9 | 1.4e37 | Linares et al. (2009b); Miller-Jones et al. (2010); Asai et al. (2012) (RXTE/ASM,PCA,MAXI,Swift/XRT) |
| Aql X-1 | 2010/7 | 2010/10 | 7.0e-9 | 1.7e37 | Gultekin et al. (2010); Asai et al. (2012); Campana et al. (2014) (MAXI,RXTE/ASM,PCA,Swift/XRT) |
| Aql X-1 | 2011/10 | 2011/10 | 1.6e-8 | 3.8e37 | Yamaoka et al. (2011); Asai et al. (2012); Güngör et al. (2014); Waterhouse et al. (2016) (MAXI,PCA) |
| Aql X-1 | 2013/6 | 2013/6 | 2.0e-8 | 4.8e37 | Degenaar & Wijnands (2013c); Waterhouse et al. (2016); Meshcheryakov et al. (2018) (MAXI) |
| Aql X-1 | 2014/6 | 2014/7 | 6.4e-9 | 1.5e37 | Meshcheryakov & Tsygankov (2014); Gandhi et al. (2014); King et al. (2016) (MAXI, Swift/XRT) |
| Aql X-1 | 2015/2 | 2015/4 | 1.8e-9 | 4.3e36 | Ueno et al. (2015); Waterhouse et al. (2016) (MAXI) |
| Aql X-1 | 2016/7 | 2016/8 | 1.8e-8 | 4.3e37 | Sanna et al. (2016); Díaz Trigo et al. (2018); Degenaar et al. (2019b); Niwano et al. (2023) (MAXI,Swift/XRT) |
| Aql X-1 | 2017/5 | 2017/6 | 1.4e-9 | 3.4e36 | Dinçer et al. (2017); Niwano et al. (2023) (MAXI) |
| Aql X-1 | 2018/2 | 2018/3 | 8.2e-9 | 2.0e37 | Spiridonova et al. (2018); Niwano et al. (2023) (MAXI) |
| Aql X-1 | 2018/11 | 2018/12 | 1e-9 | 2e36 | Lin et al. (2018); Niwano et al. (2023) (MAXI, Swift/XRT) |
| Aql X-1 | 2019/8 | 2019/9 | 1.1e-8 | 2.7e37 | Motta et al. (2019); Niwano et al. (2023) (MAXI) |
| Aql X-1 | 2020/5 | 2020/5 | 3.1e-9 | 7.5e36 | MAXI |
| Aql X-1 | 2020/8 | 2020/10 | 1.0e-8 | 2.4e37 | Saikia et al. (2020); Bahramian et al. (2020a) (MAXI,Swift/XRT) |
| Aql X-1 | 2021/3 | 2021/4 | 1.7e-9 | 4.0e36 | Kong & Zwicky Transient Facility Collaboration (2021), MAXI, Swift/XRT |
| Aql X-1 | 2021/11 | 2022/1 | 1.2e-8 | 2.8e37 | Niwano et al. (2021), MAXI |
| Aql X-1 | 2022/5 | 2022/6 | 6.5e-9 | 1.6e37 | Pawar (2022),MAXI |
| Aql X-1 | 2023/7 | 2023/9 | 1.8e-8 | 4.3e37 | Alabarta et al. (2023a, b); Yan et al. (2025),MAXI |
| Aql X-1 | 2024/9 | 2024/10 | 1.1e-8 | 2.7e37 | Liu et al. (2024); Mandal et al. (2025), MAXI |
Note. — Outbursts of Aquila X-1. in 2-10 keV ergs cm-2 s-1, in 2-10 keV ergs s-1. is about cm-2. D=4.5 kpc (Galloway et al., 2008). Reference to an instrument (e.g. MAXI) indicates we have made use of that instrument’s archive to generate some information for that line.
Cir X-1 has a distance estimate of kpc from the geometry of light echoes (Heinz et al., 2015). Its eccentric orbit makes its behavior different from other systems here, and it is arguable whether it is a low-mass or high-mass X-ray binary. Fig. 11 illustrates its overall behavior since 1996, using RXTE/ASM, MAXI/GSC and Swift/XRT data. (See Saz Parkinson et al. 2003 for an excellent summary of earlier data.). In Table B we give details on its short outbursts.
| NS XRB | Date | MJD | Peak | References | |
| 2-10 keV | 2-10 keV | ||||
| Cir X-1 | 1965/4 | 38875 | 5.8e-9 | 6.1e37 | Amnuel et al. (1982) (rocket) |
| Cir X-1 | 1969/6 | 40386 | 1.2e-8 | 1.3e38 | Amnuel et al. (1982) (rocket) |
| Cir X-1 | 1970/5 | 40719 | 1.2e-8 | 1.3e38 | Amnuel et al. (1982) (rocket) |
| Cir X-1 | 1971/2 | 40996 | 2.8e-8 | 3.0e38 | Forman et al. (1976b) |
| Cir X-1 | 1972/5 | 41459 | 1.5e-8 | 1.6e38 | Canizares et al. (1974) |
| Cir X-1 | 1973/8 | - | 6.3e-10 | 6.7e36 | Amnuel et al. (1982) |
| Cir X-1 | 1974/4 | - | 5.8e-9 | 6.1e37 | Amnuel et al. (1982); Davison & Tuohy (1975) |
| Cir X-1 | 1976/1 | 42793 | 2.0e-8 | 2.1e38 | Buff et al. (1977) |
| Cir X-1 | 1976/2 | 42810 | 2.6e-8 | 2.7e38 | Kaluzienski et al. (1976) |
| Cir X-1 | 1976/2 | 42825 | 1.4e-8 | 1.5e38 | Wilson & Carpenter (1976) |
| Cir X-1 | 1976/10 | 43075 | 1.6e-8 | 1.7e38 | Dower et al. (1982) |
| Cir X-1 | 1976/12 | 43124 | 6.7e-9 | 7.1e37 | Dower et al. (1982) |
| Cir X-1 | 1977/1 | 43158 | 1.4e-8 | 1.5e38 | Dower et al. (1982) |
| Cir X-1 | 1977/5 | 43273 | 4.5e-9 | 4.8e37 | Dower et al. (1982) |
| Cir X-1 | 1977/8 | 43383 | 3.1e-8 | 3.3e38 | Kaluzienski et al. (1977b) |
| Cir X-1 | 1977/9 | 43397 | 2.3e-8 | 2.4e38 | Kaluzienski & Holt (1977) |
| Cir X-1 | 1977/9 | 43407 | 3.2e-8 | 3.4e38 | Dower et al. (1982) |
| Cir X-1 | 1977/10 | 43423 | 1.2e-8 | 1.3e38 | Dower et al. (1982) |
| Cir X-1 | 1978/1 | 43524 | 6.8e-8 | 7.2e38 | Kaluzienski & Holt (1978) |
| Cir X-1 | 1978/2 | 43541 | 9.1e-9 | 9.6e37 | Kaluzienski & Holt (1978); Murdin et al. (1980) |
| Cir X-1 | 1978/8 | 43737 | 9.2e-9 | 9.7e37 | Dower et al. (1982) |
| Cir X-1 | 1979/1 | 43888 | 4.6e-8 | 4.9e38 | Kaluzienski & Holt (1979c) |
| Cir X-1 | 1984/5 | 45843 | 2.3e-8 | 2.4e38 | Inoue (1989); Miyamoto & Kitamoto (1985) |
| Cir X-1 | 1987/2 | - | 3.0e-8 | 3.2e38 | Tsunemi et al. (1989) |
| Cir X-1 | 1989/4 | 47617 | 8.6e-8 | 9.1e38 | Stewart et al. (1991) |
| Cir X-1 | 2006/1 | 53761 | 3.4e-9 | 3.6e37 | RXTE/ASM |
| Cir X-1 | 2006/3 | 53805 | 8.3e-9 | 8.8e37 | RXTE/ASM |
| Cir X-1 | 2006/6 | 53890 | 4.2e-9 | 4.4e37 | Jonker et al. (2006b),RXTE/ASM |
| Cir X-1 | 2006/7 | 53938 | 1.5e-8 | 1.6e38 | RXTE/ASM |
| Cir X-1 | 2006/8 | 53968 | 3.4e-8 | 3.6e38 | RXTE/ASM |
| Cir X-1 | 2006/9 | 54001 | 4.8e-8 | 5.0e38 | RXTE/ASM |
| Cir X-1 | 2006/12 | 54074 | 5.0e-9 | 5.2e37 | RXTE/ASM |
| Cir X-1 | 2006/12 | 54085 | 8.1e-9 | 8.6e37 | RXTE/ASM |
| Cir X-1 | 2007/1 | 54106 | 4.9e-9 | 5.2e37 | Nicolson (2007),RXTE/ASM |
| Cir X-1 | 2007/11 | 54430 | 1.6e-8 | 1.7e38 | RXTE/ASM |
| Cir X-1 | 2008/4 | 54558 | 4.6e-9 | 4.9e37 | RXTE/ASM |
| Cir X-1 | 2009/9 | 55100 | 4.4e-10 | 4.7e36 | MAXI |
| Cir X-1 | 2009/11 | 55156 | 8.0e-10 | 8.5e36 | Asai et al. (2014), MAXI |
| Cir X-1 | 2010/5 | 55324 | 6.6e-9 | 7.0e37 | Asai et al. (2014); D’Aì et al. (2012); Nakajima et al. (2010) |
| Cir X-1 | 2010/6 | 55360 | 1.6e-9 | 1.7e37 | D’Aì et al. (2012); Asai et al. (2014); Isobe et al. (2010) |
| Cir X-1 | 2010/7 | 55382 | 5.8e-10 | 6.1e36 | Krimm et al. (2011), MAXI |
| Cir X-1 | 2010/8 | 55438 | 8.0e-9 | 8.4e37 | MAXI, RXTE/ASM |
| Cir X-1 | 2010/11 | 55520 | 6.7e-10 | 7.1e36 | MAXI |
| Cir X-1 | 2011/2 | 55593 | 1.4e-8 | 1.5e38 | Asai et al. (2014),RXTE/ASM,Krimm et al. (2011) |
| Cir X-1 | 2011/3 | 55642 | 1.2e-9 | 1.3e37 | Asai et al. (2014),MAXI |
| Cir X-1 | 2011/5 | 55679 | 2.9e-10 | 3.1e36 | MAXI,Krimm et al. (2011) |
| Cir X-1 | 2011/6 | 55732 | 9.5e-10 | 1.0e37 | MAXI |
| Cir X-1 | 2011/7 | 55753 | 3.2e-8 | 3.4e38 | Asai et al. (2014); Kimura et al. (2012),RXTE/ASM,MAXI |
| Cir X-1 | 2011/8 | 55800 | 9.7e-9 | 1.0e38 | Asai et al. (2014),RXTE/ASM |
| Cir X-1 | 2011/10 | 55840 | 5.8e-10 | 6.1e36 | Asai et al. (2014),MAXI |
| Cir X-1 | 2011/10 | 55862 | 1.7e-9 | 1.8e37 | Asai et al. (2014),MAXI |
| Cir X-1 | 2011/12 | 55908 | 4.7e-9 | 5.0e37 | Asai et al. (2014) |
| Cir X-1 | 2012/1 | 55936 | 3.4e-9 | 3.6e37 | Asai et al. (2014) |
| Cir X-1 | 2012/4 | 56021 | 9.9e-10 | 1.0e37 | MAXI |
| Cir X-1 | 2012/6 | 56090 | 1.1e-9 | 1.2e37 | MAXI |
| Cir X-1 | 2012/6 | 56105 | 5.1e-10 | 5.4e36 | Swift/XRT |
| Cir X-1 | 2012/7 | 56112 | 8.6e-10 | 9.1e36 | Asai et al. (2014) |
| Cir X-1 | 2012/8 | 56154 | 1.0e-8 | 1.1e38 | Asai et al. (2014) |
| Cir X-1 | 2012/12 | 56280 | 3.5e-9 | 3.7e37 | MAXI,Kimura et al. (2012) |
| Cir X-1 | 2012/12 | 56287 | 1.6e-9 | 1.7e37 | Asai et al. (2014) |
| Cir X-1 | 2013/3 | 56379 | 1.3e-9 | 1.4e37 | MAXI |
| Cir X-1 | 2013/8 | 56528 | 1.2e-9 | 1.3e37 | MAXI |
| Cir X-1 | 2013/10 | 56576 | 2.0e-9 | 2.1e37 | Asai et al. (2014) |
| Cir X-1 | 2013/10 | 56585 | 2.4e-8 | 2.5e38 | Asai et al. (2014) |
| Cir X-1 | 2013/11 | 56612 | 3.2e-8 | 3.4e38 | Asai et al. (2014) |
| Cir X-1 | 2014/8 | 56891 | 1.1e-9 | 1.1e37 | MAXI |
| Cir X-1 | 2014/11 | 56964 | 1.0e-9 | 1.1e37 | MAXI |
| Cir X-1 | 2015/1 | 57033 | 1.1e-9 | 1.2e37 | MAXI |
| Cir X-1 | 2015/3 | 57111 | 9.5e-9 | 1.0e38 | Zhang et al. (2015),MAXI |
| Cir X-1 | 2015/4 | 57119 | 1.7e-8 | 1.8e38 | MAXI |
| Cir X-1 | 2015/8 | 57248 | 6.4e-9 | 6.8e37 | MAXI |
| Cir X-1 | 2015/12 | 57367 | 9.4e-9 | 9.9e37 | MAXI |
| Cir X-1 | 2016/1 | 57404 | 1.0e-8 | 1.1e38 | MAXI |
| Cir X-1 | 2016/2 | 57436 | 8.8e-9 | 9.3e37 | MAXI |
| Cir X-1 | 2016/5 | 57533 | 5.8e-9 | 6.1e37 | MAXI |
| Cir X-1 | 2016/7 | 57598 | 3.4e-9 | 3.5e37 | Swift/XRT |
| Cir X-1 | 2016/9 | 57655 | 9.0e-9 | 9.5e37 | Sanna & Motta (2016),Swift/XRT |
| Cir X-1 | 2016/12 | 57731 | 3.8e-9 | 4.0e37 | MAXI |
| Cir X-1 | 2017/1 | 57777 | 3.9e-9 | 4.1e37 | MAXI |
| Cir X-1 | 2017/4 | 57846 | 6.3e-9 | 6.6e37 | MAXI |
| Cir X-1 | 2017/8 | 57976 | 1.4e-9 | 1.4e37 | MAXI |
| Cir X-1 | 2017/9 | 58024 | 1.2e-8 | 1.3e38 | MAXI |
| Cir X-1 | 2017/10 | 58037 | 1.4e-8 | 1.5e38 | MAXI |
| Cir X-1 | 2017/12 | 58112 | 2.1e-9 | 2.2e37 | MAXI |
| Cir X-1 | 2018/1 | 58135 | 1.4e-9 | 1.5e37 | MAXI |
| Cir X-1 | 2018/2 | 58172 | 4.0e-9 | 4.2e37 | MAXI |
| Cir X-1 | 2018/3 | 58193 | 1.3e-9 | 1.4e37 | MAXI |
| Cir X-1 | 2018/7 | 58317 | 7.0e-9 | 7.4e37 | MAXI |
| Cir X-1 | 2018/8 | 58358 | 1.6e-9 | 1.6e37 | MAXI |
Note. — Recorded outbursts of Cir X-1. in 2-10 keV ergs cm-2 s-1, in 2-10 keV ergs s-1. is about cm-2. D=9.4 kpc (Heinz et al., 2015). Reference to an instrument (e.g. MAXI) indicates we have made use of that instrument’s archive to generate some information for that line.
Appendix C Bursters with no outbursts
In Table 4 we give details on the four known X-ray bursters which have no recorded outbursts.
SAX J1752.3-3138; the distance is estimated at 9.2 kpc by Cocchi et al. (2001).
SAX J1818.7+1424; Cornelisse et al. (2002b) place a limit of 9.4 kpc from an X-ray burst.
AX J1824.5-2451; Gotthelf & Kulkarni (1997) identified an X-ray burst from M28 using ASCA in 1997 (identified as AX J1824.5-2451), at a time when the total X-ray luminosity of the cluster was only ergs s-1.
SAX J2224.9+5421; the distance upper limit of 7.1 kpc comes from the discovery burst (Cornelisse et al., 2002b).
| NS XRB | Year/Month | Peak | d | References | ||
| 2-10 keV | cm-2 | 2-10 keV | kpc | |||
| SAX J1752.3-3138 | 1999/9 | 1.3e-10 | 6e21 | 1.3e36 | 9.2 | Cocchi et al. (2001) |
| SAX J1818.7+1424 | 1997/8 | 1.1e-10 | 1e21 | 1.2e36 | 9.4 | Cornelisse et al. (2002b) |
| AX J1824.5-2451 | 1995/3 | 3e-13 | 6e21 | 1e33 | 5.5 | Gotthelf & Kulkarni (1997) |
| SAX J2224.9+5421 | 1999/11 | 1.3e-13 | 5e21 | 8e32 | 7.1 | Cornelisse et al. (2002b); Degenaar et al. (2014b) |
Appendix D Known persistent neutron star low-mass X-ray binaries
In Table 10 we list the neutron star X-ray binaries we take to be persistent systems.
| Persistent LMXB | Alt. name | Nature |
| 4U 0513-40 | burster | |
| 4U 0614+09 | burster | |
| 2S 0918-549 | 4U 0919-54 | burster |
| 4U 1246-588 | 1A 1246-588 | burster |
| 4U 1254-69 | burster | |
| 4U 1323-62 | burster | |
| 4U 1543-624 | burster | |
| UW CrB | 1E 1603.6+2600 | burster |
| 4U 1626-67 | pulsar | |
| 4U 1636-536 | burster | |
| 4U 1702-429 | burster | |
| 4U 1708-23 | burster | |
| 4U 1705-32 | 1RXS J170854.4-321857 | burster |
| 4U 1705-44 | burster | |
| 4U 1708-40 | burster | |
| XTE J1710-281 | burster | |
| SAX J1712.6-3739 | burster | |
| RX J1718.4-4029 | burster | |
| IGR J17254-3257 | burster | |
| 4U 1722-30 | 4U 1724-307 | burster |
| 4U 1728-34 | burster | |
| 1RXH J173523.7-354013 | burster | |
| SLX 1735-269 | burster | |
| 4U 1735-444 | burster | |
| SLX 1737-282 | burster | |
| 1A 1742-294 | burster | |
| GX 3+1 | 4U 1744-26 | burster |
| SLX 1744-299 | burster | |
| SLX 1744-300 | burster | |
| 4U 1746-37 | burster | |
| GX 13+1 | V5512 Sgr, 4U 1811-17 | burster |
| 4U 1812-12 | burster | |
| GX 17+2 | NP Ser, 4U 1813-14 | burster |
| 4U 1820-303 | 4U 1820-30 | burster |
| 4U 1822-000 | 4U 1823-00 | burster |
| 2A 1822-371 | pulsar | |
| GS 1826-24 | 4U 1826-24, GS 1826-238 | burster |
| XB 1832-330 | NGC 6652 A | burster |
| Ser X-1 | 4U 1837+04, 3A 1837+049 | burster |
| 4U 1850-086 | 4U 1850-087 | burster |
| XB 1916-053 | 4U 1916-05 | burster |
| M15 X-2 | 4U 2129+12aa4U 2129+12 refers to both bright X-ray sources in M15. AC 211 in M15 is generally believed to contain a NS as well (van Zyl et al., 2004), but we do not include it. | burster |
| Cyg X-2 | 4U 2142+38 | burster |